### 28 Dec 2023 v1

# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

Abstract
Emergence and causality are two fundamental concepts for understanding complex systems. They are interconnected. On one hand, emergence refers to the phenomenon where macroscopic properties cannot be solely attributed to the cause of individual properties. On the other hand, causality can exhibit emergence, meaning that new causal laws may arise as we increase the level of abstraction. Causal emergence theory aims to bridge these two concepts and even employs measures of causality to quantify emergence. This paper provides a comprehensive review of recent advancements in quantitative theories and applications of causal emergence. Two key problems are addressed: quantifying causal emergence and identifying it in data. Addressing the latter requires the use of machine learning techniques, thus establishing a connection between causal emergence and artificial intelligence. We highlighted that the architectures used for identifying causal emergence are shared by causal representation learning, causal model abstraction, and world model-based reinforcement learning. Consequently, progress in any of these areas can benefit the others. Potential applications and future perspectives are also discussed in the final section of the review.
keywords:
Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model
[…]
1 Introduction
Economic growth, environmental protection, sustainable development, the global climate crisis, social inequality, and many other issues are all intertwined with complex systems Ledford (2015); Mensah et al. (2018). Therefore, gaining a deep understanding of how complex systems operate, evolve, grow, stabilize, and collapse is of paramount importance. However, this task is exceptionally challenging due to the fact that complex systems consist of diverse and heterogeneous agents that interact through complex nonlinear relationships Bar-Yam (2002). Moreover, they all exhibit emergent phenomena, which are highly common in complex systems but carry a sense of mystery Holland (2000).
How did the first living cell emerge from the collisions between various molecules in the early Earth’s environment Artime and De Domenico (2022)? How does the cognitive concept of "I" emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)? How do large neural language models suddenly exhibit emergent abilities Wei et al. (2022)? These fundamental questions revolve around the concept of emergence in complex living, cognitive, and artificial systems. Emergence refers to the phenomenon where macroscopic properties and phenomena that cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000). This presents a formidable challenge to the traditional reductionist perspective, while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
However, as elucidated by Bedau’s theory of weak emergence Bedau (1997), many emergent phenomena can be comprehended through the interactions among the individuals within the system Holland (2000). Complex systems, in fact, consist of extensive networks of interacting components Albert and Barabási (2002). Within these networks, even a minor cause, such as a perturbation of a single unit, can propagate through the interconnected network, resulting in a collective effect. The phenomenon described is commonly known as the butterfly effect Wikipedia (2023a), which provides an explanation for the occurrence of emergence. On the other hand, emergent properties, such as homeostasis Wikipedia (2023b), can stabilize the system itself, thereby preserving the locality of causal effects and preventing the observation of macro-level effects. These phenomena demonstrate that complex systems achieve interactions through causal laws, where numerous local causal laws form interconnected causal networks as a whole. And this whole possesses its unique causal characteristics.
Causality, or causation, refers to the connection between a cause and its resulting effect Granger (1969); Pearl et al. (2000); Pearl (2009). It describes the phenomenon in which an event, known as the cause, leads to another event, known as the effect. Traditional studies of causality have typically focused on the causal relationship between two or a few variables. However, the unique characteristics of causality in complex systems present new challenges to classical causal science due to the vast number of variables involved and the presence of emergent phenomena. In complex systems, it is possible for one cause to have multiple effects, and conversely, one effect may be influenced by a multitude of causes. Furthermore, in complex systems, causality often exhibits cross-level properties, which are closely associated with emergence.
Emergence and causation interconnect with each other. On the one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956). On the other hand, emergent properties may have causal effects on individuals in complex systems Kim (1992); Holland (2000). For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market, while the price may also give feedback to the market: it may affect the decision making of each individual.
Furthermore, we can understand emergence through the perspective of causation. What does emergence mean is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005). Thus, emergent properties or phenomena lose their direct explanations as usual, but may be attributed to the causes on the macro-level as pointed out by Hoel et al. (2013). Therefore, new causalities can be observed on larger scales.
In conclusion, gaining a deep understanding of emergence is crucial in the field of complex system studies. Specifically, the development of a quantitative theory of emergence is on the verge of emerging. Such a theory holds the potential to address significant challenges, including the origins of life, the emergence of novel capabilities in large neural network models, and the potential for intelligence, consciousness, and free will to arise in artificial systems. Causality not only exhibits a profound connection with emergence but is also considered by many researchers as one of the most crucial perspectives for quantitatively comprehending emergence Hoel et al. (2013); Crutchfield and Young (1989); Fromm (2005); Seth (2008).
Two primary challenges take precedence in understanding emergence from a causal perspective. The first is establishing a quantitative definition of emergence, while the second involves identifying emergent behaviors or phenomena through data analysis.
To address the first challenge, two prominent quantitative theories of emergence have emerged in the past decade. The first is Erik Hoel’s theory of causal emergence Hoel et al. (2013), while the second is Rosas’ theory of emergence based on partial information decomposition Rosas et al. (2020).
Hoel’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains. It employs the concept of effective information (EI) to quantify the extent of causal influence within Markov chains and enables comparisons of EI values across different scales Hoel et al. (2013); Hoel (2017). Causal emergence is defined by the difference of EI values between macro-level and micro-level. Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than the micro-level. Hoel and other researchers further extended measures of effective information and causal emergence for dynamical systems with continuous variables in Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020). Other measures are also possible to quantify causal effects and thereafter causal emergence. In Comolatti and Hoel (2022), Hoel systematically compared a bunch of causal effect measures and concluded that causal emergence is independent on the selection of the measure.
However, it is essential to establish a coarse-graining strategy beforehand in Hoel’s theory of causal emergence. Alternatively, the strategy can be derived by maximizing the effective information (EI) Hoel et al. (2013). However, this task becomes challenging for large-scale systems due to the computational complexity involved. To address these problems, Rosas introduced a new quantitative definition of causal emergence Rosas et al. (2020) that does not depend on coarse-graining methods, drawing from Partial Information Decomposition (PID) related theory. PID is an approach developed by Williams et al., which seeks to decompose the mutual information between a target and source variables into non-overlapping information atoms: unique, redundant, and synergistic information Williams and Beer (2010). Based on this groundwork, Rosas further developed the concept and introduced a theory called $\phi$ID to decompose the mutual information between multiple target and source variables Mediano et al. (2019). This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between source and target variables based on the inherent characteristics of the system.
The second challenge pertains to the identification of emergence from data. In an effort to tackle this issue, Rosas et al. have derived a numerical method Rosas et al. (2020). However, it is important to acknowledge that this method offers only a sufficient condition for emergence and is an approximate approach. Another limitation is a coarse-grained macro-state variable should be given beforehand to apply this method. Hence, there is a need for the development of new methods. In a recent study by Zhang et al. Zhang and Liu (2022), a machine learning framework named Neural Information Squeezer (NIS) was introduced to tackle the challenge of identifying causal emergence using Hoel’s definition. Remarkably, the NIS neural network, functioning as a "machine observer" equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data. This aligns with the viewpoint suggested by Eric Bonabeau and Jean-Louis Dessalles Bonabeau and Dessalles (2011) that all emergent properties and phenomena are observer-dependent, and the epistemological nature of emergent causal powers, including downward causation Kim (1992); Crutchfield and Young (1989). However, employing a machine observer offers additional advantages in identifying emergence from data, as it helps avoid subjective judgments.
Furthermore, we find that causal emergence identification and the recent development of artificial intelligence can have deeper connections through the bridge of NIS and its extended frameworks Yang et al. (2023). First, the architecture of NIS, that is, the “encoder - dynamics learner - decoder” framework, is similar to causal representation learning Schölkopf et al. (2021) in artificial intelligence where dynamics learner is replaced as causal mechanisms. Therefore, there is a deep connection between the two emerging fields: causal emergence and causal representation learning. The former tries to characterise and understand the emergent causality in a complex system and the latter tries to learn the latent causal structures and mechanisms from observational data. They are connected by the problem of causal emergence identification. That is to identify causal emergence from data is essentially equivalent to learning the latent causality inherent in the data.
Second, there are more connections between causal emergence and other areas in artificial intelligence. For example, causal model abstraction also discussed the interplay between causality and the abstraction (coarse-graining) of causal models Beckers et al. (2020). In reinforcement learning based on world models Ha and Schmidhuber (2018), the similar framework, “encoder - dynamics learner - decoder”, is also used. Thus, the similarity and the shared features between the two fields can help us to lend the ideas and techniques in one field to the other field. For instance, EI may be used to measure the causal effect of neural networks Marrow et al. (2020), and the techniques in causal representation learning can be employed to detect and identify causal emergence. Zhang and Liu (2022).
This paper tries to review the latest works on the quantitative theory and applications of causal emergence, and its relative works. We also discuss its connections with artificial intelligence and complex systems. In the second part, we introduce the background of causal emergence, particularly on the connections between causation and emergence in complex systems. The third part introduces several quantitative theoretic frameworks including: Crutchfield et al’s computational mechanics Crutchfield and Young (1989), Seth et al’s granger causal emergence Seth (2008), Hoel et al’s causal emergence theory and Rosas et al’ theory of emergence based on information decomposition. Several related conceptions are discussed, such as coarse-graining strategy, measure of effective information, and partial information decomposition. We also compare these two theories from several aspects. The forth part discusses the problem of causal emergence identification and then introduces the connection with artificial intelligence. Causal representation learning Schölkopf et al. (2021), abstraction of causal models Beckers et al. (2020), and world models in reinforcement learning Ha and Schmidhuber (2018) are introduced and the relations with causal emergence identification are discussed. Finally, we discuss the connections between causal emergence and other important topics in complex systems such as the origins of life, consciousness, etc.
2 Background: from Causality to Emergence
2.1 Causality
Causality, which is a fundamental concept in many fields, including philosophy, natural science, and social science, refers to the relationship between a pair of events, where the first event (cause) can influence the second one (effect) Granger (1969); Pearl et al. (2000); Pearl (2009). The relationship between cause and effect can exhibit either deterministic or probabilistic characteristics. In deterministic causality, the cause will always produce the same effect, while in probabilistic causality, the cause will only produce the effect with a certain probability.
In most physical systems governed by differential equations or Markovian dynamics, whether deterministic or probabilistic, causality is inherent. This arises from the fact that manipulating a variable within the system can result in observable changes in other variables, albeit with probabilistic outcomes. While, one physical mechanism may correspond several possible causal structures as pointed by Y. Iwasaki and H. A. Simon Iwasaki and Simon (1994) because the causal graphs describe local causal interactions while mechanisms are the global constraints for all variables.
Figure 1: The causal models (left column) and the corresponding causal hierarchy (right column) and .
However, if we do not know the physical mechanism of a system, how can we model the causality behind the system? There are several parallel theoretical frameworks to do this. For example, Judea Pearl tries probabilistic graphic models (e.g. Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al. (2000); Pearl (2009). In which nodes represent random variables, acyclic links represent causal interactions. Three different causal hierarchies are distinguished and quantified in different models. The first level of causality is association, i.e., the correlation between variables, which can be modelled by Bayesian networks Sucar (2015). The second level is intervention, which can be characterised by causal graph models Pearl (2009). The difference between a causal graph and a Bayesian network is the former allows defining a special operator which is called "do" operator to simulate the interventions done by the experimenter. The last level is courterfactual, which can be quantified by structural causal models Pearl et al. (2000). The distinguishing characteristic of a structural causal model is that all interactions within a causal graph can be described by a set of deterministic functions that involve the input variables and unknown components, thereby introducing uncertainty. Counterfactual causal inferences, that is the inferences under an imaginary scenario, can be conducted only on structural causal models. The different types of causal models and the corresponding causal hierarchy in Pearl’s theory are shown in Figure 1.
Another quantitative framework for causal inference is Donald B Rubin’s "potential outcome" theory which is directly defined on the level of counterfactual Rubin (2005). According to this theory, each individual has two potential outcomes: the outcome that would be observed when individuals receive the treatment and the outcome that would be observed when they do not receive the treatment. Rubin’s theory suggests that the causal effect of a treatment can be estimated by comparing the difference between the observed outcome and the unobserved potential outcome for each individual.
Researchers also developed a bundle of methods, which are named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data. For example, constraint-based algorithms Spirtes et al. (2000) and score-based algorithms Chickering (2002) for causal discovery are two typical representations . Causal discovery has become a hot interdisciplinary area in the middle of machine learning, causal inference, and statistics.
Due to the uncertainty and ambiguity behind the discovered causal relations, measuring the degree of causal effect between two variables is another important problem. Numerous independent historical studies have tackled the issue of causation measurement. These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009), etc. The effective information (EI) in the framework of causal emergence is also an indicator to measure causal effect. All these measures can almost be decomposed into two terms which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
In summary, causality is a fascinating inter-disciplinary topic across for several subjects. Quantitative framework and measures have been developed in last decades.
2.2 Emergence
Emergence, known as "the whole is greater than the sum of parts" Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro and macro levels of organization. Therefore, the scale of a system (micro or macro) and the cross-level interactions must be considered when we discuss emergence. This may increase the complexity of study for emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al. (2019); Ross et al. (1908); Holland (1996)
Emergence has deep connections with causation Fromm (2005); Bedau (1997); Bedau and Humphreys (2008); Bedau (2002). First, emergence emphasizes the new properties arising from the interactions between the components in the system, rather than the components themselves. Consequently, all emergent properties and processes at the macro-level can, in principle, be comprehended as the causal effects resulting from interactions between individuals at the micro-level. This is what the philosophical notion “supervenience” tries to describe.
However, it does not mean that all the emergent phenomena on macro-level can be easily attributed to micro-level individuals and their interactions. Two major reasons can account for this. The first reason is due to the principle of computational irreducibility proposed by Wolfram Wolfram et al. (2002) which claims that although we know the emergent properties are out of a set of simple rules, we can not find a short-cut way in principle to predict the results except to implement the rules. For example, complex behaviors (eg. herding and aligning) of the Boid model are the results of simple rules, but these behaviors can not be reduced to individual behaviors in a simple way except to simulate it Reynolds (1987).
The second reason is that emergent behaviors may be determined by other emergent properties on macro-level Hoel et al. (2013); Rosas et al. (2020). For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because the severe lack of fossil oil may cause inflation. Therefore, the cause of an emergent property (rice price) is another emergent macro-level coarse-grained variable and it means new causation on macro-level coarse-grained variables can be observed as pointed by the theory of causal emergence. And, interestingly, the causation in different levels must be consistent to explain the micro-level event of the price of rice. Otherwise, double causation Merton (1973) fallacy may appear. This consistency between different scale causality is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
In order to better understand emergence, Bedau et al Bedau (1997, 2002) classify emergence into three categories according to the causal interactions between micro and macro, which are nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996) and strong emergence O’Connor (1994); Kim (1992).
In all these categories, the notion of nominal emergence is least controversial, which can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994). For example, the patterns which are composed of many pixels on a screen are nominal emergence. We can say nominal emergent patterns as “supervenient” because all the macro-level properties are from individuals.
While weak emergence refers to macro-level properties or processes that can be derived through interactions between individual components but in a complicated way such that they are hardly reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996). For the weak emergent patterns or processes, the causes may come from both micro-level and macro-level, therefore, emergent causation may coexist with micro-level causation.
There are many debates on strong emergence, which refers to properties or processes on macro-level that can not be reduced to micro-level properties in principle, including the interactions of individuals O’Connor (1994); Kim (1992). Thus, the collective behaviors of boids are not strong emergent because they are the results of interactions between Boid model. This notion is controversial because it rejects any mechanistic explanations for strong emergent properties, which can not be explained by micro-level variables. Second it raises some long standing philosophical debates about “causal fundamentalism”  Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
Jochen Fromm further explains strong emergence as the causal effect of downward causation Fromm (2005). Consider a system contains three-different scales: micro, meso, and macro. Downward causation refers to the causal power from the macro-level to the messo-level or from the messo-level to the micro-level. Thereafter, although the strong emergent properties or processes on the meso-level are supervenient to micro-level properties and interactions, their causes are from the macro-level and therefore a second order patterns or processes. However, there are many debates on the notion of downward causation itself, e.g.  Yurchenko (2023); Bedau and Humphreys (2008).
Figure 2: Various types of causation that exist within multi-scale complex systems. In this figure, solid arrows represent common causation, which is generally accepted without much controversy. Dotted arrowed lines indicate a form of “causation” driven by supervenience, while dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
In summary, there are three types of causation based on their cross-level properties, which are associated with emergence, as depicted in Figure 2:
1. Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
2. Intra-level causation: This refers to causal effects occurring within the same scale or level.
3. Downward causation: Here, macro-level properties have influences on micro-level properties.
In Figure 2, the black solid lines represent causation without controversy. However, according to a recent paper by Yurchenko  Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality, but rather reasons or explanations. Yurchenko introduces the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers. Instead, they serve as reasons or explanations from the perspective of an observer. While, Yurchenko accepts the intra-level causation in the meso- or macro-level as “real” as the ones in micro-level. He even coined the term “causal equivalent principle” to represent the idea that all the intra-level causation except cross-leveled ones should be accepted and are equivalent each other.
However, other scholars think that the emergent causation, the intra-level causation in the meso- or macro- level, depends on observers Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022). There are much controversy on emergent cross-leveled causation.
3 Quantifying Emergence by Causality
Since causality and emergence have a strong connection, and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence. In this section, we will review several frameworks to quantify multi-scale causality and emergence.
3.1 Early Related Works
Before the theory of causal emergence was proposed by Hoel et al, there were some related works that have introduced very similar ideas as causal emergence theory. For example, Crutchfield et al’s computational mechanics theory considered causal states which are the partitions of the state space and may have good predictions. While, Seth et al proposed G-causality to quantify emergence by using Granger’s causality. We will introduce them in detail.
3.2 Computational Mechanics
The theory of computational mechanics proposed by Crutchfield, Shalizi, and Feldman et al. tries to formulate this kind of emergent causation in a quantative framework Crutchfield and Young (1989). In some sense, computational mechanics can be understood as the inverse of statistical mechanics because what statistical mechanics do is to derive macro-level consequences from micro-dynamics, and the inverse process done by computational mechanics is to construct a minimal causal model from the observations of a stochastic process that can generate the observed time series.
Let’s assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$. We can divide it into two segments beginning from time step $t$: the history before $t$ which is denoted as $\overleftarrow{s_{t}}$ and the future after $t$ which is denoted as $\overrightarrow{s_{t}}$. If the process is stationary, we can drop the denotation of $t$. Thus, all the possible histories $\overleftarrow{s}$ form a set which is denoted as $\overleftarrow{S}$ and all the futures $\overrightarrow{s}$ form a set $\overrightarrow{S}$.
We aim to establish a model that can reconstruct and predict observed random sequences, with a higher accuracy being desirable. However, the randomness of the sequences prevents us from achieving a perfect reconstruction unless we record every randomly occurring character. This would make the model excessively long. To preserve useful information as concisely as possible, we need a coarse-grained mapping that captures the ordered structure in the random sequences, known as patterns Crutchfield (1994). We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets which form a set $\mathcal{R}$. Any subset in $R\in\mathcal{R}$ is a state. We define a function from histories to states, $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$. Thus $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
For a set of states $\mathcal{R}$, we can measure its simplicity using a complexity metric. Intuitively, the larger the cardinality of $\mathcal{R}$, the more complex it is. Additionally, we need to consider its distribution. For example, if one state occurs frequently while others occur rarely, it is less complex compared to a situation with a uniform distribution. Therefore, we can define the statistical complexity $C_{\mu}$ of a set of states using Shannon entropy Shalizi (2001):
$C_{\mu}(\mathcal{R})\triangleq-\sum_{\rho\in\mathcal{R}}P(\mathcal{R}=\rho)\log_{2}P(\mathcal{R}=\rho).$ (1)
When constructing predictive models using a set of states, the statistical complexity refers to the size of the model.
What kind of state set can achieve the best balance between predictiveness and parsimony? We can introduce an equivalent relationship which is called causal equivalence Crutchfield (1994). Concretely, we say $\overleftarrow{s}$ and $\overleftarrow{s}^{\prime}$ are causally equivalent if and only if:
$P(\overrightarrow{s}|\overleftarrow{s})=P(\overrightarrow{s}|\overleftarrow{s}^{\prime})$
This equivalent relationship can partition all the histories into equivalent classes, and they are defined as causal states. We denote all the causal states of a history $\overleftarrow{s}$ as: $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}}$ which is a subset of the histories.
For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that make the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
$T^{(s)}_{ij}\equiv\text{P}(\mathcal{S}^{\prime}=\mathcal{S}_{j},\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s|\mathcal{S}=\mathcal{S}_{i}),$
where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and the notation $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$ in which the first symbol emitted is $s$111we use slightly different symbols compared with reference  Shalizi (2001). Hence, by combining: $\mathcal{S^{\prime}}=\mathcal{S}_{j}$ and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$, we ensure that all subsequent causal states will have the identical initial emitted symbol $s$. We denote the set $\{T^{(s)}_{ij}:s\in\mathcal{A}\}$ by $T$. And the definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$) and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
Up to now we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories. The remaining task is to show that it is in some sense optimal. It has the important properties of being minimally predictive, maximally statistically complex, and minimally stochastic Shalizi (2001).
We shall compare the causal state $\mathcal{S}$ with any state $R\in 2^{\overleftarrow{\mathcal{S}}}$, to show that for the purpose of predicting the future, the causal states does a better job, i.e. gives more information. This property can be formulated as a mathematical theorem which is called the maximal predictability theorem. This is stated as if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
$H[{\overrightarrow{\mathcal{S}}}^{L}|R]\geq H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}].$
Where, $H(\cdot|\cdot)$ is the conditional entropy, $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future. Thus, the inequality shows that the uncertainty of $\overrightarrow{\mathcal{S}}^{L}$ given the causal state $\mathcal{S}$ is less than all other states.
After achieving optimal predictiveness, the causal state set remains the one with minimal statistical complexity. We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$ which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
Next, we present the minimum statistical complexity theorem: for all prescient rivals $\hat{\mathcal{R}}$,
$C_{\mu}(\hat{\mathcal{R}})\geq C_{\mu}(\mathcal{S})$ (2)
We next show that the causal states are minimally stochastic. That is to say, compared with other competitors with the same ability of predicting the future, the causal states and their transition dynamics, have the least uncertainty. Then we have minimal stochasticity theorem which is expressed as: for all prescient rivals $\hat{\mathcal{R}}$,
$H[\hat{\mathcal{R}^{\prime}}|\hat{\mathcal{R}}]\geq H[\mathcal{S}^{\prime}|{\mathcal{S}}].$
Where $\mathcal{S}^{\prime}$ and $\hat{\mathcal{R}}^{\prime}$ are the next causal state of the process and the next state, respectively. This means causal state and $\epsilon$-machine gives the best intrinsic determinism.
Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from observed data? The authors introduced a hierarchical machine reconstruction algorithm in Crutchfield (1994), details are ignored.
Although this algorithm may not be applicable in all operational scenarios, the authors have presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
It is interesting to compare the theory of computational mechanics with causal emergence. Indeed, we can understand all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states. Thus the function $\eta$ that can map a history $\overleftarrow{s}$ to a state $R$ is a possible coarse-graining strategy.
It is worth to be pointed out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can has at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history. Therefore, $\epsilon$ is similar to the notion of effective coarse-graining strategy in reference  Zhang and Liu (2022) (see section 4.1) and the causal transitions $T$ are the corresponding effective macro-level dynamics. And the feature of minimal stochasticity characterize the deterministic property of the macro-dynamic. This is characterized by effective information (EI) in causal emergence theory.
Although the clear definition and quantitative theory of emergence are not given, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003). In Crutchfield (1994), the authors explain that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer. Additionally, they differentiate intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.
3.3 G-emergence Theory
G-emergence theory proposed by Seth in 2008 Seth (2008) is one of the earliest works on quantitative measure of emergence. His basic idea is to use nonlinear granger causality to measure weak emergence in complex systems.
Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$ along, then $B$ is said to be "Granger-cause" of $A$. This implies the existence of Granger causality between $A$ and $B$.
When applying a bivariate autoregressive model for predictions, residual terms are included in the equations of the two variables. Then, the residuals can be utilized to quantify the extent of the causal effect in G-causality. The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals, one is residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other one is the residual of the full prediction model, which shown in Figure 3. Besides, in Seth (2008), the author also defined G-autonomous as if the past values of a time series can help to predict its own future values. And the degree of G-autonomous can be measured in a similar way as G-causality.
Figure 3: G-causality and G-autonomous
With these two basic notions in G-causality, the author defines G-emergence on macro-variables in Figure 4: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if: (i) $M$ is G-autonomous with respect to $\{m\}$ and (ii) $\{m\}$ is G-cause of $M$. The degree of G-emergence can also be quantitatively measured by multiplying the degree of G-autonomous of $M$ and the average of G-causes of $\{m\}$ on $M$.
Figure 4: G-autonomous
The author tests G-emergence theory on the Boid model. This model is a famous artificial life model to simulate the flocking behavior of birds by using three simple rules: cohesion, alignment, and separation Reynolds (1987). All these basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled. The author found that with the strength of cohesion increased, the G-emergence of the entire flock is also increased. Here, the macro-variable is selected as the center of the flock (center-of-mass), and each bird is a micro-variable. The author also discovered a downward causality phenomenon in this simple model that is the center of mass can be used to predict each individual bird. However, the author does not distinguish downward causation from other common causality in this work because Granger’s causality is not a real causal relationship.
Seth’s G-emergence theory is the first attempt to quantify emergence phenomenon via causality measure. However, the causality measure that the author used is granger causality which is not a strict causal measure and it is also depends on the regression method to be used. Furthermore, the measure is defined on variables but not dynamics which means the result depends on the selection of variables.
3.4 Other Quantitative Theories of Emergence
There are alternative means of measuring emergence that do not require relying on causality. Two different ways are discussed by the scholars, one is to understand emergence as a process from disorder to order, another one is to understand emergence from the view of "the whole is greater than the sum of parts".
For example, Moez Mnif and Christian Müller-Schloer used Shannon’s entropy  Mnif and Müller-Schloer (2011) to measure order and disorder. In a self-organized process, emergence occurs when there is an increase in order. This increase can be quantified by measuring the Shannon entropy difference between the initial state and the final state, denoted as $H_{start}-H_{end}$. However, this definition has two limitations. Firstly, the measurement of entropy, denoted as $H$, is dependent on the abstract level of observation. Therefore, it is necessary to account for the entropy increase resulting from a change in the observer’s abstract level. Secondly, the choice of the initial condition of the system is arbitrary. To address this limitation, one approach is to measure the relative level of Shannon entropy compared to the maximum entropy distribution. Finally, emergence can be quantified by:
$Emergence=H_{max}-H-\Delta H_{view},$ (3)
where, $H_{max}$ is the maximized entropy of the system. If no prior information is available, the maximum entropy, denoted as $H_{max}$, corresponds to the Shannon entropy of an equal probability distribution. On the other hand, $H$ represents the Shannon entropy of the system at the final moment of a self-organization process. Additionally, $\Delta H_{view}$ represents the entropy increase during this process resulting from a change in the observer’s abstraction level. It should be noted that if the observer does not alter their abstraction level, this term would be zero. We can also normalize this quantity by dividing $H_{max}$ such that different features can be compared by the normalized quality. For a multivariate system, the authors suggest employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables. Then the authors apply this method to a simulated system of chickens. M. Tang and X. Mao applied this indicator on artificial society models  Tang and Mao (2014).
Inspired by Moez Mnif and Christian Müller-Schloer’s work,  Fisch et al. (2010) suggests to use the divergence measure between two probability distributions to better quantify emergence. They understand emergence as an unexpected or unpredictable change of the distribution underlying the observed samples. However, this method has the defect of computational complexity and estimation accuracy. To conquer the problems,  Fisch et al. (2012) further proposed an approximating method by using Gaussian mixture model to estimate the density and introducing Mahalanobis distance to characterize the divergence between data and Gausian component, better results can be obtained. In Fisch et al. (2011), the authors systematically compare the three mentioned methods, and apply them on a simple test example. Another Shannon entropy based emergence measure is given by Holzer and de Meer et al. Holzer et al. (2008); Holzer and de Meer (2011). They consider a complex system as a self-organization process, in which different individuals interact with communications. Then emergence can be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
Different from all the mentioned methods above, Teo et al. (2013); Szabo and Teo (2015) proposes a method to quantify emergence based on the idea of "the whole is greater than the sum of its parts", and it defines emergence from the interaction rules and states of agents instead of on the overall statistical measure of the whole system. Specifically, this measure consists of two terms that are subtracted from each other. The first term characterizes the collective states of the entire system, whereas the second term represents the summation of the individual states of all its constituent parts. This measure emphasize the emergence that arises from the interactions and collective behavior of the system. Then this method is tested on the example of bird flocks simulation.
3.5 Erik Hoel’s Causal Emergence Theory
3.5.1 Basic idea
The first quantitative emergence theory defined on Markov dynamics and causality measure based on intervention is Erik Hoel’s causal emergence theory.
In this framework, system properties can be characterized at various levels, ranging from micro to macro. If a system exhibits stronger causality at the macro level than at the micro level, it demonstrates causal emergence. Causality is reflected between successive states during the system’s evolution. The strength of a system’s causality reveals the extent to which its future state is influenced by its current state. The basic idea of causal emergence can be illustrated by Figure 5.
Figure 5: The basic idea of Erik Hoel’s causal emergence theory.
For example, statistical mechanics is a typical theory of causal emergence. On the micro-level, a huge number of molecules collide each other and exhibit random behaviors such that probabilistic language must be used to describe them. However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure and temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors. Therefore, thermodynamic laws have stronger causal effect than micro-level molecular dynamics.
Formal tools are used to describe the elements in Hoel’s causal emergence theory. Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro systems, and the inherent strength of causal effect of the Markov model can be measured with effective information (EI) which indicates how effectively a particular state influences the future state of a system.
EI is an intrinsic property of a system’s dynamics which can be quantified by the transitional probability matrix (TPM). A coarse-graining strategy is a function that can map a set of micro-states into a particular macro-state, and therefore a new dynamical model which can be described by TPM can naturally be derived from the micro-level TPM, and the effective information of the coarse-grained model can also be computed. The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase, and at a certain macroscopic scale of coarse-graining, the effective information reaches a maximum: this is the scale at which the system state has maximum causal power to predict future states in the most reliable and effective way.
To illustrate the causal emergence, let’s look at a particular Markov chain model with 4 possible states, whose transition probability matrix (TPM) is:
$S_{m}=\begin{bmatrix}1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 0&0&0&1\end{bmatrix}$ (4)
In this model, if we take the first 3 states as a group 1 and the last state as group 2, anyone in group 1 can transit to all 3 states of the same group with equal probability at the next moment, while because there’s only 1 state in group 2, so the 4th state will always stay in its position. Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty is mainly from group 1.
However, if we merge the first 3 states in group 1 into 1 new state, and keep the 4th state in group 2 as it is, then we get a coarse-grained Markov model with 2 states, with TPM:
[…]
Now the future states of the new system can be fully determined by the current states. It shows that we can eliminate the uncertainty of an undeterministic system by doing coarse-graining over the system states.
3.5.2 About Coarse-graining
The example above illustrates how a system’s determinism can increase as it is coarse-grained. Coarse-graining is a process that simplifies the description of a system by grouping its components into larger, more slowly varying units. It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of microscale interactions. Unlike dimension reduction techniques, such as PCA and SVD, coarse-graining takes into account the system’s features at different spatial and temporal scales. Coarse-graining also differs from renormalization. In physics, the method of renormalization group is invented to avoid the infinity in integrals, it is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro- levels Christensen and Moloney (2005); McComb (2003). However, coarse-graining methohd does not have this requirement in general. Although both techniques are targeting to describe a system from a coarse-grained level, however, coarse-graining are more focused on the states of the system, while renormalization cares of the dynamics, the system rules or partition function etc, which is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking. These are aspects that coarse-graining typically does not consider.
Given a transitional probability matrix, there are a large amount of ways to partition the state space in order to coarse-grain the system. What kind of coarse-graining strategies are more reasonable? In Hoel et al. (2013) and Hoel (2017), the authors suggest to select an appropriate coarse-graining method by maximization effective information of the coarsed TPM. In history, there is another criterion to select coarse-graining strategy called lumpability. Where lumpability of a TPM refers to the coarse-grained TPM exhibits similar dynamics and observation statistics Kemeny and Snell (1983). Commutativity is another requirement which needs the operations of coarse-graining and state transition being commutative Pfante et al. (2014). There are also other discussions or reasonable model reduction of Markov or hidden Markov models Kotsalis et al. (2008); White et al. (2000); Wolpert et al. (2014)
3.5.3 About Effective Information
In order to quantify emergence, a widely used method is to calculate the mutual information between the past states and future states of a system. This mutual information defines the upper limit of the information being transferred from past to future. When the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
$I(X_{t};X_{t+1)}=\sum P(X_{t},X_{t+1})\log\frac{P(X_{t},X_{t+1})}{P(X_{t})P(X_{t+1})}$ (6)
One limitation of utilizing mutual information is that the resulting value may fluctuate in response to changes in the joint probability of data. This can make it difficult to obtain a consistent outcome when the system is exposed to different inputs.
To address this limitation, Hoel introduced the concept of effective information. This concept measures the causal effect of the current state of a system on its future state, also based on mutual information. Effective information is a scoring metric that is calculated from the system’s transition probability matrix (TPM). The TPM does not change solely based on input data.
Assuming that the current state of the system is in the maximum entropy distribution $I_{D}$ (uniform distribution), the TPM predicts that the state distribution of the system at the next moment (i.e., in the future) will be $E_{D}$. The effective information $EI$ of the system is then defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
$EI=I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{X_{t}\sim U}$ (7)
$X_{t}\sim U$ represents the probability of the random variable $X_{t}$ following a uniform distribution. It can be readily demonstrated that the effective information ($EI$) can be fully described by the transition probability matrix ($TPM$) (see Appendix 7).
[…]
Effective information offers a solution to the aforementioned limitation of mutual information. Given that TPM can capture the inherent nature of a system, effective information is likewise an intrinsic property of the system.
Causal emergence involves assuming a uniform distribution for the input data, which can be seen as a way of manipulating the input data to isolate the interference of the original distribution of input data. This manipulation is formalized using Judea Pearl’s causation theory, specifically through the "do" operation. By imposing this assumption of a uniform distribution, we create a controlled environment with equal probabilities of $X_{t}$’s states and allows us to better understand the causal relationships between $X_{t}$ and $X_{t+1}$. Therefore, the term "causal emergence" is used.
We can calculate EIs for both TPMs of macro- and micro- dynamics, and their difference is defined as the measure of causal emergence, that is:
$CE=EI(TPM_{M})-EI(TPM_{m}),$ (9)
where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the one of micro-level. $CE$ measures the degree of causal emergence. If $CE>0$, then causal emergence occurs, otherwise, does not occur.
However, there is a limitation for EI and CE when comparing two dynamics that differ significantly in size. This is because the value of effective information relies on the number of possible states within the system, denoted as $N$, with an upper bound of $\log_{2}(N)$. To facilitate the comparisons between different coarse-graining strategies and scales, effective information is often normalized resulting in a metric known as the effect coefficient $Eff$.
$Eff=\frac{EI}{\log_{2}(N)}$ (10)
The value of effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states. If the information is fully transferred, the effect coefficient is $1$.
In addition to characterizing causal emergence, effect coefficient can be further broken down into two meaningful components: "Determinism," which represents the certainty of a current state evolving into a certain state, or diverging into multiple states in the next moment, and "Degeneracy," which represents the possibility of multiple current states converging into one state in the next moment.
$Eff=Determinism-Degeneracy.$ (11)
It should be noted that the determinism of a system is always greater than its degeneracy, as the lower bound of the effect coefficient is 00. The following examples illustrate what determinism and degeneracy look like in systems with varying TPMs.
Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graning can change them. (a) Deterministic and zero degenerative system; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from c or d, resulting in a fully deterministic system with high effect coefficient.
In Figure 6, the square cells represent for the elements of TPM matrix, and the grey scale represents for the value of TPM elements. Example (a) is a bijective system, meaning that all information from current states can be transferred to future states without loss. It is fully deterministic with zero degeneracy. Example (b) is an extreme case where all current states lead to only one future state, illustrating that high determinism does not necessarily imply a high effect coefficient. Systems (c) and (d) differ but have the same effect coefficients. Finally, system (e) is a coarse-grained version of either system (c) or (d), demonstrating two important points: first, different microscopic systems can be coarse-grained to the same macroscopic system, and second, causal emergence can be quantitatively captured from coarse-graining by the increased effect coefficient.
Despite the fact that effective information can identify and quantify causal emergence, as previously mentioned, there are other independently developed measures of causation, such as Humean constant conjunction Hume (2016), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009) etc. Hoel found all these measures, including effective information can be related to a small set of "causal primatives" Comolatti and Hoel (2022): the probability representations of sufficiency and necessity (and its variant) of cause $c$ and effect $e$.
$suff(e,c)=P(e|c)$ (12)
$nec^{\dagger}(e,c)=P(e|C)$ (13)
$-\𝑐nec(e,c)=1-P(e|C\verb|\|c)$ (14)
Where $C$ stands for the set of all causes and $-\𝑐C\verb|\|c$ stands for the set of all causes but $c$. Hoel pointed out that determinism is actually a measure of sufficiency, while degeneracy is an inverse measure of necessity. The consistency in causal primitives indicates there is a scientific foundation of these individually developed measures of causation.
One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation 8. Therefore, Varley (2022) proposed the conception of flickering emergence to decompose EI by each term being added in Equation 8. This new feature can characterize the local properties of a Markov dynamics.
3.5.4 Examples of Causal Emergence
The example above demonstrates a coarse-graining strategy that is intuitive, as it involves aggregation at the state level. However, in reality, coarse-graining often occurs among variables, and the combination of these variables produces different states. Let’s look at a more complicated example of a boolean network that given in Hoel et al. (2013) and we will inspect the causal emergence when coarse-graining its state transition mechanism. A, B, C and D are 4 binary variables whose state transational relations are as shown in Figure  7(a).
Figure 7: The mechanism and transition probability matrix of a boolean network. Each variable’s next state is determined by the combination of the current states of two other variables, as shown in the boolean network (a). If we examine an atomic subsystem with two inputs and one output, such as AB -> C, the next state of C depends on the present states of A and B. If A=0 and B=0, then the probability of C being 1 is 70%, while the probability of it being 0 is 30%, as indicated in table (b). This same microscopic dynamics applies to other atomic subsystems such as AB -> D, CD -> A, and CD -> B. When we expand to a subsystem of two inputs and two outputs, AB -> CD and CD -> AB, we obtain the transition probability matrix (TPM) demonstrated in table (c). Finally, we extend to the dynamics of the entire system ABCD -> ABCD, whose TPM is represented in a heatmap (d). The effective information and relative metrics of the entire system are as below: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
[…]
Figure 8: An example to show how different coarse-graining strategies on the same boolean network shown in Figure 7 can lead to different boolean dynamics. (a) Coarse-graining strategy 1. (b) Macro-dynamics leading by strategy 1. (c) The corresponding heatmap of TPM and effective information metrics of the macroscopic system by method 1. (d) Coarse-graining strategy 2. (e) and (f) are the macro-dynamics and TPM correspond to the coarse-graining strategy 2 in (d) respectively.
The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.05$, respectively. It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), while it has decreased in the second case ($EI_{ma2}(0.05)<EI_{mi}(1.15)$). Thus, causal emergence is observed in the first coarse-graining instance. This implies that 1) proper coarse-graining can lead to causal emergence in a macroscopic system, and 2) not all coarse-graining approaches can result in higher effective information. This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well defined quantification method. This gives us the opportunity to identify causal emergence in a given system. There is a question coming next: what coarse-graining strategy can generate the maximum effective information? We will address this question in section 4.
3.6 Extensions and Applications
We will discuss the extensions and applications of Hoel’s theory of causal emergence in this sub-section.
3.6.1 Extensions on Continuous Systems
Effective information and its associated metrics have been shown to be useful in quantifying causal emergence. However, there are some limitations to be aware of. First, the method is only applicable to discrete-state systems. Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate effective information metrics. Efforts have been made to apply or extend the concepts of effective information to continuous systems, such as Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
Ordinal Partition Network
One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones by ordinal partition network (OPN) method Varley (2020). The basic idea is to transform the time series data generated by a continuous dynamical system into sub-time series of dimensionality $d$ with discrete values. And then to establish a Transition Probability Matrix (TPM) between these discrete series. This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series. By employing rank orders, this method ensures that the values are restricted to integers within a certain range.
In Varley (2020), the author applied OPN method on Rossler chaotic attractor system. By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) is sensitive to the critical phase transition between period and chaos Varley (2020).
Causal Geometry
Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the conception of causal geometry Chvykov and Hoel (2020).
To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is:
$y=f(x)+\varepsilon,$ (15)
where $x$ is the input variable defined on the interval $[-L/2,L/2]$ and $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a gaussion noise with zero mean and standard deviation $\epsilon\ll L$. Then, an approximate form of the effective information for such continuous system can be obtained by Gaussian integral.
$EI\approx\ln(\frac{L}{\sqrt{2\pi e}})+\frac{1}{2L}\int_{-L/2}^{L/2}\ln\left(\frac{f^{\prime}(x)}{\epsilon}\right)^{2}dx.$ (16)
The reason why we set the domain of $x$ on $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented, and also facilitate to define a uniform distribution on it.
This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ as $x(t)$ and $x(t+1)$ respectively. The formula also can be extended to higher dimension. Suppose $\mathbf{x}\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers, Equation 16 can be generalized to the following form:
$EI\approx\ln(\frac{L^{n}}{(2\pi e)^{m/2}})+\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2},$ (17)
where, $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, and ||$\cdot$|| ⋅ | is the operation to take absolute value, and $\det$ is to take the determinant.
Notice that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$, thus the term in the expectation in Equation 17 can be written as:
$\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2}=\left|\det\left(\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right]\right)\right|,$ (18)
and this is the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
$g_{\mu\nu}\equiv-\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right],$ (19)
it measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ on $\mathbf{x}$. Where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$. Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, and it is the collections of all possible distributions of $p(\mathbf{y}|\mathbf{x})$. This is where the term of “geometry” in this framework comes from.
Finally, we can obtain an expression for EI with Fisher information metric:
$EI\approx\ln\frac{L^{n}}{(2\pi e)^{m/2}}-\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln|\det(g_{\mu\nu})|.$ (20)
This formula may be generalized to the cases where $p(\mathbf{y}|\mathbf{x})$ is non-Gaussian distribution. Once the distribution function is known, we can obtain its Fisher information metric, then EI can be calculated via Equation 20. The reason behind is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
While, in Chvykov and Hoel (2020), the authors consider a more general case, in which the intervention noise is added on the input(intervention) variable $x$ in Equation 16. The intervention noise is $\xi\sim\mathcal{N}(0,\delta^{2})$, and $\delta$ is the standard deviation of $\xi$. While, $\varepsilon$ is called observational noise to differentiate with $\xi$, and it is added on the output(observational) variable $y$ according to Equation 16. Finally, if two kinds of noises are both considered, and if $L=1$ and $\epsilon\ll 1$, Equation 16 becomes:
[…]
This is the formula of EI for continuous mapping function (Equation 15) given by Chvykov and Hoel (2020). Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation 15) and a more discrete map with an adjustable parameter in the logistic function. They discovered that when the noise level is low, the continuous map may exhibit higher EI than the discrete map. However, as the noise level increases, discretizing the mapping function can lead to a model with higher EI. This phenomenon helps to explain why digital circuits eventually outperform analog circuits in mitigating noise interference: the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we can not control $\mathbf{y}$ by intervening $\mathbf{x}$ directly. Instead, we can intervene $\mathbf{x}$ to influence $\theta$, and to affect $\mathbf{y}$ indirectly. Therefore, the three variables form a Markov chain: $\mathbf{x}\rightarrow\theta\rightarrow\mathbf{y}$.
[…]
Finally, the EI calculation formula for the causal geometry is:
$EI_{g}=\ln\frac{V_{I}}{(2\pi e)^{n/2}}-\frac{1}{2V_{I}}\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}\ln\left|\det\left(I_{n}+\frac{h_{\mu\nu}}{g_{\mu\nu}}\right)\right|d^{l}\theta,$ (22)
in which, we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
EI calculation for neural networks
One of the application areas for continuous EI is on neural networks. In Marrow et al. (2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compare the results with information bottleneck theory. The method to calculate EI that they used is to convert continuous variables into discrete ones by dividing the domains of input and output variables into small regions. However, this method has high computational complexity, and is hard to extend to high dimensions.
In Zhang and Liu (2022), the authors propose a new numeric method for calculating EI on neural networks. The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping following:
$\mathbf{y}=f(\mathbf{x})+\varepsilon,$ (23)
where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, and $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network, $\varepsilon$ is a Gaussian noise with zero means and the covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ on the whole training data set. In this way, the neural network can be regarded as a Gaussian distribution $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
Therefore, Equation 17 can be used on this neural network to calculate EI. In Zhang and Liu (2022), the authors use Monte Carlo integration method to estimate the expectation in Equation 17, this technique reduces the computational complexity in a large degree.
However, when the authors compared the results calculated by Equation 17 for different neural networks with different dimensions, they obtained unreasonable results: EI increases with the number of dimensions, thus, causal emergence is always not be observed because micro-dynamics has unreasonably larger EI. Another defect of Equation 17 is that the parameter $L$ always dominates the value of the expression which should be removed from the equation. One way to solve this issue is to use Eff instead of EI (see Equation 10), however, not all $L$s in the Equation 17 can be eliminated. Therefore, the authors invented a notion of dimension averaged EI,
$dEI\equiv=\frac{EI}{m}$ (24)
which simply divide EI by the dimension of the output variable $\mathbf{y}$. By using dimension averaged EI, $L$ can be also removed if we calculate dimensional averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ for macro- and micro- dynamics respectively, that is:
$dCE(f_{M},f_{m})\equiv\frac{EI(f_{M})}{n_{M}}-\frac{EI(f_{m})}{n_{m}},$ (25)
where, dCE is the dimension averaged causal emergence measure. If we bring Equation 17 into Equation 25, we get:
$dCE\approx\frac{\mathbb{E}_{\mathbf{x}_{M}}\log|\det(\partial_{\mathbf{x}_{M}}f_{M})|+\sum_{i=1}^{x_{M}}\log\sigma_{i,M}^{2}}{n_{M}}-\frac{\mathbb{E}_{\mathbf{x}_{m}}\log|\det(\partial_{\mathbf{x}_{m}}f_{m})|+\sum_{i=1}^{x_{m}}\log\sigma_{i,m}^{2}}{n_{m}},$ (26)
in which, $L$ is eliminated.
3.6.2 Causal emergence in complex network
Many complex systems can be represented by networks, therefore, to apply the framework of causal emergence on networks is necessary. However, two problems should be solved in advance to apply Hoel’s framework, one is to assign a dynamics on the studied network because networks are static structures without any dynamical properties, another one is to coarse-grain the network.
Klein and Hoel Klein and Hoel (2020) solved the first problem by considering random walks on complex networks. Subsequently, the TPM of the system is defined based on the transfer probabilities between nodes of a large number of random walkers. Due to the good mathematical properties of the random walk dynamics on graphs, the authors can establish an explicit expression of EI on networks, and the final form only relies on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network. The expression is as follow:
$EI=H(\langle W^{out}_{i}\rangle)-\langle H(W^{out}_{i})\rangle.$ (27)
Where, $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,...,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes which characterizes the degeneracy of the dynamics. And $H(W^{out}_{i})$ is the uncertainty of node $i$.
To tackle the second problem, the greedy algorithm is used to group nodes to form a macro network in Klein and Hoel (2020). It is worth noting that when merging micro nodes into macro nodes, the TPM of the corresponding macro network can be derived by merging the probabilities in TPM of the micro-level. To ensure that the grouped macro network maintains the same random walk dynamics as the original one, dynamic consistency test is implemented.
In the experimental section, the paper explores the application of greedy search for identifying causal emergence in complex networks. The experiments are conducted on artificial networks, as well as four types of real networks. In the case of ER random networks, the size of effective information is solely dependent on the connection probability $p$, and converges to $-\log_{2}p$ with the increase of network size. Additionally, a significant finding reveals the presence of a phase transition point, where the average degree of the network reaches approximately $\log_{2}N$. Beyond this point, the random network structure does not contain additional information with increasing size. In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increases as the network size expands. Conversely, when $\alpha>1.0$, the opposite conclusion is reached. The scale-free network corresponding to $\alpha=1.0$ represents the critical boundary of growth. Regarding real networks, the authors find that biological networks exhibit the lowest EI due to the presence of significant noise, which can be removed by effective coarse-grained, and causal emergence is the most significant compared with other types of networks. However, technical networks exhibit sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships. Consequently, they possess the highest EI among the networks studied.
The network coarse-graining method mentioned above is the greedy algorithm. However, when the network is very large, the efficiency of this method remains considerably low. Following that, Griebenow et al. Griebenow et al. (2019) introduced a spectral clustering-based approach to identify causal emergence within networks. More specifically, the method involved performing eigenvalue decomposition of the TPM, followed by constructing a similarity matrix using the eigenvectors of the nodes. The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro node. Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyper-parameter $\epsilon$. In this paper, the author additionally proposed a gradient descent algorithm based on deep learning. This approach encompasses several steps, beginning with the random initialization of a grouping matrix. Then this matrix is used to construct a macro network by combining the micro network. Finally, the grouping strategies are automatically learned through the process of maximizing effective information within the macro network. However, this method often falls into the local optimal solution.
3.6.3 Other Applications
Once the method for quantifying causal emergence in complex systems is developed, it can be applied across various fields that possess abundant network data. The first type of networks which was studied is biological networks.
As previously discussed, biological networks are full of noise, which poses challenges in comprehending their internal operating principles. On the one hand, such noise arises from the inherent fluctuations within the system itself, while on the other hand, it can be introduced through measurement or observation processes. Consequently, Klein et al. Klein et al. (2022) further explored the relationship among noise, degeneracy and certainty in biological networks and their specific meanings. For instance, in gene expression networks, highly deterministic relationships indicate that the expression of one gene almost invariably leads to the expression of another gene. Simultaneously, degeneracy is a prevalent phenomenon in the evolutionary processes of biological systems. Due to these two factors, it remains unclear at which scale biological systems should be analyzed to gain a deeper understanding of their functions.
To address this, Klein et al. Klein et al. (2021) conducted an analysis of protein interaction networks across more than 1,800 species. They employed EI as a measure to assess the levels of noise and uncertainty in protein interactions. The findings revealed that the macro-scale network exhibited lower levels of noise and degeneracy. Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to nodes that were not part of the macro-scale interaction group. Through robust analysis, the authors demonstrate that eukaryotes exhibit a stronger degree of causal emergence compared to archaea. Additionally, to address the ’deterministic paradox,’ the authors introduce the concepts of the neutral process and the selection process in biological evolution. The neutral process operates at the micro scale, leveraging mutations to promote interactions and enhance species diversity. On the other hand, the selection process operates at the macro scale, effectively eliminating noise that hampers system operation and efficiency. Therefore, in order to adapt to the demands of evolution, it becomes essential for evolved biological systems to function across multiple scales.
Hoel et al. Hoel and Levin (2020) conducted further research on the causal emergence within biological systems. The authors elaborated that macro and micro systems exist widely in biological systems, for example, the micro scale of a group of cells can be the potential ion channel changes, and the macro scale corresponds to the membrane potential changes of cells. Furthermore, the authors utilized EI to analyze gene regulatory networks, aiming to identify the most informative models for controlling mammalian heart development. By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it is revealed that informative macro-scale structures are prevalent across biological systems. Additionally, the authors emphasize the importance of evolutionary systems operating at multiple scales, as it offers significant advantages. Natural selection requires variation between populations, and degradation observed in biological systems serves as a crucial factor for evolution. Degradation provides the conditions for the evolutionary process. However, organisms also need to maintain predicted consistency in phenotype, behavior, and structure to ensure survival and reproduction. Consequently, evolved systems need to operate at multiple scales, and the function of multi-scale systems is malleable in changing environments.
Swain et al. Swain et al. (2022) conducted an investigation into the impact of ant colony interaction history on task assignment and task switching. They employed effective information to examine how noise information propagates among ants, while also exploring the relationship between EI and the proportion of ants assigned to different tasks within ant colonies. The study revealed that the extent of historical interaction between ant colonies influenced task assignment. Additionally, the specific type of ant colony involved in an interaction determined the level of noise present within that interaction. For example, the interactions between foragers displayed significantly higher levels of noise when contrasted with interactions between nurses or cleaners. Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also play different roles in maintaining group cohesion.
EI indicator and causal emergence theoretic framework can also be applied to artificial systems. For example, Marrow et al. Marrow et al. (2020) quantify and monitor the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer. By observing the changes of EI, including determinism(sensitivity) and degeneracy, throughout model training, the generalization ability of the model can be determined, thus helping us to better understand and explain the working principle of neural networks.
Varley et al. Varley (2020) attempted to apply the causal emergence framework to both discrete cellular automata and continuous Rossler systems. In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic and complex. By considering each state as a node, where each state determines the subsequent state, a directed state transition graph is constructed. The analysis revealed that rules 1, 2, and 4 correspond to the strongest causal emergence of dynamics. Notably, the network constructed by these three rules exhibited a significant presence of star and spoke motifs. In addition, the authors also draw some quantitative conclusions, for example, among the 17 rules belonging to the third and fourth categories, 30% exist causal emergence, 70% show causal degradation, and the CE of cellular automata with the same rules remained relatively consistent across different sizes.
Furthermore, the author employed the OPN algorithm to transform the continuous system into a discrete state transition graph for comparative analysis (see section 3.6.1). They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
3.7 Critiques
In the extensive literature on causality and emergence, Hoel’s theory has gained attention for its association of emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner. However, Dewhurst  Dewhurst (2021) provides some philosophical clarity of Hoel’s theory, arguing that it is epistemological rather than metaphysical. It suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving "genuinely novel causal powers." This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
The computation of Hoel’s effective information relies on two premises: 1) knowledge of the system’s microscopic dynamics, and 2) knowledge of the coarse-graining scheme. However, in practice, it is rare for both conditions to be simultaneously satisfied, especially in observational studies where both are unknown. This limitation hinders the practical applicability of Hoel’s theory.
It is pointed out that Hoel’s theory negleted the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022). Additionally, the combination of some coarse-graining operations over states and coarse-graining operations over time does not exhibit commutativity. For instance, assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold. This indicates that certain coarse-graining operations can result in a discrepancy between the evolution of macroscopic states and the coarse-grained states of the evolved microscopic systems. It implies the need for consistent constraints on coarse-graining strategies.
That means solely maximizing EI may raise some problems, and further constraints must be added in the framework. We will discuss this problem in the section of 4.1.2.
3.8 Rosas Fernando’s Quantifying Causal Emergence
3.8.1 Basic idea
In Hoel’s framework, it is essential to find a coarse-graining strategy in order to determine the occurrence of causal emergence, and the outcome is influenced by the choice of the coarse-graining method. While it has been suggested by Hoel Hoel et al. (2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised as proposed by Dewhurst Dewhurst (2021). Third, EI is an global measure because it requires to intervene the input variable $X$ to be a uniform distribution on the whole domain Rosas et al. (2020). However, many regions are not be able to be observed from data. Consequently, there is a pressing need for an alternative theoretical framework of causal emergence that does not rely on a coarse-graining method.
In response, Fernando E. Rosas took an approach that did not require a coarse-graining strategy as prerequisite, attempting to break down excess entropy - the mutual information between a system’s past and future states - into non-overlapping parts to identify the information components most relevant to causal emergence. To accomplish this, he relied on the Partial Information Decomposition (PID) framework proposed by Williams and Beer, which provides a method for non-overlapping decomposition of joint mutual information Williams and Beer (2010). First, we will introduce Williams and Beer’s theoretic framework.
3.8.2 Partial Information Decomposition
The Partial Information Decomposition (PID) framework investigates the general informational relationship between source variables and a target variable. To make the problem description simpler without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
Figure 9: Venn diagram of PID. (a) Relationship of mutual information between Y and X1, X2. (b-d) Examples that joint mutual information is solely contributed by redundant, synergistic, and unique information. (e) Redundancy lattice representation of (a).
The mutual information between the target variable and individual source variables,$I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibit a complex relationship and can not be intuitively converted to one another. Nonetheless, it’s an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
$I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)+Un(X_{1};Y|X_{2})+Un(X_{2};Y|X_{1})+Syn(X_{1},X_{2};Y)$ (28)
Specifically:
  • $Red$: Redundant Information refers to the information held by both sources;
  • $Un$: Unique Information refers to the information held by one source but not the other;
  • $Syn$: Synergistic Information is the information held by all sources together, but not any individual one.
If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates that they must be independent of one another.
For a more intuitive understanding, let’s consider a few simple toy examples.
Case 1, $Y=X_{1}=X_{2}$: This is a scenario in which a single source variable can predict the target variable, and the addition of another source variable does not enhance the prediction of the target variable. In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, Fig 9(b).
Case 2, $Y=X_{1}\oplus X_{2},X_{1}\perp{X_{2}}$: In another scenario, neither of the source variables can predict the target variable individually, but together, they can predict the target variable synergistically. In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, Fig 9(c).
Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, Fig 9(d).
For general cases, Williams and Beer presented a method for calculating redundant information defined as $Red_{min}$. This method reflects the concept of redundancy, identifying it as the shared information across all sources, is equivalent to the minimum amount of information contributed by any single source:
$Red_{min}(X_{1},X_{2};Y)\sum_{y\in\mathcal{Y}}P(Y=y)\min_{Xi}I(Y=y;X_{i})$ (29)
Fig 9(e) presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010). In this representation, {12} denotes synergistic information, $\left\{1\right\}$ and $\left\{2\right\}$ denote unique information, and $\left\{1\right\}$$\left\{2\right\}$ denotes redundant information. In the subsequent sections, we will also see that the $\phi$ID framework employs the notation of the redundancy lattice.
Although the PID framework remains compatible with scenarios involving more than two source variables, it should be noted that the corresponding Venn diagrams and redundancy lattices for these scenarios can become substantially more complicated and difficult to decipher, as discussed in Williams and Beer (2010).
3.8.3 Integrated Information Decomposition
The PID framework provides a useful framework for analyzing the non-overlapping information composition in a multivariate system. However, its application to causal analysis of dynamic systems is limited by the fact that it only allows for a single target variable. This limitation prevents the framework from fully capturing the transitions of multiple states across time steps. To address this challenge, Fernando E. Rosas etc have developed Integrated Information Decomposition ($\phi$ID) Mediano et al. (2019), which takes its name from Integrated Information Theory (IIT) Oizumi et al. (2014). This extension of PID provides a more comprehensive means for analyzing dynamic systems.
To introduce Rosas’s framework clearly, we only consider a system with two variables. All the definitions and calculations can be generalized to the systems with more variables.
The objective of the $\phi$ID framework is to decompose excess entropy into non-overlapping information components. In a two-variable Markovian system, the excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states. From the causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, while $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent effects.
Rosas first noted that mathematically, the PID framework is symmetric with respect to source and target variables. There are two viewpoints for analyzing the above Markovian system with PID. One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the "cause" to elements), while the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the "effect" to elements). The redundancy lattices of these perspectives are illustrated in Fig 10 (a) and (b), and respectively, are referred to as "forward PID" and " PID".
Figure 10: (a) Redundancy lattice of "forward PID", and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of "backward PID" and its target variable $X^{1}_{t}X^{2}_{t}$.
Next, Rosas introduced the $\phi$ID framework, which consolidates the forward PID and backward PID into a single framework. In this framework, the one-to-many relationship in PID is expanded to include many-to-many relationships. He built full connections between the elements of redundancy lattices of the forward PID and the backward PID. This approach generates 16 relations between the source and target, as depicted by the colored lines in Fig 11 (a). (Each color corresponds to a specific source element of $X_{t}$.) These relations are referred to as "$\phi$ID atoms" and can be represented as vertices in a lattice, as shown in Fig 11 (b). This more complex lattice is referred to as the double-redundancy lattice because it is a "product" of two redundancy lattices (of forward PID and backward PID).
Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID. (b) Double-redundancy lattice, whose vertices correspond to the edges in (a)
In the double-redundancy lattice, each vertex is defined as a "$\phi$ID atom," denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, let $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is demoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$. $\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system. Consequently, the formal definition of Integrated Information Decomposition ($\phi$ID) is:
$\mathbb{E}=I(\mathbb{X_{t}};\mathbb{X_{t+1}})=\sum_{\alpha,\beta\in\mathcal{A}}{I^{\alpha\rightarrow\beta}_{\partial}}$ (30)
The main advantage of $\phi$ID is that it can provide more fine-grained analysis of complex systems than PID by introducing the temporal dynamic. By decomposing the mutual information of the system at different times, $\phi$ID provides another approach for the quantitative study of Causal Emergence.
3.8.4 Reconciling Emergences
Based on the concept of synergistic information in the PID framework, Rosas proposed a quantitative definition of causal emergence and further classified it using the $\phi$ID framework, which addresses the problem of finding a coarse-graining strategy. The definition includes two aspects: the first is whether the system has the ability to produce causal emergence objectively. Then, on this basis, whether there is causal emergence happening with specific macroscopic features.
In terms of the capacity to generate causal emergence for a system, this definition links causal emergence to synergistic relationships among variables in different times. Therefore, a system $X_{t}$ is said to have capacity for causally emergent features if and only if that:
$Syn(X_{t};X_{t+1})>0$ (31)
This means that causal emergence is regarded as the synergy effect among variables between the previous and subsequent moments in the system with Markovian dynamics.
Then, Rosas further classify causal emergence by the $\phi$ID framework. Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ by the $\phi$ID, here are four information atoms corresponding to synergistic effect, which is regarded as causal emergence, are $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$. Since $\phi$ID includes forward and backward PID, he further divides causal emergence into downward causality and causal decoupling according to the different characteristics of those information atoms. The downward causality is defined by $\mathcal{D}(X_{t},X_{t+1})$, which encapsulating the information atoms that manifest a synergistic effect exclusively in the forward PID:
$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (32)
Conversely, the causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
$\mathcal{G}(X_{t},X_{t+1}):={I^{\{12\}\rightarrow\{12\}}_{\partial}}(X_{t},X_{t+1})$ (33)
As the $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that
$Syn(X_{t};X_{t+1})=\mathcal{G}(X_{t},X_{t+1})+\mathcal{D}(X_{t},X_{t+1}).$ (34)
Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) represented in double redundancy lattice.
Rosas also gives out a version of causal emergence quantification depending on a macro-variable, i.e., a coarse-graining strategy. If a system has the capacity to generate causal emergence, then there can be some macroscopic features that exhibit causal emergence. A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ if the complete state $X$ of the system at time $t$ is known with perfect precision, which is equivalent to $V_{t}$ is statistically independent of $X_{t+1}$ when $X_{t}$ is given. Then, for a system described by $X_{t}$, a supervenient feature $V_{t}$ is said to exhibit causal causation if:
$Un(V_{t};X_{t+1}|X_{t})>0.$ (35)
For this definition, the system’s capacity of causal emergence is required, where $Syn(X_{t};X_{t+1})>0$, since $Un(V_{t};X_{t+1}|X_{t})\leq Syn(X_{t};X_{t+1})$ holds for any supervenient feature $V_{t}$. Corresponding to the classification of system’s capacity, the Downward Causation exists of a feature variable $V$ when:
$Un(V_{t};$ $X_{t+1}^{1}|X_{t})>0$ (36)
$or$
$Un(V_{t};$ $X_{t+1}^{2}|X_{t})>0$
and the Causal Decoupling exists when
$Un(V_{t};V_{t+1}|X_{t},X_{t+1})>0,$ (37)
which also needs to be premised on the capacity of the system. Furthermore, $V_{t}$ is said to have pure causal decoupling if $Un(V_{t};X_{t+1}^{\alpha}|X_{t})=0$ and $Un(V_{t};X_{t+1}|X_{t})>0$. If all the emergent features exhibit pure causal decoupling, then the system is said to be perfectly decoupled.
Figure 13: Synergistic causes can be further divided into causal decoupling and downward causation.
Although a rigorous quantitative definition for causal emergence is proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems. Also, there are many calculation methods in the PID framework, which makes the quantitative definition determined by the specific quantitative method when applied. Therefore, to solve those problems, Rosas also relaxed the calculation methods of causal emergence, thus obtaining sufficient conditions for the two phenomena as identification criteria.
Specifically, in order to avoid exploring the specific quantitative relationship of synergistic and redundant information, this criteria repeatedly subtracts redundant information to make positive results a sufficient condition for emergent phenomena. Although this makes the quantitative method lose some of its generalizability, it improves the reliability of the results. The three indicators to be used are:
$\Psi_{t,t+1}(V):={I(V_{t};V_{t+1})}-\sum_{\begin{subarray}{c}j\end{subarray}}I(X_{t}^{j};V_{t+1}),$ (38)
which measures the mutual information between macro-variables across two time steps minus the information between micro-state to macro-state, where $V$ is a predefined macro-variable, but how to find such macro-variable is never discussed in Rosas et al. (2020).
$\Delta_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}\left({I(V_{t};X_{t+1}^{j})}-\sum_{\begin{subarray}{c}i\end{subarray}}I(X_{t}^{i};X_{t+1}^{j}))\right),$ (39)
which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$ in next step minus all the summation of mutual information between $X_{t}^{i}$ and $X_{t+1}^{j}$.
$\Gamma_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}{I(V_{t};X_{t+1}^{j})},$ (40)
this is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$
The specific functions of the indicators are as follows: $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$. Similarly, $\Delta_{t,t+1}(V)>0$ is as sufficient condition for $V_{t}$ to show downward causation. Finally, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ is a sufficient condition for causal decoupling to occur.
In summary, Rosas proposes an approach to quantitatively describe causal emergence by establishing the relationship between causal emergence and the synergistic effect of variables in different times, and further classified causal emergence by $\phi$ID. The definition not only objectively judges whether the system has the capacity to produce causal emergence, but can also measure the characteristics of causal emergence of specific macro features. The most notable contributions of his work were to bring the study of emergence closer to quantitative and empirical scientific research, to classify types of causal emergence, and to complement philosophical discussions about emergence.
There are some applications of PID or $\phi$ID. Recently, Luppi et al. apply $\phi$ID method on brain BOLD signals Luppi et al. (2022) to identify the synergistic core of the brain. They found synergistic information is for the integration of different parts to a whole, and redundancy is for robustness.
With the development of partial information decomposition techniques, we can further decompose the mutual information between two variables. This may help us to reveal the properties of systems with more aspects. For example, in  Varley and Hoel (2022), the authors first decompose the mutual information of a system, and calculate an indicator of synergy bias according to partial information decomposition method Williams and Beer (2010) to characterize how synergistic information distribute on different levels of the system. Higher synergy bias indicates that more partial information is in synergistic relationships. Then, the authors found that in some systems with occurrence of causal emergence, if the system is coarsen to a reduced system, the synergy bias will increase. Therefore, partial information is converted from redundancy to synergy as we coarse-grain the system. The conclusion is emergence is a kind of information conversion.
3.8.5 Comparison with Hoel’s framework
When comparing Hoel’s framework to Rosas’ quantification framework of causal emergence, several clear advantages can be observed in the latter. Firstly, Rosas’ theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal. Secondly, it offers a detailed decomposition of causal emergence, specifically downward causation and causal decoupling. Lastly, it effectively avoids cases of fake causal emergence, where the macro-variable depends solely on unique or redundant information from the micro-variables.
However, there are also some disadvantages to consider. Firstly, in order to obtain the full information lattice, a systematic iteration of all variable compositions is required. Additionally, despite the use of formula 35, it is necessary to define a macro-variable. Unfortunately, the authors do not provide any method to identify such a variable. Secondly, all the mutual information and its decomposition are based on correlations rather than causality. It is crucial to discuss how to incorporate causal elements, such as intervention and counterfactuals, into the framework.
Finally, the issue of identifying whether causal emergence occurs in a system based on the given time series data of its behaviors has not been addressed in the preceding discussion. To tackle this problem, the application of emerging technologies in machine learning and artificial intelligence is required. These technologies can provide valuable tools and techniques for detecting and analyzing causal emergence.
4 Causal Emergence and Artificial Intelligence
Recently, newly emerging artificial intelligence technologies have made significant breakthroughs in tackling a range of important and challenging problems. Examples include defeating human champions in complex games like Go Silver et al. (2016); Dabney et al. (2020), and predicting the intricate structures of protein folding Senior et al. (2020). These achievements have been made possible through the application of machine learning methods, which leverages well-designed neural network architectures and automatic differentiation techniques. Machine learning methods, with their diverse and well-designed neural network structures, along with automatic differentiation techniques, have demonstrated the ability to efficiently approximate complex functions within a vast functional space. Given this capability, it raises the question of whether we can leverage these emerging techniques to address the challenging problem of finding effective coarse-graining strategies or optimal decomposition methods for complex systems using time series data?
The answer is yes. We will introduce two latest works on causal emergence identification from data based on neural networks. Interestingly, these works indicate that causal emergence identification problem can also provide us with insight into connections with another related field, causal representation (machine) learning although causal emergence identification and causal representation learning are totally different research areas for distinct problems from appearance, they share commonalities. Furthermore, we also review the works on causal model abstractions and reinforcement learning based on world models as extensions because similar machine learning architecture are shared by these machine learning frameworks.
4.1 Causal Emergence Identification
In the previous section we have introduced several works on quantifying emergence by causality and other information theoretic conceptions. Among which, all works try to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence. However, in practical applications, in order for the theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time series data of dynamic systems, and give explanations for the results.
4.1.1 Rosas’ Identification Method
The first method for causal emergence identification is given by Rosas’ paper in causal emergence Rosas et al. (2020). Actually, the three indicators, the Equations 38, 39 and 40 as mentioned in section 3.8.4. The identification criteria have been exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
Although these three indicators avoid the problem of redundant information calculation, they are all sufficient conditions, that is, indicators greater than 0 can indicate emergence. The construction of this indicator repeatedly cuts out redundant information, which makes it almost impossible to identify the emergence of this indicator in systems with large redundant information or systems with many variables, and this is the case for most of the systems we face. Another defect of this method is that the coarse-grained variable $V$ must be given in prior, and different $V$ affects the results in a large degree. Therefore, the automatic coarse-graining strategy design according to data is still remain unsolved. New methods are required.
Therefore, the current theoretical frameworks of causal emergence do not give a good workable identification algorithm. Although existing works have tried to propose some workable methods based on static network structures and approximation for information decomposition, a general method for general Markovian dynamic systems is still lacking. One of the important reasons is to identify causal emergence within data needs to search for all the functions of coarse-graining or decomposing subsets for a complex system. However, the conventional numeric methods can not afford the computational cost for such search in a huge functional space. New methods are required.
4.1.2 Neural Information Squeezer (NIS)
Here we introduce one of the latest works on this direction which is just published in this special issue. In this paper, the authors formulate the problem of causal emergence identification on continuous Markov dynamics. Suppose the time series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
$\frac{d\boldsymbol{x}}{dt}=g\left(\boldsymbol{x}(t),\xi\right),$ (41)
Where $\xi$ is a Gaussian noise. The time series $\{\boldsymbol{x}_{t}\}$ are the observations of the micro-states of the system, and the problem that we confront is identifying if the causal emergence occurs for the original system. And we also want to know the appropriate coarse-graining strategies, the macro dynamics which correspond to the coarse-graining method, and the dimension of the phase space on macro-level if the causal emergence occurs.
To solve the problem, the authors propose a mathematical framework by converting the original issue into an optimization problem for functions. That is, we can find the appropriate coarse-graining strategy $\phi_{q}$, which is defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where, $q<p$ is the dimension of the macro-states, and the macro-level Markov dynamics $f$, which is defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$, the dimension averaged EI of the macro-dynamics (see Equation 24), that is,
$\max\mathcal{J}(\hat{f}_{\phi_{q}}),\qquad$ (42)
However, this problem has a trivial solution, e.g. $\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity. Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy within all $q$ dimensional functions. This is trivial because too much information has been abandoned by the coarse-graining strategy such that the macro-dynamic is useless.
To conquer the problem, the authors add a reasonable constraint to the original optimization framework, which is called effectiveness requirement. A coarse-graining strategy and the corresponding macro-dynamic are effective, which means there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state instead of the macro-state in the next time step according to the state in the previous step, the constraint can be written with,
$||\phi^{\dagger}(f(\phi(\boldsymbol{x}_{t})))-\boldsymbol{x}_{t+1}||<\epsilon,$ (43)
for all time step $t$, where, $\epsilon$ is a given constant. In this way, an effective coarse-graining strategy and the macro-dynamic can not only maximize the EI but also reproduce the original micro-dynamic as much as it can.
This constraint is an important complimentary for the framework of EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states which can solve the problem of ambiguity of macro-states. This constraint is also a guarantee of commutativity. To see this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ on both sides of Equation 43 to obtain $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
However, the problem is still hard to solve. The paper Zhang and Liu (2022) proposes a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step, and maximize EI for different $q$ in the second step. The first step can be solved by training a neural network easily, and the second step converts the complex functional optimization problem into a line search in one dimensional space.
Figure 14: The architecture and workflow of the Neural Information Squeezer (NIS). NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$. The encoder $\phi$ comprises an invertible neural network $\psi$ and a projection operator $\chi$, which simply discards fixed dimensions output by $\phi$. On the other hand, the decoder $\phi^{{\dagger}}$ incorporates the same invertible neural network as the encoder but implements its inverse $\psi^{-1}$. In cases where the input is incomplete, it can be complemented using Gaussian random noise.
To realize the abstract mathematical framework, the authors propose an encoder-dynamics learner-decoder framework of neural network, named neural information squeezer (NIS), which describes the basic working principle of the framework vividly. To reduce the complexity of the encoder-decoder which corresponds to $\phi$ and $\phi^{\dagger}$, respectively, the authors use an invertible neural network to implement the encoder, and invert the input-output workflow of the same neural network to realize the computation of the decoder. Concretely, the encoding process is decomposed into two step computation, information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on $p$ dimensional input space, and information discarding with the simple projection operation, that is:
$\phi(\boldsymbol{x})=\chi_{q}(\psi(\boldsymbol{x})),$ (44)
where $\chi_{q}$ is the projection operator to project a $p>q$ dimensional vector into a $q$ dimensional vector with the dimensions from $1$ to $q$ left. The decoding process also contains two stages, random sampling a Gaussian noise to complete the details of macro states, and information conversion but in an inverted way, that is:
$\phi^{{\dagger}}(\boldsymbol{x})=\psi^{-1}(\boldsymbol{x}\bigoplus\zeta),$ (45)
where, $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$ dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.
The adoption of an invertible neural network in the framework serves two main purposes. Firstly, by inverting the workflow of the neural network $\psi$ to obtain the decoder $\phi^{{\dagger}}$, a significant reduction in the number of parameters is achieved. This reduction can lead to improved efficiency and computational benefits. Secondly, the use of invertible functions in the neural network provides favorable mathematical properties that facilitate the analysis of the entire framework in a tractable manner. These properties enable researchers to gain insights and understand the characteristics of the framework more easily.
This architectural design with invertible neural network can not only reduce the complexity of computation, but can also separate the input information from the micro-states into noise and EI for prediction clearly. The authors proved a mathematical theorem stating that, as the neural network is converged by training, the mutual information on the adjacent macro states approaches to the mutual information of the two adjacent-step micro states in the data. This means, all the information dropped by the encoder is almost irrelevant with the prediction. All the useful information is squeezed into the information bottleneck of the dynamics learner at the macro-level. And another theorem has proved that the degree of information squeezing for micro-states is increased as we reduce the dimension of the macro states, and thus more and more useless information in the original micro-state data is thrown out by the narrower and narrower information bottleneck, the macro-state dynamics learner.
To solve the problem of causal emergence identification, we first optimize the neural networks, and the dimension averaged EI, $\mathcal{J}_{q}$ can be obtained for different $q$. When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension averaged EI for the micro-dynamics. Finally, the calculation of
$\Delta\mathcal{J}=\mathcal{J}_{q}-\mathcal{J}_{p}$ (46)
is the measure of dimension averaged causal emergence (see Equation 25. If $\Delta\mathcal{J}>0$, we can judge that causal emergence occurs.
Numeric examples are given to show the effectiveness of the whole framework. Interestingly, the neural information squeezer can work on the time series data with discrete micro-state although the whole framework is built based on continuous dynamical systems. Particularly, the model can work out the same coarse-graining strategy and macro-level dynamic as in the example of boolean network with four nodes in Hoel’s original paper, although any information about the node grouping and state mapping is not given to the framework.
It is also interesting to compare NIS with the mentioned frameworks or models in previous sections. Compare to the theory of computational mechanics, NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as the states, and when the whole framework is well-trained such that it can predict the future micro-states precisely, the encoded macro-states converge to effective states which can be treated as causal states in computational mechanics. However, the objective that maximize the effective information (EI) in NIS has no correspondence.
NIS also has similarity with G-emergence because it also adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step. However, there are several obvious differences between the two frameworks. In theory of G-emergence, the macro-state must be selected by hand at first, however, it is optimized automatically. Second, NIS uses neural networks to predict the future states, while, G-emergence uses auto-regressive techniques to fit the data. Due to the universal approximate theorem(see  Hornik (1991); Teshima et al. (2020)), neural networks are more advantage.
NIS provides a solution for causal emergence identification from data, and it is the first time to solve the problem of the automatic construction of coarse-graining strategy and the macro-dynamics from data for causal emergence identification.
Another important benefit of NIS which is to be noticed is that, what NIS has solved actually is a generalized causal emergence identification problem but not limited on a justification. Because after training, NIS can not only identify if causal emergence occurs, but also can obtain a neural network for coarse-graining the data and a neural network to simulate macro-dynamics. This may provide us richer information than just a justification.
4.1.3 Neural Information Squeezer Plus (NIS+)
One of the biggest problem of NIS is that it does not directly solve the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner). Instead, it searchs for the optimal dimension of the macro-state space $q$ such that EI can be maximized. Therefore, for given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimize Equation 43) but not for EI. Second, NIS can not support complex computations such as the coarse-graining functions containing multiple steps of information conversion and discarding.
To tackle these problems, NIS+ is proposed in ref Yang et al. (2023). Mathematically, the problem of EI maximization can be converted into a machine learning problem according to the definition of EI and the variational inequality and probability reweighting technique. After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},g_{\theta^{\prime}}$ with parameters $\omega,\theta$ and $\theta^{\prime}$, respectively. Formally, the problem of minimization of loss function problem without constraints can be written:
$\min_{\omega,\theta,\theta^{\prime}}\sum_{t=1}^{T-1}w(\boldsymbol{x}_{t})||\boldsymbol{y}_{t}-g_{\theta^{\prime}}(\boldsymbol{y}_{t+1})||+\lambda||\hat{\boldsymbol{x}}_{t+1}-\boldsymbol{x}_{t+1}||,$ (47)
where $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t}))$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t+1}))$ are the macro-states. $\lambda$ is a Lagrangian multiplier which will be taken as a hyper-parameter in experiments. $w(\boldsymbol{x}_{t})$ is the inverse probability weights which is defined as:
$w(\boldsymbol{x}_{t})=\frac{\tilde{p}(\boldsymbol{y}_{t})}{p(\boldsymbol{y}_{t})}=\frac{\tilde{p}(\phi(\boldsymbol{x}_{t}))}{p(\phi(\boldsymbol{x}_{t}))},$ (48)
where $\tilde{p}$ is the new distribution of macro-states $\boldsymbol{y}_{t}$ after intervention for $do(\boldsymbol{y}_{t}\sim U_{q})$, and $p$ is the natural distribution of the data. In practice, $p(\boldsymbol{y}_{t})$ is estimated through kernel density estimation (KDE) Rosenblatt (1956). The approximated distribution, $\tilde{p}(\boldsymbol{y}_{t})$, is assumed to be a uniform distribution, which is characterized by a constant value. Consequently, the weight $w$ is computed as the ratio of these two distributions. Theorems are proved to guarantee the correctness of this conversion. The overall framework of NIS+ can be depicted in Figure 15.
Figure 15: The workflow and architecture of our proposed framework, Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data. (a) Various forms of input data. (b) The framework of our proposed model, NIS+, which incorporates our previous model, NIS. The boxes represent functions or neural networks, and the arrow pointing to a cross represents the operation of information discarding. $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, while $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state. $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})$ represent the macro-states obtained by encoding the micro-states using the encoder. $\hat{\boldsymbol{y}}_{t}=\phi(\hat{\boldsymbol{x}}_{t})$ and $\hat{\boldsymbol{y}}_{t+1}=\phi(\hat{\boldsymbol{x}}_{t+1})$ represent the predicted macro-states obtained by encoding the predictions of micro-states. The mathematical problems that each framework aims to solve are also illustrated in the figure. (c) The various output forms of NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
Additionally, to apply NIS+ to the data generated by the complex systems of multi-agent systems and cellular automata, the structures of encoder can be extended by defining two kinds of combinations: one is to stack the basic encoder, the other is to concatenate the basic encoder. Here, a basic encoder is the composition of an invertible neural network and a projection operation. By inverting the directions of all the basic encoder, and supplements the empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
Numerical experiments demonstrate the effectiveness of NIS+ in capturing emergent behaviors and finding causal emergence in various simulations, including SIR dynamics, bird flocking (Boid model), and the Game of Life. Additionally, NIS+ is applied to find emergent features in real brain data from 830 subjects while watching the same movie. Experiments also demonstrate that the learned dynamical model exhibits improved performance in terms of out-of-distribution generalization compared to other models.
4.2 Implications in Causal Representation Learning and Out-of Distribution Generalization
We found that, as a framework of machine learning, neural information squeezer and plus can be regarded as a kind of framework for causal feature learning Chalupka et al. (2017); Spirtes et al. (2000); Chalupka et al. (2015) or causal representation learning Schölkopf et al. (2021). Causal feature learning aims that to find a high-level causal relations from the micro-level data. In causal representation learning the macro states are causal variables or causal representations Schölkopf et al. (2021), macro-dynamics are causal mechanisms, and the learned coarse-graining strategy is the representation or encoding process. In this sub-section, we will introduce causal representation learning in details at first, and compare the two emerging fields.
4.2.1 Introduction of Causal Representation Learning
Causal representation learning is an emerging field in artificial intelligence, it tries to combine two important fields in machine learning: representation learning and causal inference, together.
Representation learning aims to extract important features (or representations) hidden in data to make predictions automatically Bengio et al. (2013). Deep neural networks are always adopted, important features in data are represented by the activation modes of neurons LeCun et al. (2015); Goodfellow et al. (2016). The modes in different layers usually represent significant features in various scales, and deeper layers always correspond to larger scales Alzubaidi et al. (2021). The features in different layers are combined together to make predictions or classifications. For example, a deep neural network for human face recognition may have two components. The first one is a deep neural network with stacked convolution units LeCun et al. (1998), and it tries to extract features of faces on different scales. The second component is always a common feed-forward neural network which can accept the input features of faces and output a label to identify the character represented by the input image . The first component is for representation, and the second one is for classification. The neural network as a whole must learn to extract important features from images to make high precision classification.
Although representation learning has achieved remarkable success in various domains such as image classification LeCun et al. (2015), face recognition Wang et al. (2018), language understanding Mikolov et al. (2013); Devlin et al. (2018); Vaswani et al. (2017), and game playing Mnih et al. (2015); Dabney et al. (2020), it suffers from a critical limitation: it can only capture the information or associations within the data, but not the underlying causal relationships. According to the theory of causal hierarchy by Judea Pearl Pearl et al. (2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship which is invariant in different environments, while the latter may be more dependent on contexts and reflect the limitation of the data. For instance, in the example of face recognition mentioned previously, if all the images in the training data have outdoor backgrounds, e.g. grasslands, lakes, trees, etc., the neural network may not work well on the images with indoor backgrounds Cui and Athey (2022); Zhang et al. (2021). This is because the neural network may learn the fake association between backgrounds but not foregrounds and labels. However, the association is only the reflection of the bias in the training data, but not the real causal relationship.
In causal inference, people use graphs to represent causal relationships. Nodes represent causal variables, and edges which are always combined with other information like structural equations or conditional probabilities to represent causal relations Pearl et al. (2000). In conventional causal inference, while the parameters in a causal model can be adjusted based on data, all the variables and causal relationships are typically extracted manually rather than directly from data itself.
To conquer the problem, causal representation learning tries to combine the advantages of the two sub-fields to extract the important features and the relationships with causality behind data automatically Schölkopf et al. (2021).
Figure 16: Illustration for the workflow of a typical causal representation learning agent. In the top panel, we have a depiction of a real block world with a robotic arm, along with the corresponding causal graph representing the relationships between the variables in the real world, nevertheless, it is totally unknown for the agent. Additionally, we have the data in the form of images generated by the block world. In the bottom panel, we see the structure of the causal representation learning agent. It utilizes a convolutional neural network (CNN) to extract meaningful representations from the data and perform downward tasks based on these representations. The encoder plays a crucial role in capturing relevant features and patterns from the input data, enabling the agent to learn and understand causal relationships.
A typical scenario of causal representation learning can be depicted by Figure 16. Suppose there are a set of variables and causal mechanisms (including causal graphs and structural equations) between them which can describe how the world of robotic arms and colorful blocks works, whereas, they are not observable for the agent of causal representation learning. What the agent can observe is a set of images which are generated by robotic arms and blocks. The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks.
Formally, if the input data can be described by $X=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{p})$, causal representation learning will construct two functions and a directed acyclic graph, which include a causal representation function: $\tilde{\phi}:\mathcal{R}^{q}\rightarrow\mathcal{R}^{p}$, such that:
$X=\tilde{\phi}(Y),$ (49)
where, $Y=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},...,\boldsymbol{y}_{q})$ is a set of causal variables and $q\ll p$; a directed acyclic causal graph $\mathcal{G}_{S}\equiv(Y,E)$, on which the node set is $Y$, and the directed links in $E$ represent causal relations; as well as a set of causal mechanisms:
$\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i}),$ (50)
for $i=1,2,…,q$, where $f_{i}$ is the structural equation for the $i$th causal variable $\boldsymbol{y}_{i}$, which describes how the parent causal variables $PA_{i}\equiv\{j|j\rightarrow i\in E\}\subset\{\boldsymbol{y}_{i}\}$ and the external unknown variable $U_{i}$ determine the value of $\boldsymbol{y}_{i}$. And there is a requirement for $U_{i}$ for all $i,j$, that is:
$P(U_{i},U_{j})=P(U_{i})P(U_{j}).$ (51)
Thus, any two external unknown variables need to be independent each other.
However, learning the three elements including the two functions and a graph in the same time is very hard, we can convert the problem to reduce the difficulty. This is to learn disentanglement representations, in which the representations are independent(disentangled) each other Wang et al. (2022). We will show how we can derive disentanglement representations from Equations 49 to 51.
Notice that if we combine the equations $\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i})$ for all $i\in[1,q]$ together, we have $Y=F(Y,U)$, where $U=(U_{1},U_{2},...,U_{n})$, and $F=(f_{1},f_{2},...,f_{q})$. That means, the causal mechanisms actually describe how the unknown variables $U$ determine the causal variables $Y$. Suppose we can solve this equation to obtain $Y$ expressed by $U$, and we define $\tilde{F}$ such that: $Y=\tilde{F}(U)$, according to Equation 49,
$X=(\tilde{\phi}\circ\tilde{F})(U)$ (52)
if we define $\phi\equiv\tilde{\phi}\circ\tilde{F}$, then $\phi$ can be understood as a decoder which can decode a disentanglement representation to data. Further, if we could construct a encoding function $\phi^{{\dagger}}:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ which takes $X$ to $U$, and we require that:
$\phi\circ\phi^{{\dagger}}\sim I,$ (53)
where, $I$ is the identity mapping, thus, $\phi\circ\phi^{{\dagger}}$ is an auto-encoder, which can encode the input data $X$ into a latent space $U$ and reconstruct $X$ according to $U$. While, according to Equation 51, the latent variables are independent each other. That is, we represent the input data by a set of independent or disentanglement latent variables, and we can reconstruct the input data according to these independent variables. Therefore, this is called disentanglement representation learning. In practice, independent component decomposition method can be use to construct the decoder Hyvarinen and Morioka (2016); Khemakhem et al. (2020). To reduce the computational complexity, we can also use invertible neural network to model $\phi$ Dinh et al. (2014).
With the disentangled representation variables $U_{i}$, we can effectively perform downstream tasks. One prominent application of disentangled representation learning is causal discovery. By leveraging the disentangled representations, we can uncover causal relationships between variables and gain insights into the underlying causal structure of a system Monti et al. (2020). Causal discovery allows us to identify cause-and-effect relationships and understand how different factors interact and influence each other. This knowledge can be invaluable in various domains, such as understanding complex systems, making informed decisions, and designing interventions or policies.
To conclude, a typical causal machine learning or disentanglement representation learning framework is always composed of a module for causal or disentanglement representations and a module which contains sub-units for different downstream tasks. The representation module, which contains an encoder and a decoder, tries to extract abstract representations of the data, and the representations are not for prediction but for causal inference. The downstream tasks may be variant. For example, the transferable mechanism learning tries to look for independent causal mechanisms to complete prediction tasks in very different environments Peters et al. (2016). And existing studies have shown that learning causal models may help in transferring modules across different domains Cui and Athey (2022); Kuang et al. (2018). Causal representation learning has been applied to various problems.
4.2.2 Comparisons between Causal Emergence Identification and Causal Representation Learning
There are similarities and differences between causal emergence and causal representation learning (CRL). We will introduce them in detail separately.
Similarities
It is interesting to compare the tasks of causal emergence identification under the theoretic framework of Erik Hoel and causal representation learning. Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro variables are the causal variables, the macro dynamics are the causal mechanisms, the coarse-graining strategy is an encoding process from the original data to the representations. Under this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism. To identify causal emergence, we need to learn the appropriate causal variables to represent data and to discover causal mechanism on the macro-level (on representational latent space) such that the EI of the learned causal mechanism in latent space is larger than the original data. We use Table 1 to clarify the comparisons item by item.
Table 1: Comparison between causal emergence identification and causal representation learning.
Causal Emergence Identification Causal Representation Learning
Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real
Latent Variables Macro-states Causal representations
Causal mechanism Macro-dynamics Causal mechanisms
Mapping between data and latent variables Coarse-graining Representation
Optimization for causality EI maximization Prediction Loss, Disentanglement
Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has stronger causal effect Finding an optimal representation of the raw data such that the independent causal mechanism can be realized by the representations.
With these similarities, the technologies and concepts from both fields can learn from each other. For example, the techniques of causal representation learning can be applied to discover causal emergence.
On the other hand, the learned abstract causal representations can be viewed as macro-states, which enhances the explainability of causal representation learning. This perspective allows us to consider that CRL is actually uncovering the hidden causal emergent features within the data.
Moreover, these similarities between emergent phenomena and CRL contribute to a deeper understanding of emergence itself. By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract a more profound causal structures. These profound causal structures may serve as the origins of the mysteries surrounding emergence. By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we gain insights into the mechanisms behind emergent phenomena. These causal structures provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components. Exploring these intricate causal structures brings us closer to unraveling the enigma of emergence and sheds light on the fundamental principles governing complex systems. Consequently, the identification of emergent phenomena requires an observer, whether it be a human or a machine, to actively search for them.
By employing a machine observer, it is possible to potentially resolve the longstanding debates regarding whether emergent causality is epistemic or ontological. This is because the machine observer can provide a relatively objective outcome, although it is important to note that the result is still influenced by the observer. Furthermore, an intriguing area for future exploration involves investigating how the structure and capabilities of the observer can impact the emergence results that are identified by modifying the parameters of machine learning. This highlights the potential influence of the observer’s characteristics on the outcomes obtained through the rewiring of machine learning parameters.
Differences
However, there are several differences between the CRL and causal emergence in theory. The biggest gap is that CRL assumes that there is a real causal mechanism behind it, and the data is generated by this causal mechanism Schölkopf et al. (2021). While for causal emergence identification, the emergent variables and mechanisms on macro-level are just handy ways to observe and understand data, "real causality" may not exist. The difference can be illustrated by Figure 17
Figure 17: The Comparison between Causal Emergence Identification and Causal Representation Learning. All the potential co-founders are ignored. The black arrows represent causal relations, while the blue arrows represent interactions between macro-variables. The dashed directed lines represent the information flows for coarse-graining, representation or data generation.
While, if epistemological perspective is adopted, this difference disappears because all they do is to extract meaningful information from observational data to obtain representations with stronger causal effects.
Another major difference is that the causal mechanism for macro-states in causal emergence is a dynamical system. Therefore, if the dynamics in macro-state space is on a network, circular structures may exist because the state variables will be iterated in the dynamical system. As a result, feed-backs and circular indirect interactions are allowed in such model. However, the causal mechanism is always represented by a structural causal model in CRL. While, these models are always directed acyclic graphs, therefore, loops or circular structures are not allowed. The problem is that the meaning of the links in a network dynamics model is different from the links in structural causal model. Hence, the conflict can be resolved if we comprehend that the links in a structural causal model depict an average (net) causal effect between variables when they are in a steady state. On the other hand, the links in network dynamical models within causal emergence represent instantaneous interactions.
Finally, unknown co-founders are always ignored in causal emergence, while, they play very important roles in causal structural models.
4.2.3 Out-of-distribution Generalization
Causal representation learning is established in a self-supervised manner, in the sense that it reveals the intrinsic causal structure of data. Thus, the learned causal representations could not be exploited directly for supervised prediction task, where the causation between data and a specified target must be characterized. The measurement of causal emergence by vanilla Effective Information Hoel et al. (2013) is also defined for stationary dynamics systems in a self-supervised fashion, since the input(cause) space and the output(effect) space share the support. However, Hoel generalizes Effective Information to general input-output systems Chvykov and Hoel (2020)(see section 3.6.1 where the cause and the effect space take arbitrary support, which could either be discrete or continuous.
Empowered by the general notation of EI, causal emergence could be applied in supervised learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or a representation learned upon which) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause(feature) to the effect(target). It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffice for common prediction tasks with an i.i.d. assumption, implying that the training and test data are independently and identically distributed.
However, if samples are drawn out of the training distribution, it is essential to learn a representation space that generalizes from training to test environment, known as out-of-distribution (OOD) generalization. Due to the common belief that causal relations generalize better than statistical correlation Arjovsky et al. (2019), causal emergence theory could serve as a criteria for the causation embedded in the representation space. The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regarding to out-of-distribution generalization.
Introduction of OOD
Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic. Out-of-distribution generalization towards arbitrary distributional shifts is generally impossible Ye et al. (2021). Therefore, it is commonly assumed there exists an underlying causal mechanism that governs the data generation process, which remains invariant to distributional shifts. Formally, by Pearl’s Structural Equation Model Pearl (2009) which causally formulates data generation, the target $Y$ is supposed to be generated from part of the feature $X$, i.e. $Y=f(X_{c},\epsilon)$ where $X_{c}$ is selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$. Obviously $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments. On the other hand, the raw feature $X$ might be a mixture of causal variables and non-causal variables. The latter could be produced by confounding, selection bias or anti-causal mechanisms, all of which introduce spurious correlation between the target and non-causal variables. The strength and direction of spurious correlation are variant across environments, resulting in distributional shifts that hampers out-of-distribution performance of machine learning techniques. As a remedy, a prominent branch of OOD algorithms Arjovsky et al. (2019); Huh and Baidya (2022); Zhao et al. (2019); Ganin et al. (2016) learn an invariant representation space distilling the causal variables from the raw feature space. Typically a metric of the representation is devised to verify the stability of correlation between the learned representation space and the target, and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features. The paradigm for invariant representation learning could be formulated as Wang et al. (2022):
$\min_{f,\phi}\mathcal{R}(f(\phi(X)),Y)+\lambda\ell_{reg}(\phi),$ (54)
where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function and $\ell_{reg}$ is the regularization.
Representation learning based OOD algorithms are diversified in their constraint of the representation space characterized by the regularizers. Each metric of the representation fits a specific data generation pattern, and does not guarantee to filter out the causal variables when the setting is violated. However, causal emergence potentially offers a united metric of representation for OOD generalization rooted from causality theory.
Before diving into the application of causal emergence, we would first discuss representative regularizers of representation learning and their limitations. Domain invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments. The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al. (2018) or an adversarial classifier for the environment index Ganin et al. (2016). However, the alignment constraint provides generalization guarantee only for covariate shift Zhao et al. (2019) where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$. On the other hand, invariant risk minimization Arjovsky et al. (2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$. Invariant risk minimization is proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target, but is reported to fail under covariate shift Ahuja et al. (2021). Further, conditional invariant representation Li et al. (2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect. The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
Actually, it is theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models and it does not generalize across all data generation processes Kaur et al. (2022). As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, which hinders the general applicability of invariant representation learning. Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedure, increasing the data collection and labelling expenses. To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
Causal emergence and OOD
The match between OOD generalization and causal emergence is attributed to the do intervention of Effective Information. By definition Pearl (2009), intervening the feature distribution with $do(X)$ modifies the structural equations of all the variables in feature space to a constant value while the equation for the target remains intact as $Y=f(X_{c})$. On the one hand, the assumption of causal mechanism’s invariance to distributional shift is respected, since the causal correlation between causal variables and the target is reserved. On the other hand, the spurious correlation between non-causal variables and the target are removed by the intervention. For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation by $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$. As a result, $X_{n}$ is no longer caused by $Y$ and the spurious correlation vanishes. For another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$. Then the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist. In summary, the intervention introduced by EI captures the causal correlation between features and the target whereas suppresses the spurious correlation. It renders EI an ideal metric to assess the causality contained by the representation space and thus the generalization of the induced algorithm.
Recall that an increment in the Effective Information implies the occurrence of causal emergence. Therefore, we bring forward a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning. Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e. $I(g(X);Y)\leq I(X;Y)$, by data processing theorem. Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space. In contrast, EI(or normalized EI) would possibly peak at a medium stage of abstraction of raw feature, coinciding with the philosophy of OOD generalization that less could be more. Ideally, at the peak of EI(or normalized EI) where causal emergence occurs, all non-causal features are excluded and the causal features are revealed, resulting in the most informative representation while performing invariant to distributional shifts.
Actually, removing spurious correlation and recovering causal mechanisms by manipulating feature distributions is shared by the reweighting technique widely adopted in OOD literature. For instance, Stable Learning Shen et al. (2020a, b); Zhang et al. (2021) is oriented at the setting where collinearity exists in input variables, causing spurious correlation between non-causal variables and target. It reweights training samples to decorrelate the features and thus reducing collinearity and eliminating spurious correlation. Similar to EI, Stable Learning is free of environment index labeling. Further, sample reweighting and feature decorrelation shares the philosophy of distribution intervention. In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
4.3 Other Related Topics in AI
There are other related topics in artificial intelligence, e.g., abstraction of causal models and world model based reinforcement learning. We will introduce them in detail.
4.3.1 Abstraction of Causal Models
In CRL, we have the ability to extract causal relationships and construct causal models from raw data. However, in cases where the causal model becomes excessively large and complex, it becomes necessary to employ an abstraction process Rubenstein et al. (2017). Abstraction in causal modeling refers to the reduction or simplification of the causal model while preserving its essential causal relationships and important features.
Coarse-graining is the multi-scale description of the state information on different information reduction scales. While causal models also can be multi-scale Chalupka et al. (2016). The causal model in the higher level is an abstraction of the model in the lower level.
If the original causal model is given, then the key problem is how can we obtain an abstraction of the original model. Obviously, the abstraction of the causal model cannot be arbitrary, it must recover the main features of the original causal model. Several works proposed different standards for abstract models Rubenstein et al. (2017); Beckers and Halpern (2019); Chalupka et al. (2016), among which approximate abstraction is a good and reasonable proposal. The basic idea behind the method of approximation abstraction is that the differences of causal mechanism both in the abstraction and the original causal model should be minimized Rubenstein et al. (2017); Beckers and Halpern (2019); Beckers et al. (2020); Beckers (2021).
Figure 18: An approximate abstraction of the El Niño phenomenon.
For example, we can build a causal model to describe how the Equatorial Pacific Zonal wind ${W}$ can have a causal effect on the temperature of the Equatorial Pacific ${T}$ Chalupka et al. (2016). We write the causal variable $W$ as $X_{C}$, the effect variable $T$ as $X_{E}$, and this is the original causal model, which can be denoted as $\mathcal{M}_{m}=((\{{U}\},\{{X_{C}},{X_{E}}\},\mathcal{R}_{m}),\{g_{m},f_{m}\},\mathcal{I}_{m})$, where ${U}$ is the collection of all the unknown variables, ${X_{C}}$ and ${X_{E}}$ are causal variables, $g_{m}$ and $f_{m}$ are causal mechanisms to describe how ${U}$ and ${X_{C}}$ can affect other variables, $\mathcal{I}_{m}$ is the set of interventions. However, we care about higher level causality when we consider the El Nino or La Nina phenomena. Thus, we can build an abstract causal model $\mathcal{M}_{M}=((\{{U}\},\{Y_{C},Y_{E}\},\mathcal{R}_{M}),\{g_{M},f_{M}\},\mathcal{I}_{M})$, where $Y_{C},Y_{E}$ are new causal variables which are built by the coarse-graining mapping $\phi$, that is $Y_{C}=\phi({X_{C}})$, $Y_{E}=\phi({X_{E}})$. And $g_{M}$ and $f_{M}$ are causal mechanisms of $Y_{C},Y_{E}$ on the abstraction model. However, if $\phi$ is given, $g_{M}$ and $f_{M}$ are limited otherwise inconsistencies will exist. In  Beckers et al. (2020), the authors propose a limitation on $g_{M}$ and $f_{M}$. We denote $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ as the distance between different level models $\mathcal{M}_{m}$ and $\mathcal{M}_{M}$. They claim that $\mathcal{M}_{M}$ is an $\epsilon$-approximate abstraction of $\mathcal{M}_{m}$, if and only if $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ must be less than $\epsilon$ Beckers et al. (2020). That is,
$d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})=\min_{\phi_{\mathcal{U}}}\max_{do(X=x)\in\mathcal{I}_{m},{u}_{m}\in\mathcal{R}_{m}(\mathcal{U}_{m})}$
$\big{\{}d_{v_{M}}(\phi(\mathcal{M}_{m}({u}_{m},do(X=x))),\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x))))\big{\}}.$ (55)
In this equation, $\phi_{\mathcal{U}}$ is a coarse-grained mapping of unknown exogenous variables, $do(X=x)$ is an intervention on certain endogenous variables, $d_{v_{M}}$ is the distance between variables defined at a higher level, and $\omega_{\phi}$ is the mapping of interventions between different causal models at different levels. The distance between models is transformed into the distance between variables. Given a defined method of measuring the distance between variables (such as Euclidean distance), we can measure the distance between the output of a lower level model after coarse-grained projection$\big{(}\phi(\mathcal{M}_{m}({u}_{m},do(X=x)))\big{)}$ and the variables output by a higher level model$\big{(}\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x)))\big{)}$. We focus on identifying external variables and intervention methods that maximize the distance, and we do not concern ourselves with the coarse-grained mapping between external variables. Therefore, we select the optimal $\phi_{\mathcal{U}}$ and require that it be surjective Beckers et al. (2020). This yields a distance value that characterizes the inconsistency between different levels of causal models.
There are similarities between causal emergence identification problem and abstraction of causal models. If the original causal mechanism can be understood as a micro-dynamics, then the abstracted mechanism can be regarded as a macro-dynamics. Therefore, finding the optimal $\phi$ and causal mechanisms $g_{M}$ and $f_{M}$ is equivalent to finding an optimal coarse-graining strategy and macro-dynamics. In the framework of NIS, we set limitations on coarse-graining strategy and macro-dynamics by requiring that the macro-dynamics can recover the micro-dynamics with $\epsilon$ error to exclude trivial cases. This requirement is also similar to the approximation requirement for the distance.
In artificial intelligence, model abstraction is a widely used technique to obtain a compressed model. For example, the abstraction of a neural network model can be stored in cell-phone to conduct fast computation Wang et al. (2021); Pulina and Tacchella (2010). Another article discusses the importance of structural analysis methods for neural network analysis and proposes a new method based on a formal theory of causal abstraction. This method aligns neural representations with variables in interpretable causal models and uses interchange interventions to experimentally verify their causal properties. The method is utilized in a case study to examine neural models that have been trained on a complex Natural Language Inference (NLI) dataset Geiger et al. (2021). This dataset comprises numerous artificially processed, linguistically intricate English phrases and sentences. The analysis reveals that neural representations effectively capture the compositional structure of the provided examples Geiger et al. (2021). Model reduction is also important topics in control theory and dynamical systems Antoulas (2005).
The primary distinction between causal emergence identification and causal model abstraction lies in their approach to abstraction or dimension reduction. Causal emergence identification focuses on abstracting or reducing dimensions from the original data, whereas causal model abstraction focuses on abstracting from the micro-level causal models. In another word, coarse-graining operates on states or data, while abstractions operate on models. In physics literatures Christensen and Moloney (2005), abstractions of models are also called renormalization.
Another difference is unknown confounder $\mathcal{U}$ being considered in causal model abstraction is ignored in all the discussion of causal emergence. The emergence of causality requires the construction of models from a multi-scale perspective, and how to discuss mixed variables at different scales has always been a challenge. The theory of approximate abstraction takes into account the coarse-grained mapping of latent confounding variables $\phi_{\mathcal{U}}$, making it compatible with the coarse-grained function of endogenous variables. For the study of causal emergence, this theory lays a theoretical foundation for considering the confounding effects brought by environmental factors.
The approximate abstract framework currently has many limitations. Firstly, there is a lack of effective algorithms that can calculate the degree of approximation in the approximate abstract. Given the involvement of traversing intervention sets, it is difficult for machines to perform efficient numerical calculations that are not maintainable, and cannot be used as constraint conditions for automatic modeling. Additionally, the intervention space is very large, and some interventions are meaningless or impossible, so it is necessary to empirically provide an intervention distribution in advance Beckers et al. (2020), making it difficult for this framework to be applied to real-world problems.
4.3.2 World Models in Reinforcement Learning
In both the generalized causal emergence identification problem and causal representation learning, a common computational framework is shared, consisting of an encoder-decoder architecture and causal mechanisms or dynamics. This framework can also be integrated into a reinforcement learning agent, enabling the agent to learn a world model and make informed decisions and plans. Such an agent, known as a world model-based reinforcement learning agent, has demonstrated impressive performance across various environments and tasks Ha and Schmidhuber (2018).
Reinforcement learning is a significant sub-field of machine learning that possesses several distinctive features when compared to conventional supervised learning techniques Sutton and Barto (2018). Firstly, reinforcement learning agents have the capability to interact with the environment and actively acquire new information through interventions. Consequently, causal information can be obtained through these interactions Zeng et al. (2023). Secondly, in contrast to supervised learning where the processes of learning and action are clearly separated, reinforcement learning agents can seamlessly blend learning and action processes. Lastly, reinforcement learning agents demonstrate strong performance in environments that have delayed feedback, such as game playing scenarios Mnih et al. (2013); Silver et al. (2016).
The reinforcement learning algorithms can be classified into model-based and model-free types, and the world model based reinforcement learning belongs to the former Ha and Schmidhuber (2018); Hafner et al. (2023); LeCun (2022); Hafner et al. (2019a, b, 2020, 2023). As its name suggests, there is a model of the world which can simulate the dynamics of the environment the agent confronts. The dynamics of the world model can be learned via the interactions between the agent and the environment, and it can also help the agent to make plans and decisions about the uncertain environment Sekar et al. (2020). However, in order to represent a complex environment, the world model needs to provide a coarse-grained description of the environment. Consequently, a standard architecture of a world model typically includes an encoder and a decoder, as illustrated in Figure 19.
Figure 19: The architecture of a world model embedded in a model-based reinforcement learning agent, where $X_{t}$ is the environment at time $t$, and $Y_{t}$ is its representation, $H_{t}$ is the internal variable of the agent. Different from NIS and causal representation learning agent, another variable, the action in the previous step $A_{t-1}$, should be input to the world model to implement the one step dynamics. And the world model can be trained in a supervised learning manner by comparing the predicted representation of the environment and the prediction.
The world model can also be treated as a macro-dynamics or causal mechanism Zhang et al. (2020); Allshire et al. (2021); Jiang et al. (2022); Zhang et al. (2020); Wang et al. (2022). All the states of the environment can be regarded as macro-states or causal representations. These states are constructed to ignore irrelevant information and capture the most important causal features of the environment for the agent to make decisions. In the process of planning, the agent can use the world model to simulate the running of the real world dynamics. Numeric results show that the reinforcement learning algorithms with world models can perform better than the classic reinforcement learning algorithms.
World model-based reinforcement learning techniques can also be employed for identifying causal emergence. When attempting to identify the causal emergence phenomenon within a highly complex system, extracting macro states and dynamics solely from observational data becomes challenging. Explorations or experiments are required to facilitate this process Sekar et al. (2020). In such cases, a reinforcement learning agent equipped with a trainable world model, specifically a macro-dynamics model, becomes essential. By treating the complex system as a unified entity, the agent with the world model can interact with it and derive emergent causal laws from these interactions.
On the other hand, the exploration of causal emergence can offer valuable insights into reinforcement learning algorithms. Firstly, the measure of causal effect, EI, can be applied to world models. Techniques of maximizing EI can be employed to enhance the causal effect of world models within reinforcement learning agents Yang et al. (2023). This, in turn, can greatly benefit planning and problem-solving tasks, improving their efficacy and performance. It is like the bi-simulation Kemertas and Aumentado-Armstrong (2021); Zhang et al. (2020) which can train the agent with an internal metric.
Second, if the reinforcement learning agent with a world model can learn the emergent causal mechanisms automatically from a complex system that the agent is interacting with, then it will have more complex behaviors and be more intelligent. The agent may use the learnt emergent causal mechanisms to make plans. From learning to planning, the whole process may take place continuously. Therefore, causal emergence can be observed from outside because new behaviors governed by different causal mechanisms emerge as time goes by. This agent seems to have its own purpose or objectives.
4.4 Summary
In summary, there are deep connections between causal emergence and some hot topics in artificial intelligence such as CRL, abstractions, and world models. On one hand, the machine learning framework for causal emergence identification has similar structures to the frameworks of CRL. And even the concepts can have one to one correspondence as shown in Table 1. Therefore, the ideas and techniques of both sides can learn from each other. On the other hand, the studies of causal emergence can provide new perspectives and insights for causal representations and world model based algorithms. For example, we can explain that what the representations and causal mechanisms learned are emergent causal laws. Maximization of EI may help to improve the efficiency of learning the causalities for both CRL and reinforcement learning agents. Finally, we found a reinforcement learning agent with the ability to learn causal emergence may exhibit the ability of actions for purpose and objectives.
5 Potential Applications in Complex Systems
In this section, we will explore how the problems discussed in the previous sections relate to our understanding of complex systems. On one hand, the methods of causal emergence identification and the associated AI algorithms can be employed in complex systems to discover emergent patterns and causal relationships. On the other hand, the concepts of causal emergence can aid in our comprehension of complex systems, particularly in addressing significant issues such as consciousness and free will. These ideas have the potential to provide valuable insights and approaches to tackle these complex and profound problems.
5.1 Potential Applications
Complex systems typically consist of a large number of interacting units, and these units and interactions form various hierarchical and self-similar structures in a complex manner Mandelbrot and Mandelbrot (1982). In particular, when we examine complex systems from different scales, various structures and patterns may emerge at different scales, and these patterns may further interact with each other. Thus, a complex network of causality may form, and the nodes in this network may not only include the individual units of the complex system but also contains the emergent patterns in distinct scales.
As a result, it is crucial to determine how we can identify and extract emergent patterns and causal relationships from vast amounts of data generated by complex systems.
5.1.1 Emergence detection
In complex dynamical systems, various higher-level patterns, such as waves, periodic oscillations, and solitons, are ubiquitous. For instance, in the climate system, typhoons and tornadoes are emergent vortex structures. In urban areas, traffic jams also emerge as a result of interactions among a large number of cars. Social riots are another example of emergent events in higher levels of human society. Identifying these emergent patterns at an early stage is crucial and significant Haugen et al. (2023); Mnif and Müller-Schloer (2011). Therefore, there is an urgent need for a method that can automatically detect these emergent patterns and even provide early warning signals.
In O’toole et al. (2017), the authors regard emergent phenomenon as a specific event happened in a complex system, and detecting emergence is equivalent to detecting the formation and evaporation of this event. Therefore, emergence detection problem can be converted as a change-point detection problem of time series data.
Two methods are discriminated to detect emergence: one is centralized method, the other is decentralized. For example, in Mnif and Müller-Schloer (2011), the authors propose that the negative Shannon entropy can be used to measure the orderness of the entire system, and it can also indicate if emergence occurs. NIS and NIS+ are also centralized methods.
Another one is decentralized emergence detection method, in which the detectors are dispersed on the complex system, and implement event detection spontaneously and separately O’Toole et al. (2014).
As the development of artificial intelligence, some deep learning methods can be used to detect emergence. For example, dynamical image segmentation can be used to find interesting patterns in the video automatically Liu et al. (2021). A complex system can always be represented as a dynamical graph, e.g., a flock of birds, and when emergence occurs, the connectivity of this interaction network always increase. Therefore, we can detect the change of the network to find emergence Pazho et al. (2022). Other deep learning methods could be helpful.
5.1.2 Discovering emergent causality
Complex systems can be understood as a large network with events connected by causal links. A small disturb of one event may give rise a large roam of changes as butterfly effect indicates. While, sometimes, one single event happens with destiny which means it may be the effect of many different causes. How to discover the complex causal relationships from data in an automatic way is a challenging problem.
The research area of causal discovery tries to tackle this problem in many different ways Glymour et al. (2019); Zanga et al. (2022). For example, some studies try to use well-defined statistical indicators to characterize the causality between two variables, e.g., the measures of mutual information Donges et al. (2009), transfer entropy Stetter et al. (2012), etc. Nevertheless, these methods can not distinguish causal relationship from correlations. Another widely used method is independent causal discovery method Glymour et al. (2019), which tries to find the conditional independence among three variables in data, and derive Markov equivalent sub-graphs accordingly. Finally, binary causal interactions can be identified from these sub-graphs Kalisch and Bühlman (2007). Score based method tries to maximize or minimize a score function to find the optimal causal structures Zheng et al. (2018); Zhu et al. (2019). These methods always require a large amount of data and computation.
However, more challenges will be confronted when these causal discovery methods are applied in complex systems because circular causal structures and cross-level causality exist.
Circular causal structures widely exist in complex dynamical systems due to the feedback effect. That is, one variable may take effect on itself via a direct self-loop feedback or a circle with long chain. This circular causal structure may challenge the existed study methods of causality because most of these methods study directed acyclic causal graphs. Recent progress has developed some methods to discover these circular structures e.g.  Richardson (2013); Lacerda et al. (2012); Forré and Mooij (2018); Rantanen et al. (2020); Zhang et al. (2019); Pamfil et al. (2020); Vowels et al. (2022).
Another aspect is higher level or cross-level causation if the ingredients of scale and coarse-graining are considered. For example, downward causation describes the causal effect between macro-level and micro-level variables. Thus, the causal connections may be cross-leveled. If causal emergence occurs in a complex system, some strong causality may be found between macro-variables. All of these factors in complex systems present new challenges. The machine learning methods discussed in section 4.1.2 can address the question of whether there exists higher-level causality at the macro-level. However, it’s important to note that this is a global property. It must be further developed and extended to find local causal relationship between macro-variables or macro-variable and micro-variable Varley (2022). While, existed methods of causal discovery must be extended to take the method of grouping a set of variables or coarse-graining a system into consideration. For example, the article Cheng et al. (2023) proposes a coarse-to-fine causal discovery algorithm based on Granger causality and graph neural network. The grouping process of variables is from coarse-to-find which can improve the efficiency of the algorithm. Other multi-level causal discovery methods are also proposed and discussed in Wang et al. (2023); Zhang et al. (2023); Fan et al. (2023).
One of the potential applications of these complicated causal discovery method is to enhance the interpretabiity of large and complex neural networks. As we know, interpretability is one of the notorious problem for large neural networks, although they are very useful to solve many real problems. Thus, we can apply the methods of causal discovery, especially, the methods to discover emergent causality on these neural network models to find hidden and emergent causal relationships between neurons or group of neurons Chattopadhyay et al. (2019); Ren et al. (2023). If the degrees of causal effect can be computed and compared among these discovered causal relationships, the most important causation can be pointed out to solve a concrete problem. This may enhance the interpretability of the neural network. In this way, the misery of emergent ability of large language models may be dismissed by uncovering the hidden causation within large models Michaud et al. (2023); Wei et al. (2022).
5.1.3 Beyond causality
In this review, we mainly focus on the emergence of causality, however, we do not need to be bounded on this kind of special emergence, more macroscopic or global level properties could also be emergent although causal emergence may be the most important one. For example, individuality is one of the emergent property with no any connection with causality which has been discussed in previous section. In  Chang et al. (2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world. The authors then argued that this agent with information closure be regarded as a model of consciousness.
Symmetry may be another interesting property within some complex systems, and it may be emergent. For example, any single trajectory of a large number of random walkers in a two dimensional Euclidean space is random, nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic. This kind of symmetry can only be found on the macro-level, thus, is emergent.
It should be noted that EI can also be viewed as a unique indicator of symmetry, specifically, it is time reversal symmetry. The reason behind this is that when the macro-dynamics approaches a time reversible map (a bijection between states across a time step), EI is increased. And the maximum EI can be arrived when the Markov dynamics is reversible. This interesting fact should be further studied in the future.
It is interesting and useful to deploy a machine learning method to automatically discover this kind of emergent symmetry. The framework of NIS may be extended to account for the problem of emergent symmetry identification if the optimized objective, EI, is replaced with some measure of isotope or more general symmetry Liu and Tegmark (2022). Similarly, we can also find other emergent symmetry with this framework once the symmetry can be measured and optimized.
5.1.4 Designing Emergence
We even can design a set of interaction rules for a complex system such that the desired property can emerge from the interactions once the desired property can be defined in a clear and quantitative way. The basic idea to realize emergence design is by optimization Adamatzky et al. ; Kreyssig and Dittrich (2011). Actually, the early studies of evolutionary computation or genetic algorithm have solved this problem for some simple and small cases Adamatzky et al. . For example, M. Mitchell et al. used genetic algorithm to search the interaction rules of one dimensional cellular automata such that the automata can solve some simple mathematical problems Mitchell et al. (1993). Recently, as the development of machine learning and automatic differential, searching such interaction rules may become easier and more efficient if the rules can be converted into differentiable operations Mordvintsev et al. (2022); Gala et al. (2023); Zhang and Walker (2022).
However, if the space of possible interaction rules is very large, all the methods based on search or optimization may be invalid. Thus, some ideas developed in the field of deep learning can be borrowed here to design emergence. For example, the framework of generative adversarial network (GAN) can be used to design complex systems to realize some emergent properties Goodfellow et al. (2014). On one side, we should design a generative model for a complex system or the interaction rules for the system such that a large number of candidate rules or systems can be generated. On the other side, a discriminator or a critic can be designed to evaluate the candidates if the desired emergent properties have been realized. Similar ideas can be used on other possible generative models, e.g., diffusion model, normalizing flow Cao et al. (2022); Yang et al. (2022); Papamakarios et al. (2021); Kobyzev et al. (2020), etc.
However, the difficulty of these methods lies in the demand for a large amount of training data and computations. The lack of data may hamper the applications. In general, designing emergence is a useful problem but it is also difficult.
5.2 Understanding Complex Systems from Causal Emergence
Deep understanding of causal emergence and emergent causality will help us to understand the hidden mechanism of complex systems including living systems, social systems, climate systems, ecosystems, etc. Particularly, some big problems like free will, consciousness Tononi and Koch (2015), and life are all related to emergent causation Walker (2014). For example, free will can be understood as an emergent downward causation Yurchenko (2023). In Chinese medicine, Yin and Yang can be understood as macro-states and the properties of their interactions are emergent causality. Social phenomena can be understood from the viewpoint of causal emergence Elder-Vass (2010). For instance, the increase of the rate of suicide may be the emergent causal effect of economic crises. Interestingly, both EI and $\phi$ID frameworks on causal emergence have relation with one of the consciousness theory: integrated information theory Tononi et al. (2016). As a result, further studies of causal emergence and the connections with other big problems are required.
For a particular system, there are three problems that should be addressed in future studies: 1. When does causal emergence occurs? 2. How emergent causality take functional effects on the system? 3. How emergent causality may change when the system is changed to adapt the environment? We will discuss these problems one by one in brief.
We still do not know when will causal emergence occurs and how does the measure of causal emergence change with some key parameters of a system. In Yang et al. (2023), the authors have shown how the measure of causal emergence change with different noises in an extended Boid model. It increases as the observational noise increases because coarse-graining operation can eliminate the observational noise. However, the same coarse-graining operation can hardly reduce the intrinsic noise (the noise within the dynamical equations), which leads to the degree of causal emergence decreases. It is reasonable to expect that there is a phase transition of the emergent causality within a complex system when some key parameter changes because the causality or the strength of causal effect is also a global property and it may dependent on some order parameters.
If causal emergence occurs in the system, how does the emergent causation take effects on the part and the whole of the system? For example, the emergent "I" can be understood as an emergent macro-variable Yurchenko (2023). How does "I" can influence the part, say my foot, to move? This problem is non-trivial because it relates to the problem of mind-body interaction. One of the difficulty is that how can we build a model for such systems? Should we directly model the emergent properties and causation or model the underlying interaction rules and ignore the causation in higher level? If we directly model the emergent "I", the model may become very complex because the properties of "I" are difficult. However, if we model the system from the micro-level interactions, we can still hardly understand the emergent causality. Shall we switch the viewpoints between micro- and macro-? If yes, in what manner to switch?
Finally, what is the relationship between adaptation and emergence Holland (1995, 2000, 1992)? Sometimes, when we refer to a property as emergent, we actually mean that this property can be developed through adaptation. Therefore, adaptation or evolution serves as a causal force for certain emergent properties. This concept also applies to emergent causality. For instance, downward causation, which is commonly found in complex systems, emerges as a result of adaptation and evolution. The next crucial question is: how can we evolve an emergent property or causation? This problem resembles the issue of designing emergence discussed in Section  5.1.4. However, our aim here is to seek an explanation rather than a design. We want to understand the specific environment and manner in which the observed emergent causation can be evolved through adaptation.
All these problems desire further studies in future. However, these problems are just name a few, more interesting problems need to be addressed.
6 Conclusion
Quantifying the notion of emergence has been a challenge for a long time. Causal emergence serves as an important step in the development of a quantitative theory of emergence, making it a crucial concept in the field of complex system studies. In this paper, we provide a systematic review of papers on causal emergence and other related quantitative studies pertaining to emergence and causality. Our analysis highlights two crucial problem areas: emergence quantification and causal emergence identification in data. We discovered intriguing connections between the latter problem and artificial intelligence, particularly in sub-fields such as causal representation learning and world-model based reinforcement learning. As artificial intelligence continues to advance, we anticipate that more puzzles related to causality and cross-level interactions within complex systems will be solved.
7 Appendix A - Calculation of EI with TPM
Calculating effective information ($EI$) requires knowledge of the joint probability distribution of the system. The state transition probability matrix (TPM) defines the probability of transitioning from one state of the system to another. Specifically, the $j_{th}$ element of the $i_{th}$ row in the TPM, denoted by $TPM(i,j)$, represents the probability of the system being in state $s_{j}$ at time $t+1$, given that it is in state $s_{i}$ at time $t$ (i.e., $P(X_{t+1}=s_{j}|X_{t}=s_{i}))$.
$TPM(i,j)=P(X_{t+1}=s_{j}|X_{t}=s_{i})\\ =\frac{P(X_{t+1}=s_{j},X_{t}=s_{i})}{P(X_{t}=s_{i})}$ (56)
Assuming we can "force" the system to conform to a maximum entropy distribution (i.e., uniform distribution) with $n$ states, the probability of being in any given state $i$ at time $t$, denoted by $P_{u}(X_{t}=s_{i})$, is $1\over{n}$.
As a result, the TPM can be expressed as the product of the joint probability of the system’s maximum entropy distribution, denoted by $P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})$, and the total number of states, which is $n$, as demonstrated below.
$TPM(i,j)=\frac{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}{P_{u}(X_{t}=s_{i})}\\ =\frac{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}{1/n}\\ =n\times{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}$ (57)
Hence, the TPM of the system can be used to express both the joint probability and marginal probability at time $t+1$ of the system’s maximum entropy distribution.
[…]
Using this relationship, $EI$ can be calculated directly with the TPM.
[…]
Reference

### 25 Feb 2024 v3

@@
-Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model
+causality; emergence; causal emergence; causal emergence identification; effective information; machine learning
 Abstract
@@
 On the other hand, causality can exhibit emergence, meaning that new causal laws may arise as we increase the level of abstraction.
-Causal emergence theory aims to bridge these two concepts and even employs measures of causality to quantify emergence.
-This paper provides a comprehensive review of recent advancements in quantitative theories and applications of causal emergence.
-Two key problems are addressed: quantifying causal emergence and identifying it in data.
-Addressing the latter requires the use of machine learning techniques, thus establishing a connection between causal emergence and artificial intelligence.
-We highlighted that the architectures used for identifying causal emergence are shared by causal representation learning, causal model abstraction, and world model-based reinforcement learning.
-Consequently, progress in any of these areas can benefit the others.
-Potential applications and future perspectives are also discussed in the final section of the review.
+Causal emergence (CE) theory aims to bridge these two concepts and even employs measures of causality to quantify emergence.
+This paper provides a comprehensive review of recent advancements in quantitative theories and applications of CE.
+It focuses on two primary challenges: quantifying CE and identifying it from data.
+The latter task requires the integration of machine learning and neural network techniques, establishing a significant link between causal emergence and machine learning.
+We highlight two problem categories: CE with machine learning and CE for machine learning, both of which emphasize the crucial role of effective information (EI) as a measure of causal emergence.
+The final section of this review explores potential applications and provides insights into future perspectives.
 keywords:
-Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model
+causality; emergence; causal emergence; causal emergence identification; effective information; machine learning
 […]
@@
 Moreover, they all exhibit emergent phenomena, which are highly common in complex systems but carry a sense of mystery Holland (2000).
-How did the first living cell emerge from the collisions between various molecules in the early Earth’s environment Artime and De Domenico (2022)?
-How does the cognitive concept of "I" emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)?
+How did the first living cell emerge from the collisions between various molecules in the Earth’s early environment Artime and De Domenico (2022)?
+How does the cognitive concept of “I” emerge from the intricate interactions among countless neurons in our brain Lagercrantz and Changeux (2009)?
 How do large neural language models suddenly exhibit emergent abilities Wei et al.
@@
 These fundamental questions revolve around the concept of emergence in complex living, cognitive, and artificial systems.
-Emergence refers to the phenomenon where macroscopic properties and phenomena that cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000).
-This presents a formidable challenge to the traditional reductionist perspective, while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
+Emergence refers to the phenomenon where macroscopic properties and phenomena cannot be solely attributed to or explained by the properties of individual components Anderson (1972); Meehl and Sellars (1956); Bedau (1997); Holland (2000).
+This presents a formidable challenge to the traditional reductionist perspective while also shedding light on the underlying reasons for the enigmatic nature of emergent phenomena.
 However, as elucidated by Bedau’s theory of weak emergence Bedau (1997), many emergent phenomena can be comprehended through the interactions among the individuals within the system Holland (2000).
@@
 These phenomena demonstrate that complex systems achieve interactions through causal laws, where numerous local causal laws form interconnected causal networks as a whole.
-And this whole possesses its unique causal characteristics.
+And this whole possesses unique causal characteristics.
 Causality, or causation, refers to the connection between a cause and its resulting effect Granger (1969); Pearl et al.
@@
 Emergence and causation interconnect with each other.
-On the one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956).
+On one hand, emergence is the causal effect of complex and nonlinear interactions between components in complex systems Anderson (1972); Meehl and Sellars (1956).
 On the other hand, emergent properties may have causal effects on individuals in complex systems Kim (1992); Holland (2000).
-For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market, while the price may also give feedback to the market: it may affect the decision making of each individual.
+For example, the price of fossil fuel is the emerging result of the interactions between buyers and sellers in the market.
+At the same time, the price may also provide feedback to the market: it can affect the decision making of each individual.
 Furthermore, we can understand emergence through the perspective of causation.
-What does emergence mean is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005).
-Thus, emergent properties or phenomena lose their direct explanations as usual, but may be attributed to the causes on the macro-level as pointed out by Hoel et al.
+What emergence means is that some phenomena and properties on the macroscopic level can not be attributed to the microscopic properties Fromm (2005).
+Thus, emergent properties or phenomena lose their direct explanations as usual but may be attributed to the causes on the macro-level, as pointed out by Hoel et al.
 (2013).
@@
 Specifically, the development of a quantitative theory of emergence is on the verge of emerging.
-Such a theory holds the potential to address significant challenges, including the origins of life, the emergence of novel capabilities in large neural network models, and the potential for intelligence, consciousness, and free will to arise in artificial systems.
+Such a theory holds the potential to address significant challenges, including the origins of life Ellis (2023), the emergence of novel capabilities in large neural network models Wei et al.
+(2022), and the potential for intelligence, consciousness, and free will to arise in artificial systems Oizumi et al.
+(2014).
 Causality not only exhibits a profound connection with emergence but is also considered by many researchers as one of the most crucial perspectives for quantitatively comprehending emergence Hoel et al.
@@
 Two primary challenges take precedence in understanding emergence from a causal perspective.
-The first is establishing a quantitative definition of emergence, while the second involves identifying emergent behaviors or phenomena through data analysis.
+The first is establishing a quantitative definition of emergence, whereas the second involves identifying emergent behaviors or phenomena through data analysis.
 To address the first challenge, two prominent quantitative theories of emergence have emerged in the past decade.
-The first is Erik Hoel’s theory of causal emergence Hoel et al.
-(2013), while the second is Rosas’ theory of emergence based on partial information decomposition Rosas et al.
+The first is Erik Hoel et al.’s theory of causal emergence Hoel et al.
+(2013), whereas the second is Fernando E.
+Rosas et al.’s theory of emergence based on partial information decomposition Rosas et al.
 (2020).
-Hoel’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains.
+Hoel et al.’s theory of causal emergence specifically addresses complex systems that are modeled using Markov chains.
 It employs the concept of effective information (EI) to quantify the extent of causal influence within Markov chains and enables comparisons of EI values across different scales Hoel et al.
 (2013); Hoel (2017).
-Causal emergence is defined by the difference of EI values between macro-level and micro-level.
-Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than the micro-level.
-Hoel and other researchers further extended measures of effective information and causal emergence for dynamical systems with continuous variables in Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020).
-Other measures are also possible to quantify causal effects and thereafter causal emergence.
-In Comolatti and Hoel (2022), Hoel systematically compared a bunch of causal effect measures and concluded that causal emergence is independent on the selection of the measure.
-However, it is essential to establish a coarse-graining strategy beforehand in Hoel’s theory of causal emergence.
+Causal emergence is defined by the difference in the EI values between the macro-level and micro-level.
+Several examples of discrete Markov chains have been shown to exhibit causal emergence if their EI values of macro-level dynamics are larger than those of micro-level dynamics.
+Hoel and other researchers further extended the measures of effective information and causal emergence for dynamical systems with continuous variables Chvykov and Hoel (2020) and complex networks Klein and Hoel (2020).
+Other measures are also possible for quantifying causal effects and, consequently, causal emergence.
+In Comolatti and Hoel (2022), Comolatti and Hoel systematically compared several causal effect measures and concluded that causal emergence is independent of the selection of the measure.
+However, in Hoel’s theory of causal emergence, it is essential to establish a coarse-graining strategy beforehand.
 Alternatively, the strategy can be derived by maximizing the effective information (EI) Hoel et al.
@@
 However, this task becomes challenging for large-scale systems due to the computational complexity involved.
-To address these problems, Rosas introduced a new quantitative definition of causal emergence Rosas et al.
-(2020) that does not depend on coarse-graining methods, drawing from Partial Information Decomposition (PID) related theory.
+To address these problems, Rosas et al.
+introduced a new quantitative definition of causal emergence Rosas et al.
+(2020) that does not depend on coarse-graining methods, drawing from partial information decomposition (PID)-related theory.
 PID is an approach developed by Williams et al., which seeks to decompose the mutual information between a target and source variables into non-overlapping information atoms: unique, redundant, and synergistic information Williams and Beer (2010).
@@
 (2019).
-This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between source and target variables based on the inherent characteristics of the system.
+This framework provides a quantitative definition of causal emergence by measuring the positive synergy information between the source and target variables based on the inherent characteristics of the system.
 The second challenge pertains to the identification of emergence from data.
-In an effort to tackle this issue, Rosas et al.
-have derived a numerical method Rosas et al.
+In an effort to address this issue, Rosas et al.
+derived a numerical method Rosas et al.
 (2020).
 However, it is important to acknowledge that this method offers only a sufficient condition for emergence and is an approximate approach.
-Another limitation is a coarse-grained macro-state variable should be given beforehand to apply this method.
+Another limitation is that a coarse-grained macro-state variable should be given beforehand to apply this method.
 Hence, there is a need for the development of new methods.
+Recently, artificial intelligence, propelled by the rapid advancements in machine learning and deep neural network technology, has witnessed significant progress.
+In the context of causal emergence, there are two key aspects to consider.
+Firstly, machine learning and neural network technology can be employed to address the challenge of identifying causal emergence.
+By leveraging these tools, we can develop approaches to effectively detect and analyze causal emergence phenomena.
+Secondly, the concepts and techniques from causal emergence can be introduced into machine learning to enhance the generalization capabilities of models.
+This integration can potentially improve the ability of machine learning algorithms to generalize well beyond the training data, leading to more robust and adaptable systems.
 In a recent study by Zhang et al.
-Zhang and Liu (2022), a machine learning framework named Neural Information Squeezer (NIS) was introduced to tackle the challenge of identifying causal emergence using Hoel’s definition.
-Remarkably, the NIS neural network, functioning as a "machine observer" equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data.
-This aligns with the viewpoint suggested by Eric Bonabeau and Jean-Louis Dessalles Bonabeau and Dessalles (2011) that all emergent properties and phenomena are observer-dependent, and the epistemological nature of emergent causal powers, including downward causation Kim (1992); Crutchfield and Young (1989).
-However, employing a machine observer offers additional advantages in identifying emergence from data, as it helps avoid subjective judgments.
-Furthermore, we find that causal emergence identification and the recent development of artificial intelligence can have deeper connections through the bridge of NIS and its extended frameworks Yang et al.
+Zhang and Liu (2022), a machine learning framework named the Neural Information Squeezer (NIS) was introduced to address the challenge of identifying causal emergence using Hoel et al.’s framework.
+Remarkably, the NIS neural network, functioning as a “machine observer” equipped with an internal model, exhibits a remarkable ability to identify causal emergence across various types of data.
+In the latest updated version of this work, the Neural Information Squeezer Plus (NIS+) has been developed to directly maximize the critical measure of causal emergence theory, namely effective information (EI) Yang et al.
 (2023).
-First, the architecture of NIS, that is, the “encoder - dynamics learner - decoder” framework, is similar to causal representation learning Schölkopf et al.
-(2021) in artificial intelligence where dynamics learner is replaced as causal mechanisms.
-Therefore, there is a deep connection between the two emerging fields: causal emergence and causal representation learning.
-The former tries to characterise and understand the emergent causality in a complex system and the latter tries to learn the latent causal structures and mechanisms from observational data.
-They are connected by the problem of causal emergence identification.
-That is to identify causal emergence from data is essentially equivalent to learning the latent causality inherent in the data.
-Second, there are more connections between causal emergence and other areas in artificial intelligence.
-For example, causal model abstraction also discussed the interplay between causality and the abstraction (coarse-graining) of causal models Beckers et al.
+Through extensive experiments conducted on both simulated data and real brain data, the NIS+ has demonstrated its ability to automatically find emergent macro-variables and macro-dynamics.
+Consequently, the NIS+ enables quantifying causal emergence in data with the learned macro-dynamics.
+The results of these experiments highlight the effectiveness and potential of the NIS+ in capturing and analyzing causal emergence phenomena.
+Furthermore, the NIS+ showcases superior performance in terms of generalization ability by directly maximizing effective information (EI).
+This brings forth a second question: can we leverage the measure of causation, EI, in the context of causal emergence, to enhance the generalization capability of neural networks for out-of-distribution scenarios?
+This concept is referred to as causal emergence for machine learning.
+By exploring this idea, we aim to bridge the gap between causal emergence and machine learning, potentially unlocking new avenues for improving the generalization abilities of machine learning.
+Finally, in Section 5, we address several important and related issues.
+Firstly, we explore the similarities and differences between two emerging fields: causal emergence and causal representation learning Schölkopf et al.
 (2021).
-In reinforcement learning based on world models Ha and Schmidhuber (2018), the similar framework, “encoder - dynamics learner - decoder”, is also used.
-Thus, the similarity and the shared features between the two fields can help us to lend the ideas and techniques in one field to the other field.
-For instance, EI may be used to measure the causal effect of neural networks Marrow et al.
-(2020), and the techniques in causal representation learning can be employed to detect and identify causal emergence.
-Zhang and Liu (2022).
-This paper tries to review the latest works on the quantitative theory and applications of causal emergence, and its relative works.
-We also discuss its connections with artificial intelligence and complex systems.
-In the second part, we introduce the background of causal emergence, particularly on the connections between causation and emergence in complex systems.
-The third part introduces several quantitative theoretic frameworks including: Crutchfield et al’s computational mechanics Crutchfield and Young (1989), Seth et al’s granger causal emergence Seth (2008), Hoel et al’s causal emergence theory and Rosas et al’ theory of emergence based on information decomposition.
-Several related conceptions are discussed, such as coarse-graining strategy, measure of effective information, and partial information decomposition.
-We also compare these two theories from several aspects.
-The forth part discusses the problem of causal emergence identification and then introduces the connection with artificial intelligence.
-Causal representation learning Schölkopf et al.
-(2021), abstraction of causal models Beckers et al.
-(2020), and world models in reinforcement learning Ha and Schmidhuber (2018) are introduced and the relations with causal emergence identification are discussed.
-Finally, we discuss the connections between causal emergence and other important topics in complex systems such as the origins of life, consciousness, etc.
+This comparison sheds light on the interplay between these two domains.
+Secondly, we delve into a philosophical problem concerning ontological or epistemological causality and emergence, providing insights into the underlying philosophical implications.
+Lastly, we discuss the potential applications of causal emergence in complex systems and how it contributes to our understanding of complex systems from a causal emergence perspective.
+These discussions broaden the scope of this paper and offer intriguing avenues for future research.
+This paper aims to provide a comprehensive review of the latest research on the quantitative theory and applications of causal emergence and related works.
+It also explores the connections between causal emergence, machine learning, and complex systems.
+The subsequent section delves into the background of causal emergence, with a particular focus on the interplay between causation and emergence in complex systems.
+In Section 3, various quantitative theoretical frameworks are introduced, including Crutchfield et al.’s computational mechanics Crutchfield and Young (1989), Seth et al.’s Granger causal emergence Seth (2008), Hoel et al.’s causal emergence theory, and Rosas et al.’s theory of emergence based on information decomposition.
+Additionally, related concepts such as the coarse-graining strategy, measures of effective information, and partial information decomposition are discussed, and a comparative analysis of these theories is presented.
+Section 4 addresses the connection between causal emergence theory and machine learning.
+It explores the use of machine learning and neural network techniques for identifying causal emergence and extends the measure of effective information (EI) to machine learning problems.
+Finally, this paper delves into other important topics and potential applications in the fields of machine learning and complex systems.
 2 Background: From Causality to Emergence
@@
 The relationship between cause and effect can exhibit either deterministic or probabilistic characteristics.
-In deterministic causality, the cause will always produce the same effect, while in probabilistic causality, the cause will only produce the effect with a certain probability.
+In deterministic causality, the cause will always produce the same effect, whereas in probabilistic causality, the cause will only produce the effect with a certain probability.
 In most physical systems governed by differential equations or Markovian dynamics, whether deterministic or probabilistic, causality is inherent.
 This arises from the fact that manipulating a variable within the system can result in observable changes in other variables, albeit with probabilistic outcomes.
-While, one physical mechanism may correspond several possible causal structures as pointed by Y.
+As pointed out by Y.
 Iwasaki and H.
 A.
-Simon Iwasaki and Simon (1994) because the causal graphs describe local causal interactions while mechanisms are the global constraints for all variables.
-Figure 1: The causal models (left column) and the corresponding causal hierarchy (right column) and .
-However, if we do not know the physical mechanism of a system, how can we model the causality behind the system?
-There are several parallel theoretical frameworks to do this.
-For example, Judea Pearl tries probabilistic graphic models (e.g.
-Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al.
+Simon Iwasaki and Simon (1994), one physical mechanism may correspond to several possible causal structures because causal graphs describe local causal interactions, whereas mechanisms are the global constraints for all variables.
+We further discuss the relationship between Markovian dynamics and causal models in detail below.
+If we do not know the physical mechanism of a system, how can we model the causality behind the system?
+There are several parallel theoretical frameworks that can be used to do this.
+For example, Judea Pearl utilized probabilistic graphic models (e.g., Bayesian network, causal graph, and structural causal model) to characterize causal interactions Pearl et al.
 (2000); Pearl (2009).
-In which nodes represent random variables, acyclic links represent causal interactions.
-Three different causal hierarchies are distinguished and quantified in different models.
-The first level of causality is association, i.e., the correlation between variables, which can be modelled by Bayesian networks Sucar (2015).
-The second level is intervention, which can be characterised by causal graph models Pearl (2009).
-The difference between a causal graph and a Bayesian network is the former allows defining a special operator which is called "do" operator to simulate the interventions done by the experimenter.
-The last level is courterfactual, which can be quantified by structural causal models Pearl et al.
+In these models, nodes represent random variables and acyclic links represent causal interactions.
+Pearl distinguished and quantified three levels of causality in different models.
+The first level is association, i.e., the correlation between variables, which can be modeled using Bayesian networks Sucar (2015).
+The second level is intervention, which can be characterized using causal graph models Pearl (2009).
+The difference between a causal graph and a Bayesian network is that the former enables defining a special operator, called a “do” operator, to simulate the interventions performed by the experimenter.
+The last level is counterfactual, which can be quantified using structural causal models Pearl et al.
 (2000).
 The distinguishing characteristic of a structural causal model is that all interactions within a causal graph can be described by a set of deterministic functions that involve the input variables and unknown components, thereby introducing uncertainty.
-Counterfactual causal inferences, that is the inferences under an imaginary scenario, can be conducted only on structural causal models.
+Counterfactual causal inferences, that is, inferences under an imaginary scenario, can be conducted only on structural causal models.
 The different types of causal models and the corresponding causal hierarchy in Pearl’s theory are shown in Figure 1.
-Another quantitative framework for causal inference is Donald B Rubin’s "potential outcome" theory which is directly defined on the level of counterfactual Rubin (2005).
-According to this theory, each individual has two potential outcomes: the outcome that would be observed when individuals receive the treatment and the outcome that would be observed when they do not receive the treatment.
+Figure 1: The causal models (left) column and the corresponding causal hierarchy (right) column.
+Another quantitative framework for causal inference is Donald B Rubin’s “potential outcome” theory, which operates at the counterfactual level Rubin (2005).
+According to this theory, each individual has two potential outcomes: the outcome that would be observed if the individual receives the treatment, and the outcome that would be observed if the individual does not receive the treatment.
 Rubin’s theory suggests that the causal effect of a treatment can be estimated by comparing the difference between the observed outcome and the unobserved potential outcome for each individual.
-Researchers also developed a bundle of methods, which are named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data.
+Researchers have also developed a bundle of methods, named causal discovery algorithms Malinsky and Danks (2018), to automatically discover the causal relations between variables merely from data.
 For example, constraint-based algorithms Spirtes et al.
 (2000) and score-based algorithms Chickering (2002) for causal discovery are two typical representations.
-Causal discovery has become a hot interdisciplinary area in the middle of machine learning, causal inference, and statistics.
+Causal discovery has emerged as an interdisciplinary area at the intersection of machine learning, causal inference, and statistics.
 Due to the uncertainty and ambiguity behind the discovered causal relations, measuring the degree of causal effect between two variables is another important problem.
-Numerous independent historical studies have tackled the issue of causation measurement.
-These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009), etc.
-The effective information (EI) in the framework of causal emergence is also an indicator to measure causal effect.
-All these measures can almost be decomposed into two terms which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
-In summary, causality is a fascinating inter-disciplinary topic across for several subjects. Quantitative framework and measures have been developed in last decades.
+Numerous independent historical studies have addressed the issue of causation measurement.
+These include Hume’s concept of constant conjunction Hume (2016), Eells’ and Suppes’ measures of causation as probability raising Eells (1991); Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009).
+Effective information (EI) in the framework for causal emergence is also an indicator for measuring causal effect, and we explain this point in Section 3.5.3.
+All these measures can be roughly decomposed into two causal primitives, which characterize the necessary causation and sufficient causation, respectively Hoel and Levin (2020).
+It is crucial to differentiate between Markovian dynamics or dynamical mechanism models and the causal models introduced above.
+In current causal emergence theories Hoel et al.
+(2013); Rosas et al.
+(2020), Markovian dynamics, rather than causal models, including causal graphs and structural causal models, play a central role.
+There are several distinctions between Markovian dynamics and these causal models.
+Firstly, in Markovian dynamics, the temporal evolution of one or multiple variables is typically described, with causal relationships between these variables extending across different points in time.
+Consequently, all variables are dependent on time, and a variable at one time step can causally impact another variable at the subsequent time step.
+Conversely, the aforementioned causal models do not explicitly incorporate the element of time.
+In these models, neither the variables nor the causal relationships are explicitly defined as functions of time.
+Secondly, it is worth noting that circular interaction structures, including self-looped interactions, may exist in Markovian dynamics.
+However, such circular structures are typically not permitted in the causal models we introduced.
+This distinction arises from the fact that both causal graph models and structural causal models are constructed based on directed acyclic graphs (DAGs) Pearl (2009).
+However, this issue can be resolved by converting Markovian dynamics into DAGs.
+The initial step involves introducing static causal variables at each time step.
+Consequently, the same variable in Markovian dynamics can be transformed into a set of different variables at different time steps.
+This expansion allows for causal relationships across time steps while eliminating circular structures in causal graphs Koller and Friedman (2009).
+Given the limitations of conventional causal models based on directed acyclic graphs (DAGs) Textor et al.
+(2016), there is a need for new representation methods.
+For instance, Richardson Richardson (2013) expanded the causal modeling approach to include DAGs.
+Spirtes Spirtes (2013) discussed DAG representations of feedback models and applied them to economic processes.
+Lacerda et al.
+also developed a method utilizing independent components analysis to discover cyclic causal models Lacerda et al.
+(2012).
+Furthermore, Forré and Mooij introduced a novel causal discovery method for nonlinear structural causal models with cycles and latent confounders Forré and Mooij (2018).
+George T.H.
+Ellison proposed a methodology that leverages domain knowledge to extract DAGs from temporal relationships among multiple variables Ellison (2023).
+Causal discovery methods can also be extended to temporal data by segregating various temporal data analysis tasks, such as classification, clustering, and prediction.
+Recently, Gong et al.
+provided a comprehensive review of these techniques in Gong et al.
+(2023).
+These advancements have the potential to broaden the range of available causal models.
+In summary, causality is a fascinating interdisciplinary topic spanning several subjects. Numerous quantitative frameworks and measures have been developed in recent decades.
 2.2 Emergence
-Emergence, known as "the whole is greater than the sum of parts" Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro and macro levels of organization.
+Emergence, known as “the whole is greater than the sum of parts” Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro- and macro-levels of organization.
 Therefore, the scale of a system (micro or macro) and the cross-level interactions must be considered when we discuss emergence.
-This may increase the complexity of study for emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al.
+This may increase the complexity of the study of emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al.
 (2019); Ross et al.
-(1908); Holland (1996)
+(1908); Holland (1996).
+Although many examples of emergence have been pointed out in various fields Holland (1996, 2000), including the herding behaviors of birds Reynolds (1987), collective behaviors of simple computer programs Gardner (1970), and emerging ability of large language models Wei et al.
+(2022), a unifying understanding of this phenomenon does not exist right now.
+People try to understand emergence from different aspects, including self-organization Luhmann (1995), order out of disorder Crutchfield (1994), and causality.
+In this review, we mainly focus on the last point of view.
 Emergence has deep connections with causation Fromm (2005); Bedau (1997); Bedau and Humphreys (2008); Bedau (2002).
@@
 Consequently, all emergent properties and processes at the macro-level can, in principle, be comprehended as the causal effects resulting from interactions between individuals at the micro-level.
-This is what the philosophical notion “supervenience” tries to describe.
-However, it does not mean that all the emergent phenomena on macro-level can be easily attributed to micro-level individuals and their interactions.
+This is what the philosophical notion of “supervenience” tries to describe.
+However, this does not mean that all emergent phenomena at the macro-level can be easily attributed to micro-level individuals and their interactions.
 Two major reasons can account for this.
 The first reason is due to the principle of computational irreducibility proposed by Wolfram Wolfram et al.
-(2002) which claims that although we know the emergent properties are out of a set of simple rules, we can not find a short-cut way in principle to predict the results except to implement the rules.
-For example, complex behaviors (eg.
-herding and aligning) of the Boid model are the results of simple rules, but these behaviors can not be reduced to individual behaviors in a simple way except to simulate it Reynolds (1987).
-The second reason is that emergent behaviors may be determined by other emergent properties on macro-level Hoel et al.
+(2002), which claims that although we know emergent properties arise from a set of simple rules, there is no shortcut, in principle, to predict the results other than implementing the rules.
+For example, the complex behaviors (e.g., herding and aligning) of the Boid model are the results of simple rules, but these behaviors cannot be reduced to individual behaviors simply without simulating them Reynolds (1987).
+The second reason is that emergent behaviors may be determined by other emergent properties at the macro-level Hoel et al.
 (2013); Rosas et al.
 (2020).
-For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because the severe lack of fossil oil may cause inflation.
-Therefore, the cause of an emergent property (rice price) is another emergent macro-level coarse-grained variable and it means new causation on macro-level coarse-grained variables can be observed as pointed by the theory of causal emergence.
-And, interestingly, the causation in different levels must be consistent to explain the micro-level event of the price of rice.
-Otherwise, double causation Merton (1973) fallacy may appear.
-This consistency between different scale causality is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
-In order to better understand emergence, Bedau et al Bedau (1997, 2002) classify emergence into three categories according to the causal interactions between micro and macro, which are nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996) and strong emergence O’Connor (1994); Kim (1992).
-In all these categories, the notion of nominal emergence is least controversial, which can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994).
-For example, the patterns which are composed of many pixels on a screen are nominal emergence.
-We can say nominal emergent patterns as “supervenient” because all the macro-level properties are from individuals.
-While weak emergence refers to macro-level properties or processes that can be derived through interactions between individual components but in a complicated way such that they are hardly reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996).
-For the weak emergent patterns or processes, the causes may come from both micro-level and macro-level, therefore, emergent causation may coexist with micro-level causation.
-There are many debates on strong emergence, which refers to properties or processes on macro-level that can not be reduced to micro-level properties in principle, including the interactions of individuals O’Connor (1994); Kim (1992).
-Thus, the collective behaviors of boids are not strong emergent because they are the results of interactions between Boid model.
+For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because severe shortages of fossil oil can cause inflation.
+Therefore, the cause of one emergent property (rice price) is another emergent macro-level coarse-grained variable.
+This means new causation can be observed between macro-level coarse-grained variables, as pointed out by the theory of causal emergence.
+Interestingly, the causation at different levels must be consistent to explain the micro-level event of fluctuating rice prices.
+Otherwise, a double causation Merton (1973) fallacy may appear.
+This consistency between causality at different scales is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
+In order to better understand emergence, Bedau et al.
+Bedau (1997, 2002) classified emergence into three categories according to the causal interactions between the micro- and macro-levels: nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996), and strong emergence O’Connor (1994); Kim (1992).
+Of these categories, the notion of nominal emergence is the least controversial.
+It can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994).
+For example, pixel patterns on a screen are nominal emergent properties.
+We can consider such nominal emergent patterns as “supervenient” because all macro-level properties derive from individuals.
+Weak emergence refers to macro-level properties or processes that derive through interactions between individual components in a complicated way such that they cannot be easily reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996).
+For weak emergent patterns or processes, the causes may come from both the micro- and macro-levels; therefore, emergent causation may coexist with micro-level causation.
+In the discourse surrounding emergence, weak emergence is generally more widely accepted than strong emergence.
+However, there are some concepts and descriptions that remain somewhat ambiguous Héder and Paksi .
+For instance, in Bedeau’s original definition, a property is considered weakly emergent if and only if it can be derived from micro-dynamics through simulation.
+Nevertheless, the term “simulation” itself can be interpreted in different ways—whether it refers to digital simulations or analog simulations is not explicitly implied.
+Bedeau further elaborated on this concept by stating that if a macro-level property can, in principle, be simulated without requiring an actual simulation, then it can be considered weakly emergent Bedau (1997).
+However, even this explanation is not entirely clear-cut.
+There are more debates on strong emergence, which refers to properties or processes at the macro-level that cannot, in principle, be reduced to micro-level properties, including the interactions of individuals O’Connor (1994); Kim (1992).
+Thus, Boid’s collective behaviors are not strongly emergent because they are the result of interactions within the Boid model.
 This notion is controversial because it rejects any mechanistic explanations for strong emergent properties, which cannot be explained by micro-level variables.
-Second it raises some long standing philosophical debates about “causal fundamentalism”  Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
-Jochen Fromm further explains strong emergence as the causal effect of downward causation Fromm (2005).
-Consider a system contains three-different scales: micro, meso, and macro.
-Downward causation refers to the causal power from the macro-level to the messo-level or from the messo-level to the micro-level.
-Thereafter, although the strong emergent properties or processes on the meso-level are supervenient to micro-level properties and interactions, their causes are from the macro-level and therefore a second order patterns or processes.
-However, there are many debates on the notion of downward causation itself, e.g.
-Yurchenko (2023); Bedau and Humphreys (2008).
-Figure 2: Various types of causation that exist within multi-scale complex systems.
+Furthermore, it raises some long-standing philosophical debates about “causal fundamentalism” Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
+Jochen Fromm further explained strong emergence as the causal effect of downward causation Fromm (2005).
+Consider a system that contains three different scales: micro, meso, and macro.
+Downward causation refers to the causal power from the macro-level to the meso-level or from the meso-level to the micro-level.
+Consequently, although the strong emergent properties or processes at the meso-level are supervenient to micro-level properties and interactions, their causes derive from the macro-level and, therefore, second-order patterns or processes.
+However, there are many debates on the notion of downward causation itself, e.g., Yurchenko (2023); Bedau and Humphreys (2008).
+In summary, there are three types of causation based on their cross-level properties, which are associated with emergence, as depicted in Figure 2:
+  1. 1.
+    Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
+  2. 2.
+    Intra-level causation: This refers to causal effects occurring within the same scale or level.
+  3. 3.
+    Downward causation: Here, macro-level properties influence micro-level properties.
+Figure 2: The various types of causation that exist within multi-scale complex systems.
 In this figure, solid arrows represent common causation, which is generally accepted without much controversy.
-Dotted arrowed lines indicate a form of “causation” driven by supervenience, while dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
-In summary, there are three types of causation based on their cross-level properties, which are associated with emergence, as depicted in Figure 2:
-1. Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
-2. Intra-level causation: This refers to causal effects occurring within the same scale or level.
-3. Downward causation: Here, macro-level properties have influences on micro-level properties.
+Dotted arrowed lines indicate a form of “causation” driven by supervenience, whereas dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
 In Figure 2, the black solid lines represent causation without controversy.
-However, according to a recent paper by Yurchenko  Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality, but rather reasons or explanations.
-Yurchenko introduces the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers.
+However, according to a recent paper by Yurchenko Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality but rather reasons or explanations.
+Yurchenko introduced the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers.
 Instead, they serve as reasons or explanations from the perspective of an observer.
-While, Yurchenko accepts the intra-level causation in the meso- or macro-level as “real” as the ones in micro-level.
-He even coined the term “causal equivalent principle” to represent the idea that all the intra-level causation except cross-leveled ones should be accepted and are equivalent each other.
-However, other scholars think that the emergent causation, the intra-level causation in the meso- or macro- level, depends on observers Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022).
-There are much controversy on emergent cross-leveled causation.
+Yurchenko accepted intra-level causation at the meso- or macro-level as “real”, similar to causation at the micro-level.
+He even coined the term “causal equivalent principle” to represent the idea that all intra-level causations, except those that are cross-level, should be accepted and are equivalent to each other.
+However, other scholars think that compared to the causal relationship at the micro-level, the intra-level causation at the meso- or macro-level is problematic because the coarse-graining strategy is relevant Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022).
+As demonstrated in the numeric example in Section 3.5.4, the causation measured by EI is dependent on the coarse-graining method.
+Hoel et al.
+tried to avoid this problem by introducing EI maximization.
+Indeed, EI maximization serves as an objective standard for selecting coarse-graining and macro-dynamics; however, a unique solution is required.
+We discuss this problem further in Section 5.2.
 3 Quantifying Emergence by Causality
-Since causality and emergence have a strong connection, and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence.
-In this section, we will review several frameworks to quantify multi-scale causality and emergence.
+Since causality and emergence have a strong connection and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence.
+In this section, we review several frameworks used to quantify multi-scale causality and emergence.
 3.1 Early Related Works
-Before the theory of causal emergence was proposed by Hoel et al, there were some related works that have introduced very similar ideas as causal emergence theory.
-For example, Crutchfield et al’s computational mechanics theory considered causal states which are the partitions of the state space and may have good predictions.
-While, Seth et al proposed G-causality to quantify emergence by using Granger’s causality.
-We will introduce them in detail.
+Before the theory of causal emergence was proposed by Hoel et al., some works introduced very similar ideas to causal emergence theory.
+For example, Crutchfield et al.’s computational mechanics theory considered causal states, which are the partitions of the state space and may have good predictions.
+Furthermore, Seth et al.
+proposed G-causality to quantify emergence using Granger’s causality.
+We discuss them in detail below.
 3.2 Computational Mechanics
 The theory of computational mechanics proposed by Crutchfield, Shalizi, and Feldman et al.
-tries to formulate this kind of emergent causation in a quantative framework Crutchfield and Young (1989).
-In some sense, computational mechanics can be understood as the inverse of statistical mechanics because what statistical mechanics do is to derive macro-level consequences from micro-dynamics, and the inverse process done by computational mechanics is to construct a minimal causal model from the observations of a stochastic process that can generate the observed time series.
-Let’s assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$.
-We can divide it into two segments beginning from time step $t$: the history before $t$ which is denoted as $\overleftarrow{s_{t}}$ and the future after $t$ which is denoted as $\overrightarrow{s_{t}}$.
-If the process is stationary, we can drop the denotation of $t$.
-Thus, all the possible histories $\overleftarrow{s}$ form a set which is denoted as $\overleftarrow{S}$ and all the futures $\overrightarrow{s}$ form a set $\overrightarrow{S}$.
+tried to formulate this kind of emergent causation in a quantitative framework Crutchfield and Young (1989).
+In some sense, computational mechanics can be understood as the inverse of statistical mechanics.
+This is because statistical mechanics derives macro-level consequences from micro-dynamics, whereas the inverse process is performed by computational mechanics, which constructs a minimal causal model from the observations of a stochastic process that can generate the observed time series.
+Let us assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$.
+We can divide it into two segments beginning from time step $t$: the history before $t$, denoted as $\overleftarrow{s_{t}}$, and the future after $t$, denoted as $\overrightarrow{s_{t}}$.
+If the process is stationary, we can remove the denotation of $t$.
+Thus, all possible histories $\overleftarrow{s}$ form a set, denoted as $\overleftarrow{S}$, and all futures $\overrightarrow{s}$ form a set, denoted as $\overrightarrow{S}$.
 We aim to establish a model that can reconstruct and predict observed random sequences, with a higher accuracy being desirable.
@@
 To preserve useful information as concisely as possible, we need a coarse-grained mapping that captures the ordered structure in the random sequences, known as patterns Crutchfield (1994).
-We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets which form a set $\mathcal{R}$.
-Any subset in $R\in\mathcal{R}$ is a state.
-We define a function from histories to states, $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$.
-Thus $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
+We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets that form a set $\mathcal{R}$.
+Any subset $R\in\mathcal{R}$ is called a "state".
+We define a function from histories to states as $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$.
+Thus, $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
 For a set of states $\mathcal{R}$, we can measure its simplicity using a complexity metric.
@@
 What kind of state set can achieve the best balance between predictiveness and parsimony?
-We can introduce an equivalent relationship which is called causal equivalence Crutchfield (1994).
+We can introduce an equivalent relationship called causal equivalence Crutchfield (1994).
 Concretely, we say $\overleftarrow{s}$ and $\overleftarrow{s}^{\prime}$ are causally equivalent if and only if:
@@
 This equivalent relationship can partition all the histories into equivalent classes, and they are defined as causal states.
-We denote all the causal states of a history $\overleftarrow{s}$ as: $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}}$ which is a subset of the histories.
-For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that make the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
+We denote all the causal states of a history $\overleftarrow{s}$ as $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}},$ which is a subset of the histories.
+For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that represent the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
 $T^{(s)}_{ij}\equiv\text{P}(\mathcal{S}^{\prime}=\mathcal{S}_{j},\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s|\mathcal{S}=\mathcal{S}_{i}),$
-where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and the notation $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$ in which the first symbol emitted is $s$111we use slightly different symbols compared with reference  Shalizi (2001).
+where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$, in which the first symbol emitted is $s$ (we use slightly different symbols from those in Shalizi (2001)).
 Hence, by combining: $\mathcal{S^{\prime}}=\mathcal{S}_{j}$ and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$, we ensure that all subsequent causal states will have the identical initial emitted symbol $s$.
 We denote the set $\{T^{(s)}_{ij}:s\in\mathcal{A}\}$ by $T$.
-And the definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
-The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$) and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
-Up to now we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories.
-The remaining task is to show that it is in some sense optimal.
+The definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
+The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$), and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
+Up until now, we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories.
+The remaining task is to show that it is, in some sense, optimal.
 It has the important properties of being minimally predictive, maximally statistically complex, and minimally stochastic Shalizi (2001).
-We shall compare the causal state $\mathcal{S}$ with any state $R\in 2^{\overleftarrow{\mathcal{S}}}$, to show that for the purpose of predicting the future, the causal states does a better job, i.e.
-gives more information.
-This property can be formulated as a mathematical theorem which is called the maximal predictability theorem.
-This is stated as if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
+We compare the causal state $\mathcal{S}$ to any state $R\in 2^{\overleftarrow{\mathcal{S}}}$ to show that for the purpose of predicting the future, the causal states do a better job, i.e., they provide more information.
+This property can be formulated as a mathematical theorem, called the maximal predictability theorem.
+It is stated as follows: if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
 $H[{\overrightarrow{\mathcal{S}}}^{L}|R]\geq H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}].$
-Where, $H(\cdot|\cdot)$ is the conditional entropy, $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future.
+where $H(\cdot|\cdot)$ is the conditional entropy, and $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future.
 Thus, the inequality shows that the uncertainty of $\overrightarrow{\mathcal{S}}^{L}$ given the causal state $\mathcal{S}$ is less than all other states.
 After achieving optimal predictiveness, the causal state set remains the one with minimal statistical complexity.
-We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$ which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
+We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$, which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
 Next, we present the minimum statistical complexity theorem: for all prescient rivals $\hat{\mathcal{R}}$,
 $C_{\mu}(\hat{\mathcal{R}})\geq C_{\mu}(\mathcal{S})$ (2)
-We next show that the causal states are minimally stochastic.
-That is to say, compared with other competitors with the same ability of predicting the future, the causal states and their transition dynamics, have the least uncertainty.
-Then we have minimal stochasticity theorem which is expressed as: for all prescient rivals $\hat{\mathcal{R}}$,
+Next, we show that the causal states are minimally stochastic.
+That is to say, compared with other competitors with the same ability to predict the future, the causal states and their transition dynamics have the least uncertainty.
+Then, we have the minimal stochasticity theorem, which is expressed as follows: for all prescient rivals $\hat{\mathcal{R}}$,
 $H[\hat{\mathcal{R}^{\prime}}|\hat{\mathcal{R}}]\geq H[\mathcal{S}^{\prime}|{\mathcal{S}}].$
 where $\mathcal{S}^{\prime}$ and $\hat{\mathcal{R}}^{\prime}$ are the next causal state of the process and the next state, respectively.
-This means causal state and $\epsilon$-machine gives the best intrinsic determinism.
-Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from observed data?
-The authors introduced a hierarchical machine reconstruction algorithm in Crutchfield (1994), details are ignored.
-Although this algorithm may not be applicable in all operational scenarios, the authors have presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
+This means that the causal state and $\epsilon$-machine provide the best intrinsic determinism.
+Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from the observed data?
+The authors of Crutchfield (1994) introduced a hierarchical machine reconstruction algorithm; however, the details are not reiterated here.
+Although this algorithm may not be applicable to all operational scenarios, the authors presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
 It is interesting to compare the theory of computational mechanics with causal emergence.
-Indeed, we can understand all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states.
+Indeed, we can understand that all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states.
 Thus the function $\eta$ that can map a history $\overleftarrow{s}$ to a state $R$ is a possible coarse-graining strategy.
-It is worth to be pointed out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can has at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history.
-Therefore, $\epsilon$ is similar to the notion of effective coarse-graining strategy in reference  Zhang and Liu (2022) (see section 4.1) and the causal transitions $T$ are the corresponding effective macro-level dynamics.
-And the feature of minimal stochasticity characterize the deterministic property of the macro-dynamic.
-This is characterized by effective information (EI) in causal emergence theory.
-Although the clear definition and quantitative theory of emergence are not given, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003).
-In Crutchfield (1994), the authors explain that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer.
-Additionally, they differentiate intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.
+It is worth pointing out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can have at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history.
+Therefore, $\epsilon$ is similar to the notion in the effective coarse-graining strategy in Zhang and Liu (2022) (see Section 3.6.5), and the causal transitions $T$ represent the corresponding effective macro-level dynamics.
+The feature of minimal stochasticity characterizes the deterministic property of the macro-dynamic.
+This property is characterized by effective information (EI) in causal emergence theory.
+Although a clear definition and quantitative theory of emergence were not provided, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003).
+In Crutchfield (1994), the authors explained that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer.
+Additionally, they differentiated intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.
 3.3 G-Emergence Theory
-G-emergence theory proposed by Seth in 2008 Seth (2008) is one of the earliest works on quantitative measure of emergence.
+G-emergence theory, proposed by Seth in 2008 Seth (2008), is one of the earliest works on a quantitative measure of emergence.
 His basic idea is to use nonlinear Granger causality to measure weak emergence in complex systems.
-Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$ along, then $B$ is said to be "Granger-cause" of $A$.
+Granger causality (G-causality) is formally defined as follows: Given two time series $A$ and $B$, if the past values of $B$ can help predict the future values of $A$, beyond what can be forecasted using the past values of $A$, then $B$ is said to be “Granger-cause” of $A$.
 This implies the existence of Granger causality between $A$ and $B$.
@@
 Then, the residuals can be utilized to quantify the extent of the causal effect in G-causality.
-The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals, one is residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other one is the residual of the full prediction model, which shown in Figure 3.
-Besides, in Seth (2008), the author also defined G-autonomous as if the past values of a time series can help to predict its own future values.
+The degree of $B$ being the G-cause of $A$ is quantified by the logarithmic of the ratio of the two variances of residuals.
+One is the residual of the autoregression model of $A$ if all the terms of $B$ are omitted, and the other is the residual of the full prediction model, as shown in Figure 3.
+In addition, in Seth (2008), the author also defined G-autonomous as the ability of past values in a time series to predict its own future values.
 And the degree of G-autonomous can be measured in a similar way as G-causality.
-Figure 3: G-causality and G-autonomous
-With these two basic notions in G-causality, the author defines G-emergence on macro-variables in Figure 4: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if: (i) $M$ is G-autonomous with respect to $\{m\}$ and (ii) $\{m\}$ is G-cause of $M$.
+Figure 3: G-causality and G-autonomous.
+With these two basic notions in G-causality, as shown in Figure 4, the author defined G-emergence on macro-variables as follows: a macro-variable $M$ is G-emergent from a set of micro-variables $\{m\}$ if and only if (i) $M$ is G-autonomous with respect to $\{m\}$, and (ii) $\{m\}$ is G-cause of $M$.
 The degree of G-emergence can also be quantitatively measured by multiplying the degree of G-autonomous of $M$ and the average of G-causes of $\{m\}$ on $M$.
-Figure 4: G-autonomous
-The author tests G-emergence theory on the Boid model.
-This model is a famous artificial life model to simulate the flocking behavior of birds by using three simple rules: cohesion, alignment, and separation Reynolds (1987).
-All these basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled.
-The author found that with the strength of cohesion increased, the G-emergence of the entire flock is also increased.
-Here, the macro-variable is selected as the center of the flock (center-of-mass), and each bird is a micro-variable.
-The author also discovered a downward causality phenomenon in this simple model that is the center of mass can be used to predict each individual bird.
-However, the author does not distinguish downward causation from other common causality in this work because Granger’s causality is not a real causal relationship.
-Seth’s G-emergence theory is the first attempt to quantify emergence phenomenon via causality measure.
-However, the causality measure that the author used is granger causality which is not a strict causal measure and it is also depends on the regression method to be used.
-Furthermore, the measure is defined on variables but not dynamics which means the result depends on the selection of variables.
+Figure 4: G-autonomous.
+The author tested the G-emergence theory on the Boid model.
+This model is a famous artificial life model that simulates the flocking behavior of birds using three simple rules: cohesion, alignment, and separation Reynolds (1987).
+These basic rules are realized by the virtual forces exerted on each bird, and the strength of each force can be controlled.
+The author found that by increasing the strength of cohesion, the G-emergence of the entire flock also increased.
+Here, the macro-variable is selected as the center of the flock (center of mass), and each bird is a micro-variable.
+The author also discovered a downward causality phenomenon in this simple model: the center of mass can be used to predict each individual bird.
+However, the author did not distinguish downward causation from other common causality in their work because Granger’s causality is not a real causal relationship.
+Seth’s G-emergence theory was the first attempt to quantify the emergence phenomenon via a causality measure.
+However, the causality measure that the author used was Granger causality, which is not a strict causal measure, and it also depends on the regression method to be used.
+Furthermore, the measure is defined on variables but not dynamics, which means the result depends on the selection of variables.
 3.4 Other Quantitative Theories of Emergence
 There are alternative means of measuring emergence that do not require relying on causality.
-Two different ways are discussed by the scholars, one is to understand emergence as a process from disorder to order, another one is to understand emergence from the view of "the whole is greater than the sum of parts".
-For example, Moez Mnif and Christian Müller-Schloer used Shannon’s entropy  Mnif and Müller-Schloer (2011) to measure order and disorder.
+Two different methods have been discussed by the scholars.
+One is to understand emergence as a process from disorder to order, and the other is to understand emergence from the perspective that “the whole is greater than the sum of parts”.
+For example, Moez Mnif and Christian Müller-Schloer used Shannon entropy Mnif and Müller-Schloer (2011) to measure order and disorder.
 In a self-organized process, emergence occurs when there is an increase in order.
@@
 $Emergence=H_{max}-H-\Delta H_{view},$ (3)
-where, $H_{max}$ is the maximized entropy of the system.
+where $H_{max}$ is the maximized entropy of the system.
 If no prior information is available, the maximum entropy, denoted as $H_{max}$, corresponds to the Shannon entropy of an equal probability distribution.
@@
 We can also normalize this quantity by dividing $H_{max}$ such that different features can be compared by the normalized quality.
-For a multivariate system, the authors suggest employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables.
-Then the authors apply this method to a simulated system of chickens.
+For a multivariate system, the authors suggested employing a radar plot to visualize the emergence fingerprint of the dynamical process, which showcases the normalized emergence measure across various variables.
+Then, the authors applied this method to a simulated system of chickens.
 M.
 Tang and X.
-Mao applied this indicator on artificial society models  Tang and Mao (2014).
+Mao applied this indicator to artificial society models Tang and Mao (2014).
 Inspired by Moez Mnif and Christian Müller-Schloer’s work,  Fisch et al.
-(2010) suggests to use the divergence measure between two probability distributions to better quantify emergence.
-They understand emergence as an unexpected or unpredictable change of the distribution underlying the observed samples.
-However, this method has the defect of computational complexity and estimation accuracy.
-To conquer the problems,  Fisch et al.
-(2012) further proposed an approximating method by using Gaussian mixture model to estimate the density and introducing Mahalanobis distance to characterize the divergence between data and Gausian component, better results can be obtained.
+(2010) suggested using the divergence measure between two probability distributions to better quantify emergence.
+They understood emergence as being an unexpected or unpredictable change of the distribution underlying the observed samples .
+However, this method suffers from computational complexity and estimation accuracy.
+To address these problems, Fisch et al.
+(2012) further proposed an approximating method using the Gaussian mixture model to estimate the density and introduced Mahalanobis distance to characterize the divergence between data and Gaussian components, leading to better results.
 In Fisch et al.
-(2011), the authors systematically compare the three mentioned methods, and apply them on a simple test example.
-Another Shannon entropy based emergence measure is given by Holzer and de Meer et al.
+(2011), the authors systematically compared the three aforementioned methods and applied them to a simple test example.
+Another Shannon entropy-based emergence measure was proposed by Holzer and de Meer et al.
 Holzer et al.
 (2008); Holzer and de Meer (2011).
-They consider a complex system as a self-organization process, in which different individuals interact with communications.
-Then emergence can be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
-Different from all the mentioned methods above, Teo et al.
-(2013); Szabo and Teo (2015) proposes a method to quantify emergence based on the idea of "the whole is greater than the sum of its parts", and it defines emergence from the interaction rules and states of agents instead of on the overall statistical measure of the whole system.
+They considered a complex system as a self-organization process in which different individuals interact with communications.
+Emergence can then be measured based on a ratio between the Shannon entropy measure on all communications between agents and the total summation of the Shannon entropy for each communication as a separate source.
+Unlike the aforementioned methods, Teo et al.
+(2013); Szabo and Teo (2015) proposed a method to quantify emergence based on the idea that “the whole is greater than the sum of its parts”, defining emergence from the interaction rules and states of agents instead of the overall statistical measure of the whole system.
 Specifically, this measure consists of two terms that are subtracted from each other.
 The first term characterizes the collective states of the entire system, whereas the second term represents the summation of the individual states of all its constituent parts.
-This measure emphasize the emergence that arises from the interactions and collective behavior of the system.
-Then this method is tested on the example of bird flocks simulation.
+This measure emphasizes the emergence that arises from the interactions and collective behavior of the system.
+This method was then tested on the example of bird flock simulation.
 3.5 Erik Hoel’s Causal Emergence Theory
 3.5.1 Basic Idea
-The first quantitative emergence theory defined on Markov dynamics and causality measure based on intervention is Erik Hoel’s causal emergence theory.
+The first quantitative emergence theory based on Markov dynamics and causality measures through intervention was Erik Hoel’s causal emergence theory.
 In this framework, system properties can be characterized at various levels, ranging from micro to macro.
-If a system exhibits stronger causality at the macro level than at the micro level, it demonstrates causal emergence.
+If a system exhibits stronger causality at the macro-level than at the micro-level, it demonstrates causal emergence.
 Causality is reflected between successive states during the system’s evolution.
 The strength of a system’s causality reveals the extent to which its future state is influenced by its current state.
-The basic idea of causal emergence can be illustrated by Figure 5.
-Figure 5: The basic idea of Erik Hoel’s causal emergence theory.
+The basic idea of causal emergence is illustrated in Figure 5.
+Figure 5: The basic idea of Erik Hoel’s causal emergence theory. The colored circles represent micro-states, whereas the colored squares represent macro-states.
 For example, statistical mechanics is a typical theory of causal emergence.
-On the micro-level, a huge number of molecules collide each other and exhibit random behaviors such that probabilistic language must be used to describe them.
-However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure and temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors.
-Therefore, thermodynamic laws have stronger causal effect than micro-level molecular dynamics.
+At the micro-level, a huge number of molecules collide and exhibit random behaviors such that probabilistic language must be used to describe them.
+However, if we coarse-grain the whole system into several thermodynamic physical variables like pressure, temperature, etc., we can use very concise and exact thermodynamic equations to describe their behaviors.
+Therefore, thermodynamic laws have a stronger causal effect than micro-level molecular dynamics.
 Formal tools are used to describe the elements in Hoel’s causal emergence theory.
-Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro systems, and the inherent strength of causal effect of the Markov model can be measured with effective information (EI) which indicates how effectively a particular state influences the future state of a system.
-EI is an intrinsic property of a system’s dynamics which can be quantified by the transitional probability matrix (TPM).
-A coarse-graining strategy is a function that can map a set of micro-states into a particular macro-state, and therefore a new dynamical model which can be described by TPM can naturally be derived from the micro-level TPM, and the effective information of the coarse-grained model can also be computed.
-The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase, and at a certain macroscopic scale of coarse-graining, the effective information reaches a maximum: this is the scale at which the system state has maximum causal power to predict future states in the most reliable and effective way.
-To illustrate the causal emergence, let’s look at a particular Markov chain model with 4 possible states, whose transition probability matrix (TPM) is:
+Typically, it employs discrete Markov models to describe the micro-dynamics of systems, and the corresponding macro-level systems with different Markov dynamics can be derived by coarse-graining the micro-systems.
+Additionally, the inherent strength of the causal effect of the Markov model can be measured with effective information (EI), indicating how effectively a particular state influences the future state of a system.
+EI is an intrinsic property of a system’s dynamics and can be quantified by the transitional probability matrix (TPM).
+A coarse-graining strategy is a function that maps a set of micro-states into a particular macro-state, allowing for the derivation of a new dynamical model described by TPM from the micro-level TPM.
+The effective information of the coarse-grained model can also be computed.
+The phenomenon of causal emergence implies that as we coarse-grain microscopic states, the amount of effective information transmitted from the current state to the next state can possibly increase.
+At a certain macroscopic scale of coarse-graining, the effective information reaches a maximum; this scale represents the point at which the system state has the maximum causal power to predict future states in the most reliable and effective way.
+To illustrate causal emergence, let us look at a particular Markov chain model with four possible states, whose transition probability matrix (TPM) is:
 $S_{m}=\begin{bmatrix}1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 1/3&1/3&1/3&0\\ 0&0&0&1\end{bmatrix}$ (4)
-In this model, if we take the first 3 states as a group 1 and the last state as group 2, anyone in group 1 can transit to all 3 states of the same group with equal probability at the next moment, while because there’s only 1 state in group 2, so the 4th state will always stay in its position.
-Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty is mainly from group 1.
-However, if we merge the first 3 states in group 1 into 1 new state, and keep the 4th state in group 2 as it is, then we get a coarse-grained Markov model with 2 states, with TPM:
-[…]
-Now the future states of the new system can be fully determined by the current states.
-It shows that we can eliminate the uncertainty of an undeterministic system by doing coarse-graining over the system states.
+In this model, if we take the first three states as group 1 and the last state as group 2, anyone in group 1 can transition to any of the three states in the same group with equal probability at the next moment.
+However, because there is only one state in group 2, so the fourth state will always stay in its position.
+Intuitively, we can conclude that the future states are not fully determined by the current states, and uncertainty mainly arises from group 1.
+However, if we merge the first three states in group 1 into one new state and keep the fourth state in group 2 as it is, we obtain a coarse-grained Markov model with two states, represented by the following TPM:
+$S_{M}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$ (5)
+Now, the future states of the new system can be fully determined by the current states.
+This shows that we can eliminate the uncertainty of a nondeterministic system by performing coarse-graining over the system states.
 3.5.2 About Coarse-Graining
@@
 Coarse-graining is a process that simplifies the description of a system by grouping its components into larger, more slowly varying units.
-It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of microscale interactions.
+It is often used to identify the essential features of a system that determine its macroscopic behavior, without being burdened by the details of micro-scale interactions.
 Unlike dimension reduction techniques, such as PCA and SVD, coarse-graining takes into account the system’s features at different spatial and temporal scales.
 Coarse-graining also differs from renormalization.
-In physics, the method of renormalization group is invented to avoid the infinity in integrals, it is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro- levels Christensen and Moloney (2005); McComb (2003).
-However, coarse-graining methohd does not have this requirement in general.
-Although both techniques are targeting to describe a system from a coarse-grained level, however, coarse-graining are more focused on the states of the system, while renormalization cares of the dynamics, the system rules or partition function etc, which is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking.
+In physics, the renormalization group method was invented to eliminate infinities in integrals.
+It is also used to coarse-grain a system such that the Hamiltonian or partition functions are similar between micro- and macro-levels Christensen and Moloney (2005); McComb (2003).
+However, the coarse-graining method does not have this requirement in general.
+Although both techniques are designed to describe a system from a coarse-grained level, coarse-graining focuses more on the states of the system, whereas renormalization cares about dynamics, system rules, partition functions, etc.
+Renormalization is frequently used in physics and the study of phase transitions to explore critical phenomena and symmetry breaking.
 These are aspects that coarse-graining typically does not consider.
-Given a transitional probability matrix, there are a large amount of ways to partition the state space in order to coarse-grain the system.
+Given a transitional probability matrix (TPM), there are many ways to partition the state space for coarse-graining the system.
 What kind of coarse-graining strategies are more reasonable?
 In Hoel et al.
-(2013) and Hoel (2017), the authors suggest to select an appropriate coarse-graining method by maximization effective information of the coarsed TPM.
-In history, there is another criterion to select coarse-graining strategy called lumpability.
-Where lumpability of a TPM refers to the coarse-grained TPM exhibits similar dynamics and observation statistics Kemeny and Snell (1983).
-Commutativity is another requirement which needs the operations of coarse-graining and state transition being commutative Pfante et al.
+(2013); Hoel (2017), the authors suggested selecting an appropriate coarse-graining method by maximizing the effective information of the coarse TPM.
+In the literature, there is another criterion for selecting a coarse-graining strategy called lumpability, where the lumpability of a TPM refers to the coarse-grained TPM exhibiting similar dynamics and observation statistics Kemeny and Snell (1983).
+Commutativity is another requirement that requires the operations of coarse-graining and state transitions to be commutative Pfante et al.
 (2014).
-There are also other discussions or reasonable model reduction of Markov or hidden Markov models Kotsalis et al.
+There are also other discussions or reasonable model reductions of Markov or hidden Markov models Kotsalis et al.
 (2008); White et al.
 (2000); Wolpert et al.
-(2014)
+(2014).
 3.5.3 About Effective Information
 In order to quantify emergence, a widely used method is to calculate the mutual information between the past states and future states of a system.
-This mutual information defines the upper limit of the information being transferred from past to future.
-When the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
+This mutual information defines the upper limit of the information being transferred from the past to the future.
+If the mutual information is high, it suggests that a significant amount of information about the past is retained in the future.
 $I(X_{t};X_{t+1)}=\sum P(X_{t},X_{t+1})\log\frac{P(X_{t},X_{t+1})}{P(X_{t})P(X_{t+1})}$ (6)
 One limitation of utilizing mutual information is that the resulting value may fluctuate in response to changes in the joint probability of data.
-This can make it difficult to obtain a consistent outcome when the system is exposed to different inputs.
-To address this limitation, Hoel introduced the concept of effective information.
-This concept measures the causal effect of the current state of a system on its future state, also based on mutual information.
-Effective information is a scoring metric that is calculated from the system’s transition probability matrix (TPM).
-The TPM does not change solely based on input data.
-Assuming that the current state of the system is in the maximum entropy distribution $I_{D}$ (uniform distribution), the TPM predicts that the state distribution of the system at the next moment (i.e., in the future) will be $E_{D}$.
-The effective information $EI$ of the system is then defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
-$EI=I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{X_{t}\sim U}$ (7)
-$X_{t}\sim U$ represents the probability of the random variable $X_{t}$ following a uniform distribution.
-It can be readily demonstrated that the effective information ($EI$) can be fully described by the transition probability matrix ($TPM$) (see Appendix 7).
-[…]
-Effective information offers a solution to the aforementioned limitation of mutual information.
-Given that TPM can capture the inherent nature of a system, effective information is likewise an intrinsic property of the system.
-Causal emergence involves assuming a uniform distribution for the input data, which can be seen as a way of manipulating the input data to isolate the interference of the original distribution of input data.
-This manipulation is formalized using Judea Pearl’s causation theory, specifically through the "do" operation.
-By imposing this assumption of a uniform distribution, we create a controlled environment with equal probabilities of $X_{t}$’s states and allows us to better understand the causal relationships between $X_{t}$ and $X_{t+1}$.
-Therefore, the term "causal emergence" is used.
-We can calculate EIs for both TPMs of macro- and micro- dynamics, and their difference is defined as the measure of causal emergence, that is:
+This can make it difficult to obtain a consistent outcome if the system is exposed to different inputs.
+Effective Information (EI)
+To address this limitation, Hoel introduced the concept of effective information to measure the causal effect of the current state on the future state of a system.
+Effective information is a scoring metric based on mutual information and can be calculated from the system’s transition probability matrix (TPM), which is invariant to the input data.
+By “intervening” in the current state distribution to follow the uniform distribution (the distribution of maximum entropy), denoted as $I_{D}$, the TPM enables the prediction of the future state distribution $E_{D}$ at the next moment.
+The effective information ($EI$) of the system is defined as the mutual information between $I_{D}$ and $E_{D}$, which can be expressed as
+$EI\equiv I(I_{D};E_{D})=I(X_{t};X_{t+1})|_{do(X_{t})\sim U}$ (7)
+The probability distribution of the random variable $X_{t}$, representing the initial state, is denoted as $do(X_{t})\sim U$, indicating that it follows a uniform distribution.
+Hoel adopted the notion of the “do” operator from Judea Pearl’s causal analysis framework Hoel (2017).
+It is worth noting that in Pearl’s context, the do operator is typically used to assign a specific value to the intervened variable rather than applying a distribution Pearl (2009).
+Another point that needs clarification is that the “do” operator used here is purely a mathematical construct that specifies the distribution of $X_{t}$.
+It does not imply the need for actual intervention in the system.
+However, the imaginary “do” operation is equivalent to a real intervention in the context of this scenario, given that the dynamical mechanism (TPM) is provided.
+Consequently, we can perform any desired intervention on the system, just as in computer simulations.
+The second issue pertains to why we apply the “do” operator to a uniform distribution.
+In Hoel et al.
+(2013), Hoel et al.
+claimed that the distribution should be the maximum entropy distribution, which is the most reasonable selection for the input variable $X_{t}$ if we have no prior information about the input variable Jaynes (1957).
+As we know, uniform distribution can be derived by maximizing entropy if there is no constraint.
+The authors of this review believe that applying the “do” operator to a uniform distribution ensures that the objective measured by EI solely reflects the dynamical mechanism itself, i.e., the TPM, and is independent of any input data.
+This point can be clearer if we re-express EI as a function of the TPM of the system (see Appendix A for the detailed derivation):
+$EI=\frac{1}{N}\sum_{i,j}TPM(i,j)\log_{2}\Big{(}\frac{N\times TPM(i,j)}{\sum_{k}TPM(k,j)}\Big{)},$ (8)
+where $TPM(i,j)$ represents the transitional probability of the system from the state $i$ to the state $j$.
+Therefore, effective information offers a solution to the aforementioned limitation of mutual information.
+Given that the TPM can capture the inherent nature of a system, EI is likewise an intrinsic property of the system.
+However, if we apply the “do” operator to the system using different distributions, EI will depend on the chosen distribution, and certain transition probabilities of specific rows may carry greater weight in the average calculation.
+Consequently, deriving a simple expression solely based on the transitional probability matrix (TPM) may not be possible.
+Furthermore, if there is no “do” operator, the effects of the dynamical mechanism and the input distribution in EI would not be distinguishable or separated, as the mutual information is a function of the input distribution $p(X_{t})$.
+The Derived Measures of EI
+Furthermore, we can calculate the EIs for the TPMs of both macro- and micro-dynamics, and their difference is defined as the measure of causal emergence, that is,
 $CE=EI(TPM_{M})-EI(TPM_{m}),$ (9)
-where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the one of micro-level.
+where $TPM_{M}$ represents the TPM of macro-level dynamics, and $TPM_{m}$ represents the TPM of micro-level dynamics.
 $CE$ measures the degree of causal emergence.
-If $CE>0$, then causal emergence occurs, otherwise, does not occur.
+If $CE>0$, then causal emergence occurs; otherwise, it does not occur.
 However, there is a limitation for EI and CE when comparing two dynamics that differ significantly in size.
 This is because the value of effective information relies on the number of possible states within the system, denoted as $N$, with an upper bound of $\log_{2}(N)$.
-To facilitate the comparisons between different coarse-graining strategies and scales, effective information is often normalized resulting in a metric known as the effect coefficient $Eff$.
+To facilitate comparisons between different coarse-graining strategies and scales, effective information is often normalized, resulting in a metric known as the effect coefficient $Eff$.
 $Eff=\frac{EI}{\log_{2}(N)}$ (10)
-The value of effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states.
+The value of the effect coefficient is always between 00 and $1$, representing the proportion of effective information being transferred from current states to future states.
 If the information is fully transferred, the effect coefficient is $1$.
-In addition to characterizing causal emergence, effect coefficient can be further broken down into two meaningful components: "Determinism," which represents the certainty of a current state evolving into a certain state, or diverging into multiple states in the next moment, and "Degeneracy," which represents the possibility of multiple current states converging into one state in the next moment.
-$Eff=Determinism-Degeneracy.$ (11)
+In addition to characterizing causal emergence, the effect coefficient can be further broken down into two meaningful components: “$Determinism$”, which represents the certainty of a current state evolving into a certain state or diverging into multiple states in the next moment, and “$Degeneracy$”, which represents the possibility of multiple current states converging into one state in the next moment.
+$Eff=Determinism-Degeneracy$ (11)
+where both “$Determinism$” and “$Degeneracy$” can be defined in terms of the TPM (refer to Equations (8) and (10)):
+$Determinism$ $=\frac{1}{N\log_{2}(N)}{\sum_{i,j}{TPM(i,j)\log_{2}(N\times TPM(i,j))}}$ (12)
+$Degeneracy$ $=\frac{1}{N\log_{2}(N)}{\sum_{i,j}TPM(i,j)\log_{2}(\sum_{k}{TPM(k,j)})}$
 It should be noted that the determinism of a system is always greater than its degeneracy, as the lower bound of the effect coefficient is 00.
 The following examples illustrate what determinism and degeneracy look like in systems with varying TPMs.
-Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graning can change them.
-(a) Deterministic and zero degenerative system; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from c or d, resulting in a fully deterministic system with high effect coefficient.
-In Figure 6, the square cells represent for the elements of TPM matrix, and the grey scale represents for the value of TPM elements.
+In Figure 6, the square cells represent the elements of the TPM, and the grayscale areas represent the values of the TPM elements.
 Example (a) is a bijective system, meaning that all information from current states can be transferred to future states without loss.
@@
 Finally, system (e) is a coarse-grained version of either system (c) or (d), demonstrating two important points: first, different microscopic systems can be coarse-grained to the same macroscopic system, and second, causal emergence can be quantitatively captured from coarse-graining by the increased effect coefficient.
-Despite the fact that effective information can identify and quantify causal emergence, as previously mentioned, there are other independently developed measures of causation, such as Humean constant conjunction Hume (2016), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009) etc.
-Hoel found all these measures, including effective information can be related to a small set of "causal primatives" Comolatti and Hoel (2022): the probability representations of sufficiency and necessity (and its variant) of cause $c$ and effect $e$.
-$suff(e,c)=P(e|c)$ (12)
-$nec^{\dagger}(e,c)=P(e|C)$ (13)
-$-\𝑐nec(e,c)=1-P(e|C\verb|\|c)$ (14)
-Where $C$ stands for the set of all causes and $-\𝑐C\verb|\|c$ stands for the set of all causes but $c$.
-Hoel pointed out that determinism is actually a measure of sufficiency, while degeneracy is an inverse measure of necessity.
-The consistency in causal primitives indicates there is a scientific foundation of these individually developed measures of causation.
-One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation 8.
-Therefore, Varley (2022) proposed the conception of flickering emergence to decompose EI by each term being added in Equation 8.
-This new feature can characterize the local properties of a Markov dynamics.
+One limitation of EI is its global nature, as highlighted by the fact that the summation of additional terms encompasses the entire state space, as defined in Equation (8).
+Therefore, Varley (2022) proposed the concept of flickering emergence, which decomposes EI into each term added in Equation (8).
+This new feature can characterize the local properties of Markov dynamics.
+Figure 6: Examples of effect coefficient, determinism, degeneration, and how coarse-graining can change them.
+(a) Deterministic system with zero degeneracy; (b) Deterministic system with zero effect coefficient due to high degeneracy; (c) Partially deterministic system; (d) Deterministic system with low effect coefficient; (e) Coarse-grained system from (c) or (d), resulting in a fully deterministic system with a high effect coefficient.
+Comparison with Other Measures of Causation
+Hoel’s theory of causal emergence is based on effective information, however, is the selection of EI necessary for measuring causation?
+Is causal emergence a phenomenon that depends on the selection of the measure of causation?
+To address this problem, in Comolatti and Hoel’s work Comolatti and Hoel (2022), a systematic comparison was conducted between effective information (EI) and other measures of causation that are widely applied in various fields, ranging from philosophy to genetics.
+The findings of this comparison revealed that causal emergence is not merely a peculiar phenomenon limited to a specific measure.
+Instead, it exhibits commonalities and shared characteristics across different measures and fields of study.
+EI is not the sole measure for capturing causal emergence; there are other measures of causation that can also reveal the phenomenon of causal emergence.
+This suggests that the concept of causal emergence has broader applicability and relevance.
+We introduce this work in detail below.
+Firstly, Hoel highlighted that causation is not merely a singular relationship between a cause and an effect.
+Instead, it encompasses two fundamental dimensions known as causal primitives: sufficiency and necessity.
+The sufficiency aspect of causation refers to the scenario where the occurrence of the cause guarantees the occurrence of the effect.
+In other words, whenever the cause happens, the effect is also observed.
+This sufficiency dimension, denoted as $suff$, can be formally defined as the probability of the effect $e$ occurring given the condition that the cause $c$ has occurred:
+$suff(e,c)=P(e|c).$ (13)
+In addition, a necessary ($nec$) relation in causation refers to the absence of the effect implying the absence of the cause.
+In other words, when the effect does not occur, it indicates that the cause also does not occur.
+This can be understood as a kind of causal effect measure for the counterfactual.
+This necessary dimension, denoted as $nec$, can be quantitatively defined as the probability that the effect $e$ does not occur when the condition that the cause $c$ has not occurred is given by:
+$nec(e,c)=1-P(e|C/c),$ (14)
+where $C/c$ represents the set of all possible causes in $C$ but with the particular cause $c$ being excluded.
+Therefore, $P(e|C/c)$ represents the probability that $e$ occurs if other causes except $c$ occur.
+With these causal primitives defined, Hoel compared different measures of causation, including Hume’s constant conjunction Hume (2016), Cheng’s causal attribution Cheng and Novick (1991), Eells’s measure of causation as probability raising Eells (1991), Suppes’s measure of causation as probability raising Suppes (1973), and Judea Pearl’s measures of causation Pearl (2009).
+The finding was that all of these measures can be expressed as the two causal primitives.
+For example, Cheng’s causal attribution can be expressed as:
+$CS_{cheng}=\frac{suff(e,c)+nec(e,c)-1}{nec(e,c)}$ (15)
+With this understanding, a natural question arises: Can effective information (EI) be expressed using causal primitives?
+The answer is affirmative.
+However, to clarify this point, two important distinctions need to be made.
+Firstly, in most measures of causation, the causal variables are binary, whereas EI is defined for variables with multiple values.
+Secondly, EI is an information-theoretic measure, whereas others are probabilistic measures.
+Despite these distinctions, EI can still be expressed using causal primitives.
+To understand this, let us examine the equivalent measure of EI: normalized effective information ($Eff$) in Equation (11).
+This measure contains two terms.
+The first term is $determinism$, and the second one is $degeneracy$.
+These two terms can be expressed as $suff$ and $nec$:
+$determinism=1-\frac{\sum_{c\in C}P(c)H(e|c)}{\log_{2}N},$ (16)
+where $H(e|c)\equiv-\sum_{e\in E}P(e|c)\log_{2}P(e|c)$ is the Shannon entropy of the conditional probability $P(e|c)$, which is $suff(e,c)$, and $P(c),\forall c\in C$ is the distribution of all causes.
+In the definition of EI (Equation 7), this distribution is intervened as a uniform distribution such that equal weights are assigned to causes $c\in C$.
+It is not hard to see that $determinism$ is an information metric for $suff(e,c)$ and is averaged for all causes.
+Furthermore, as $suff(e,c)=P(e|c)$ increases and approaches one, indicating the emergence of causal effect, the value of $H$ decreases, whereas determinism concurrently increases.
+Thus, the determinism term in EI plays a similar role to that of $suff$ in causal primitives.
+Another term is $degeneracy$, which can be re-written as:
+$degeneracy=1-\frac{H(e|C)}{\log_{2}N},$ (17)
+where $H(e|C)\equiv-\sum_{e\in E}P(e|C)\log_{2}P(e|C)$ is also the Shannon entropy of the conditional probability $P(e|C)$, which is calculated by averaging the causal effect for all elements in $C$ as $P(e|C)=\sum_{c\in C}P(c)P(e|c)$.
+The entropy $H(e|C)$ serves as a measure of the average causal effect for counterfactuals, as $P(e|C)$ can be interpreted as $P(e|C/c)$.
+This is due to the approximation $P(e|C)\approx P(e|C/c)$ when the number of elements in $C$ is significantly large, and $C$ certainly encompasses $C/c$.
+Consequently, $degeneracy$ in $EI$ acts as the counterpart to $nec$ in causal primitives.
+Therefore, EI or $eff$ is a valuable measure of causation, particularly in cases where the cause and effect variables are not limited to binary values.
+Comolatti and Hoel Comolatti and Hoel (2022) conducted additional experiments to investigate the phenomenon of causal emergence using various Markovian dynamics, employing different measures of causation, as discussed above.
+Their findings revealed the widespread occurrence of causal emergence, regardless of the specific measure of causation employed.
 3.5.4 Examples of Causal Emergence
@@
 However, in reality, coarse-graining often occurs among variables, and the combination of these variables produces different states.
-Let’s look at a more complicated example of a boolean network that given in Hoel et al.
-(2013) and we will inspect the causal emergence when coarse-graining its state transition mechanism.
-A, B, C and D are 4 binary variables whose state transational relations are as shown in Figure  7(a).
+Let us look at a more complicated example of a boolean network, given in Hoel et al.
+(2013), and inspect the causal emergence when coarse-graining its state transition mechanism.
+A, B, C, and D are four binary variables whose state transition relations are shown in Figure 7a.
 Figure 7: The mechanism and transition probability matrix of a boolean network.
 Each variable’s next state is determined by the combination of the current states of two other variables, as shown in the boolean network (a).
-If we examine an atomic subsystem with two inputs and one output, such as AB -> C, the next state of C depends on the present states of A and B.
-If A=0 and B=0, then the probability of C being 1 is 70%, while the probability of it being 0 is 30%, as indicated in table (b).
-This same microscopic dynamics applies to other atomic subsystems such as AB -> D, CD -> A, and CD -> B.
-When we expand to a subsystem of two inputs and two outputs, AB -> CD and CD -> AB, we obtain the transition probability matrix (TPM) demonstrated in table (c).
-Finally, we extend to the dynamics of the entire system ABCD -> ABCD, whose TPM is represented in a heatmap (d).
-The effective information and relative metrics of the entire system are as below: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
-[…]
-Figure 8: An example to show how different coarse-graining strategies on the same boolean network shown in Figure 7 can lead to different boolean dynamics.
+If we examine an atomic subsystem with two inputs and one output, such as AB →C, the next state of C depends on the present states of A and B.
+If A = 0 and B = 0, the probability of C being 1 is 70%, whereas the probability of it being 0 is 30%, as indicated in the table (b).
+This same microscopic dynamics applies to other atomic subsystems such as AB →D, CD →A, and CD →B.
+When we expand to a subsystem of two inputs and two outputs, AB →CD and CD →AB, we obtain the transition probability matrix (TPM), as demonstrated in the table (c).
+Finally, we extend to the dynamics of the entire system ABCD →ABCD, whose TPM is represented as a heatmap (d).
+The effective information and relative metrics of the entire system are as follows: $EI_{mi}=1.1486$, $Eff=0.2871$, $Determinism=0.3390$, $Degeneracy=0.0519$
+In order to demonstrate how different coarse-graining strategies can impact effective information, we can combine two micro-states into a single macro-state.
+For example, we can define $\alpha=(A,B)$ and $\beta=(C,D)$ as coarse-grained variables at the macro-level, where their values can be either ‘off’ or ‘on’.
+Using various aggregation strategies, as illustrated in (a)/(b) and (d)/(e) in Figure 8, we can construct the TPMs for the resulting coarse-grained systems.
+The TPM and the corresponding metrics are presented in Figure 8c,f, respectively.
+The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.18$, respectively.
+It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), whereas it has decreased in the second case ($EI_{ma2}(0.18)<EI_{mi}(1.15)$).
+Thus, causal emergence is observed in the first coarse-graining scenario.
+This implies that (1) proper coarse-graining can lead to causal emergence in a macroscopic system, and (2) not all coarse-graining approaches result in higher effective information.
+This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well-defined quantification method.
+This gives us the opportunity to identify causal emergence in a given system.
+This raises the following question: What coarse-graining strategy can generate the maximum effective information?
+We address this question in Section 4.
+Figure 8: An example showing how different coarse-graining strategies applied to the same boolean network, shown in Figure 7, can lead to different boolean dynamics.
 (a) Coarse-graining strategy 1.
-(b) Macro-dynamics leading by strategy 1.
-(c) The corresponding heatmap of TPM and effective information metrics of the macroscopic system by method 1.
+(b) Macro-dynamics resulting from strategy 1.
+(c) The corresponding heatmap of the TPM and the effective information metrics of the macroscopic system using method 1.
 (d) Coarse-graining strategy 2.
-(e) and (f) are the macro-dynamics and TPM correspond to the coarse-graining strategy 2 in (d) respectively.
-The effective information values for the two macroscopic systems are $EI_{ma1}=1.55$ and $EI_{ma2}=0.05$, respectively.
-It is evident that the effective information has increased in the first coarse-graining scenario ($EI_{ma1}(1.55)>EI_{mi}(1.15)$), while it has decreased in the second case ($EI_{ma2}(0.05)<EI_{mi}(1.15)$).
-Thus, causal emergence is observed in the first coarse-graining instance.
-This implies that 1) proper coarse-graining can lead to causal emergence in a macroscopic system, and 2) not all coarse-graining approaches can result in higher effective information.
-This example illustrates that coarse-graining can play an important role in emergence, and effective information is a well defined quantification method.
-This gives us the opportunity to identify causal emergence in a given system.
-There is a question coming next: what coarse-graining strategy can generate the maximum effective information?
-We will address this question in section 4.
-3.6 Extensions and Applications
-We will discuss the extensions and applications of Hoel’s theory of causal emergence in this sub-section.
-3.6.1 Extensions on Continuous Systems
+(e,f) The macro-dynamics and TPM corresponding to coarse-graining strategy 2 in (d), respectively.
+3.5.5 Extensions to Continuous Systems
 Effective information and its associated metrics have been shown to be useful in quantifying causal emergence.
@@
 First, the method is only applicable to discrete-state systems.
-Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate effective information metrics.
-Efforts have been made to apply or extend the concepts of effective information to continuous systems, such as Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
+Second, knowledge of the system’s transition probability matrix (TPM) is required to calculate the effective information metrics.
+Efforts have been made to apply or extend the concepts of effective information to continuous systems Chvykov and Hoel (2020); Varley (2020); Zhang and Liu (2022).
 Ordinal Partition Network
-One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones by ordinal partition network (OPN) method Varley (2020).
-The basic idea is to transform the time series data generated by a continuous dynamical system into sub-time series of dimensionality $d$ with discrete values.
-And then to establish a Transition Probability Matrix (TPM) between these discrete series.
-This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series.
+One approach to incorporating effective information in continuous systems is to transform the continuous variables into discrete ones using the ordinal partition network (OPN) method Varley (2020).
+The basic idea is to transform the time-series data generated by a continuous dynamical system into a sub-time series of dimensionality $d$ with discrete values and then establish a transition probability matrix (TPM) between these discrete series.
+This approach involves three steps: firstly, sampling the original time series with a specified time interval $\tau$ to create sub-time series of dimensionality $d$; secondly, ranking these sub-series based on their values; and finally, utilizing the $d$-dimensional vectors of rank orders to represent the original sub-series.
 By employing rank orders, this method ensures that the values are restricted to integers within a certain range.
-In Varley (2020), the author applied OPN method on Rossler chaotic attractor system.
-By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) is sensitive to the critical phase transition between period and chaos Varley (2020).
+In Varley (2020), the author applied the OPN method to the Rossler chaotic attractor system.
+By quantifying effective information, he noticed that effective information (such as determinism and degeneracy) was sensitive to the critical phase transition between period and chaos Varley (2020).
 Causal Geometry
-Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the conception of causal geometry Chvykov and Hoel (2020).
-To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is:
+Alternatively, Pavel Chvykov and Erik Hoel extended the causal emergence framework to continuous systems and proposed the concept of causal geometry Chvykov and Hoel (2020).
+To understand the basic idea of this work, we suppose the continuous system considered is a continuous functional map and assume that the uncertainties are small disturbances added to this deterministic function, that is,
 $y=f(x)+\varepsilon,$ (18)
-where $x$ is the input variable defined on the interval $[-L/2,L/2]$ and $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a gaussion noise with zero mean and standard deviation $\epsilon\ll L$.
-Then, an approximate form of the effective information for such continuous system can be obtained by Gaussian integral.
+where $x$ is the input variable defined on the interval $[-L/2,L/2]$, $L$ is a very large constant, $y$ is the output variable, and $\varepsilon$ is a Gaussian noise with a zero mean and standard deviation $\epsilon\ll L$.
+Then, an approximate form of the effective information for such a continuous system can be obtained using the Gaussian integral.
 $EI\approx\ln(\frac{L}{\sqrt{2\pi e}})+\frac{1}{2L}\int_{-L/2}^{L/2}\ln\left(\frac{f^{\prime}(x)}{\epsilon}\right)^{2}dx.$ (19)
-The reason why we set the domain of $x$ on $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented, and also facilitate to define a uniform distribution on it.
-This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ as $x(t)$ and $x(t+1)$ respectively.
-The formula also can be extended to higher dimension.
-Suppose $\mathbf{x}\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers, Equation 16 can be generalized to the following form:
+The reason we set the domain of $x$ to $[-L/2,L/2]$ is to guarantee that the Gaussian integral can be implemented and to simplify the definition of a uniform distribution on it.
+This form can be easily generalized to a dynamical system with discrete time steps once we replace $x$ and $y$ with $x(t)$ and $x(t+1)$, respectively.
+The formula can also be extended to higher dimensions, with $\mathbf{x}$ and $\mathbf{y}$ represented in bold.
+Suppose $\mathbf{x}$ $\in[-L/2,L/2]^{n}\subset\mathcal{R}^{n}$ and $\mathbf{y}$ $\in\mathcal{R}^{m}$, where $n$ and $m$ are positive integers.
+Equation (19) can be generalized to the following form:
 $EI\approx\ln(\frac{L^{n}}{(2\pi e)^{m/2}})+\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2},$ (20)
-where, $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, and ||$\cdot$|| ⋅ | is the operation to take absolute value, and $\det$ is to take the determinant.
-Notice that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$, thus the term in the expectation in Equation 17 can be written as:
+where $\Sigma$ is the covariance matrix of the Gaussian noise $\varepsilon$, $U([-L,L]^{n})$ represents the uniform distribution on the hypercube $[-L,L]^{n}$, ||$\cdot$|| ⋅ | is the absolute value operation, and $\det$ is the determinant.
+Note that the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ is a Gaussian distribution, that is, $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
+Thus, the term in the expectation in Equation (20) can be written as:
 $\left|\det\left(\frac{\partial_{\mathbf{x}}f(\mathbf{x})}{\Sigma^{1/2}}\right)\right|^{2}=\left|\det\left(\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right]\right)\right|,$ (21)
-and this is the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
+and this corresponds to the determinant of the negative Fisher information metric for the distribution $p(\mathbf{y}|\mathbf{x})$:
 $g_{\mu\nu}\equiv-\mathbb{E}_{\mathbf{y}|\mathbf{x}}\left[\partial_{\mu}\partial_{\nu}\ln p(\mathbf{y}|\mathbf{x})\right],$ (22)
-it measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ on $\mathbf{x}$.
-Where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$.
-Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, and it is the collections of all possible distributions of $p(\mathbf{y}|\mathbf{x})$.
-This is where the term of “geometry” in this framework comes from.
-Finally, we can obtain an expression for EI with Fisher information metric:
+which measures the sensitivity of the logarithmic $p(\mathbf{y}|\mathbf{x})$ to changes in $\mathbf{x}$, where $\partial_{\mu}\equiv\frac{\partial}{\partial\mathbf{x}_{\mu}}$ represents the partial derivative with respect to the $\mu$th component of $\mathbf{x}$.
+Therefore, we have defined a distance metric for the Riemann manifold =$\mathcal{M}=\{p(\mathbf{y}|\mathbf{x})\}$ on the parameter space $\mathbf{x}\in[-L/2,L/2]^{n}$, encompassing all possible distributions of $p(\mathbf{y}|\mathbf{x})$.
+This is the origin of the term “geometry” in this framework.
+Finally, we can obtain an expression for EI with the Fisher information metric:
 $EI\approx\ln\frac{L^{n}}{(2\pi e)^{m/2}}-\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim U([-\frac{L}{2},\frac{L}{2}]^{n})}\ln|\det(g_{\mu\nu})|.$ (23)
-This formula may be generalized to the cases where $p(\mathbf{y}|\mathbf{x})$ is non-Gaussian distribution.
-Once the distribution function is known, we can obtain its Fisher information metric, then EI can be calculated via Equation 20.
-The reason behind is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
-While, in Chvykov and Hoel (2020), the authors consider a more general case, in which the intervention noise is added on the input(intervention) variable $x$ in Equation 16.
-The intervention noise is $\xi\sim\mathcal{N}(0,\delta^{2})$, and $\delta$ is the standard deviation of $\xi$.
-While, $\varepsilon$ is called observational noise to differentiate with $\xi$, and it is added on the output(observational) variable $y$ according to Equation 16.
-Finally, if two kinds of noises are both considered, and if $L=1$ and $\epsilon\ll 1$, Equation 16 becomes:
+This formula can be generalized to cases where $p(\mathbf{y}|\mathbf{x})$ is a non-Gaussian distribution.
+Once the distribution function is known, we can obtain its Fisher information metric, and then EI can be calculated using Equation (23).
+The reason behind this is that the whole manifold $p(\mathbf{y}|\mathbf{x})$ for any $\mathbf{x}$ can be understood as a concatenation of local Gaussian distributions.
+In Chvykov and Hoel (2020), the authors considered a more general case in which intervention noise is added to the input (intervention) variable $x$ in Equation (19).
+The intervention noise is denoted as $\xi\sim\mathcal{N}(0,\delta^{2})$, where $\delta$ is the standard deviation of $\xi$.
+In contrast, $\varepsilon$ is called observational noise distinguishing it from $\xi$, and is added to the output (observational) variable $y$ according to Equation (19).
+Finally, if two kinds of noise are both considered and if $L=1$ and $\epsilon\ll 1$, Equation (19) becomes:
 […]
-This is the formula of EI for continuous mapping function (Equation 15) given by Chvykov and Hoel (2020).
-Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation 15) and a more discrete map with an adjustable parameter in the logistic function.
-They discovered that when the noise level is low, the continuous map may exhibit higher EI than the discrete map.
+This is the formula for EI for the continuous mapping function (Equation (18)) given by Chvykov and Hoel (2020).
+Using this equation and a logistic function for $f(x)$, the authors compared the EIs of a more continuous map (Equation (18)) and a more discrete map with an adjustable parameter in the logistic function.
+They discovered that when the noise level is low, the continuous map can exhibit higher EI compared to the discrete map.
 However, as the noise level increases, discretizing the mapping function can lead to a model with higher EI.
-This phenomenon helps to explain why digital circuits eventually outperform analog circuits in mitigating noise interference: the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
-To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we can not control $\mathbf{y}$ by intervening $\mathbf{x}$ directly.
-Instead, we can intervene $\mathbf{x}$ to influence $\theta$, and to affect $\mathbf{y}$ indirectly.
+This phenomenon helps explain why digital circuits eventually outperform analog circuits in mitigating noise interference; the binarization and coarse-graining strategy of digital circuits suppresses the propagation and diffusion of noise.
+To generalize the information geometry to the case with intervention noise and observational noise, let us introduce a new intermediate variable $\theta\in\Theta\subset\mathcal{R}^{l}$ with dimensionality $l$ such that we cannot control $\mathbf{y}$ by directly intervening $\mathbf{x}$.
+Instead, we can intervene $\mathbf{x}$ to influence $\theta$ and indirectly affect $\mathbf{y}$.
 Therefore, the three variables form a Markov chain: $\mathbf{x}\rightarrow\theta\rightarrow\mathbf{y}$.
 […]
-Finally, the EI calculation formula for the causal geometry is:
+Finally, the EI calculation formula for causal geometry is:
 $EI_{g}=\ln\frac{V_{I}}{(2\pi e)^{n/2}}-\frac{1}{2V_{I}}\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}\ln\left|\det\left(I_{n}+\frac{h_{\mu\nu}}{g_{\mu\nu}}\right)\right|d^{l}\theta,$ (25)
-in which, we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
+where we have set $L=1$ and $m=l=n$ to reduce the number of free parameters, and $I_{n}$ is the identity matrix with size $n$, $V_{I}=\int_{\Theta}\sqrt{|\det(h_{\mu\nu})|}d^{l}\theta$.
 EI Calculation for Neural Networks
-One of the application areas for continuous EI is on neural networks.
+One of the application areas for continuous EI is neural networks.
 In Marrow et al.
-(2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compare the results with information bottleneck theory.
-The method to calculate EI that they used is to convert continuous variables into discrete ones by dividing the domains of input and output variables into small regions.
-However, this method has high computational complexity, and is hard to extend to high dimensions.
-In Zhang and Liu (2022), the authors propose a new numeric method for calculating EI on neural networks.
-The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping following:
+(2020), the authors applied EI to analyze the causal effects for different layers of neural networks and compared the results with the information bottleneck theory.
+The method they used to calculate EI was to convert continuous variables into discrete ones by dividing the domains of the input and output variables into small regions.
+However, this method has high computational complexity and is challenging to extend to high dimensions.
+In Zhang and Liu (2022), the authors proposed a new numeric method for calculating EI on neural networks.
+The basic idea behind this method is to understand a well-trained feed-forward neural network as a stochastic mapping, following:
 $\mathbf{y}=f(\mathbf{x})+\varepsilon,$ (26)
-where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, and $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network, $\varepsilon$ is a Gaussian noise with zero means and the covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ on the whole training data set.
+where $\mathbf{x}\in\mathcal{R}^{n}$ is the input variable, $\mathbf{y}\in\mathcal{R}^{m}$ is the output variable, and $f:\mathcal{R}^{n}\rightarrow\mathcal{R}^{m}$ is the deterministic map of the neural network.
+Additionally, $\varepsilon$ is a Gaussian noise with a mean of zero and a covariance matrix $\Sigma=diag(\sigma_{1}^{2},\sigma_{2}^{2},\cdot\cdot\cdot,\sigma_{m}^{2})$, where $\sigma_{i}$ is the mean square error on dimension $i$ across the entire training dataset.
 In this way, the neural network can be regarded as a Gaussian distribution $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(f(\mathbf{x}),\Sigma)$.
-Therefore, Equation 17 can be used on this neural network to calculate EI.
-In Zhang and Liu (2022), the authors use Monte Carlo integration method to estimate the expectation in Equation 17, this technique reduces the computational complexity in a large degree.
-However, when the authors compared the results calculated by Equation 17 for different neural networks with different dimensions, they obtained unreasonable results: EI increases with the number of dimensions, thus, causal emergence is always not be observed because micro-dynamics has unreasonably larger EI.
-Another defect of Equation 17 is that the parameter $L$ always dominates the value of the expression which should be removed from the equation.
-One way to solve this issue is to use Eff instead of EI (see Equation 10), however, not all $L$s in the Equation 17 can be eliminated.
-Therefore, the authors invented a notion of dimension averaged EI,
-$dEI\equiv=\frac{EI}{m}$ (24)
-which simply divide EI by the dimension of the output variable $\mathbf{y}$.
-By using dimension averaged EI, $L$ can be also removed if we calculate dimensional averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ for macro- and micro- dynamics respectively, that is:
+Therefore, Equation (20) can be used on this neural network to calculate EI.
+In Zhang and Liu (2022), the authors used the Monte Carlo integration method to estimate the expectation in Equation (20).
+This technique significantly reduces the computational complexity.
+However, when the authors compared the results calculated using Equation (20) for various neural networks with different dimensions, they obtained unreasonable results: EI increased with the number of dimensions.
+Thus, causal emergence was not always observed because micro-dynamics exhibited unreasonably larger EI.
+Another drawback of Equation (20) is that the parameter $L$ always dominates the value of the expression and should be removed from the equation.
+One way to solve this issue is to use Eff instead of EI (see Equation (10)).
+However, not all $L$s in Equation (20) can be eliminated.
+Therefore, the authors introduced the notion of dimension-averaged EI, defined as
+$dEI\equiv\frac{EI}{m}$ (27)
+which simply divides EI by the dimension of the output variable $\mathbf{y}$.
+By using dimension-averaged EI, $L$ can be removed when calculating dimension-averaged causal emergence for the neural networks $f_{M}$ and $f_{m}$ representing macro- and micro-dynamics, respectively, given by
 $dCE(f_{M},f_{m})\equiv\frac{EI(f_{M})}{n_{M}}-\frac{EI(f_{m})}{n_{m}},$ (28)
-where, dCE is the dimension averaged causal emergence measure. If we bring Equation 17 into Equation 25, we get:
+where dCE is the dimension-averaged causal emergence measure. If we incorporate Equation (20) into Equation (28), we obtain:
 $dCE\approx\frac{\mathbb{E}_{\mathbf{x}_{M}}\log|\det(\partial_{\mathbf{x}_{M}}f_{M})|+\sum_{i=1}^{x_{M}}\log\sigma_{i,M}^{2}}{n_{M}}-\frac{\mathbb{E}_{\mathbf{x}_{m}}\log|\det(\partial_{\mathbf{x}_{m}}f_{m})|+\sum_{i=1}^{x_{m}}\log\sigma_{i,m}^{2}}{n_{m}},$ (29)
-in which, $L$ is eliminated.
-3.6.2 Causal emergence in complex network
-Many complex systems can be represented by networks, therefore, to apply the framework of causal emergence on networks is necessary.
-However, two problems should be solved in advance to apply Hoel’s framework, one is to assign a dynamics on the studied network because networks are static structures without any dynamical properties, another one is to coarse-grain the network.
-Klein and Hoel Klein and Hoel (2020) solved the first problem by considering random walks on complex networks.
-Subsequently, the TPM of the system is defined based on the transfer probabilities between nodes of a large number of random walkers.
-Due to the good mathematical properties of the random walk dynamics on graphs, the authors can establish an explicit expression of EI on networks, and the final form only relies on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network.
-The expression is as follow:
+in which $L$ is eliminated.
+3.5.6 Causal Emergence in Complex Networks
+Many complex systems can be represented by networks; therefore, applying the framework for causal emergence to networks is necessary.
+However, two problems should be solved in advance to apply Hoel’s framework: one is to assign dynamics to the studied network because networks are static structures without any dynamical properties and the other is to coarse-grain the network.
+Klein and Hoel Klein and Hoel (2020) addressed the first problem by considering random walks on complex networks.
+Subsequently, the TPM of the system was defined based on the transfer probabilities between nodes of a large number of random walkers.
+Due to the good mathematical properties of the random walk dynamics on graphs, the authors established an explicit expression for EI on networks, and the final form only relied on the weighted normalized adjacency matrix $w_{ij}$($\sum_{j}w_{ij}=1$) of the network.
+The expression is as follows:
 $EI=H(\langle W^{out}_{i}\rangle)-\langle H(W^{out}_{i})\rangle.$ (30)
-Where, $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,...,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes which characterizes the degeneracy of the dynamics.
-And $H(W^{out}_{i})$ is the uncertainty of node $i$.
-To tackle the second problem, the greedy algorithm is used to group nodes to form a macro network in Klein and Hoel (2020).
-It is worth noting that when merging micro nodes into macro nodes, the TPM of the corresponding macro network can be derived by merging the probabilities in TPM of the micro-level.
-To ensure that the grouped macro network maintains the same random walk dynamics as the original one, dynamic consistency test is implemented.
-In the experimental section, the paper explores the application of greedy search for identifying causal emergence in complex networks.
-The experiments are conducted on artificial networks, as well as four types of real networks.
-In the case of ER random networks, the size of effective information is solely dependent on the connection probability $p$, and converges to $-\log_{2}p$ with the increase of network size.
-Additionally, a significant finding reveals the presence of a phase transition point, where the average degree of the network reaches approximately $\log_{2}N$.
-Beyond this point, the random network structure does not contain additional information with increasing size.
-In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increases as the network size expands.
-Conversely, when $\alpha>1.0$, the opposite conclusion is reached.
-The scale-free network corresponding to $\alpha=1.0$ represents the critical boundary of growth.
-Regarding real networks, the authors find that biological networks exhibit the lowest EI due to the presence of significant noise, which can be removed by effective coarse-grained, and causal emergence is the most significant compared with other types of networks.
-However, technical networks exhibit sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships.
-Consequently, they possess the highest EI among the networks studied.
+where $H(\langle W^{out}_{i}\rangle)$ represents the Shannon entropy calculated from the distribution of the averaged out-weights across all nodes ($W^{out}_{i}\equiv\{w_{ij}|j=1,2,\dots,N\}$), which characterizes the determinism of the random walk dynamics; $\langle H(W^{out}_{i})\rangle$ is the average Shannon entropy for all nodes, which characterizes the degeneracy of the dynamics; and $H(W^{out}_{i})$ is the uncertainty of node $i$.
+To address the second problem, in Klein and Hoel (2020), the greedy algorithm was used to group nodes to form a macro-network.
+It is worth noting that when merging micro-nodes into macro-nodes, the TPM of the resulting macro-network can be derived by merging the probabilities from the TPM of the micro-level.
+To ensure that the grouped macro-network maintains the same random walk dynamics as the original one, a dynamic consistency test is implemented.
+In the experimental section, the aforementioned paper explored the application of a greedy search for identifying causal emergence in complex networks.
+Experiments were conducted on artificial networks as well as four types of real networks.
+In the case of ER random networks, the size of effective information was solely dependent on the connection probability $p$ and converged to $-\log_{2}p$ with the increase in network size.
+Additionally, a significant finding was the presence of a phase transition point, where the average degree of the network reached approximately $\log_{2}N$.
+Beyond this point, the random network structure did not contain additional information with increasing size.
+In the case of preferential attachment (PA) networks, when $\alpha<1.0$ ($\alpha$ represents the degree of preferential attachment), the effective information of the network increased as the network size expanded.
+Conversely, when $\alpha>1.0$, the opposite was observed.
+The scale-free network corresponding to $\alpha=1.0$ represented the critical boundary of growth.
+Regarding real networks, the authors found that biological networks exhibited the lowest EI due to the presence of significant noise, which can be removed through effective coarse-graining, and causal emergence was the most significant compared to other types of networks.
+However, technical networks exhibited sparsity and non-degeneracy, resulting in higher average efficiency and more specific node relationships.
+Consequently, they exhibited the highest EIs among the studied networks.
 The network coarse-graining method mentioned above is the greedy algorithm.
@@
 Griebenow et al.
-(2019) introduced a spectral clustering-based approach to identify causal emergence within networks.
+(2019) introduced a spectral clustering-based approach for identifying causal emergence within networks.
 More specifically, the method involved performing eigenvalue decomposition of the TPM, followed by constructing a similarity matrix using the eigenvectors of the nodes.
-The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro node.
-Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyper-parameter $\epsilon$.
-In this paper, the author additionally proposed a gradient descent algorithm based on deep learning.
-This approach encompasses several steps, beginning with the random initialization of a grouping matrix.
-Then this matrix is used to construct a macro network by combining the micro network.
-Finally, the grouping strategies are automatically learned through the process of maximizing effective information within the macro network.
+The OPTICS algorithm was employed to cluster the nodes, and nodes belonging to the same cluster were aggregated into a macro-node.
+Subsequently, the maximum value of EI was selected by utilizing the linear search distance hyperparameter $\epsilon$.
+In their paper, the authors additionally proposed a gradient descent algorithm based on deep learning.
+This approach encompassed several steps, beginning with the random initialization of a grouping matrix.
+Then, this matrix was used to construct a macro-network by combining the micro-networks.
+Finally, the grouping strategies were automatically learned through the process of maximizing effective information within the macro-network.
 However, this method often falls into the local optimal solution.
@@
 Once the method for quantifying causal emergence in complex systems is developed, it can be applied across various fields that possess abundant network data.
-The first type of networks which was studied is biological networks.
+The first type of network studied was biological networks.
 As previously discussed, biological networks are full of noise, which poses challenges in comprehending their internal operating principles.
-On the one hand, such noise arises from the inherent fluctuations within the system itself, while on the other hand, it can be introduced through measurement or observation processes.
+On one hand, such noise arises from inherent fluctuations within the system itself, whereas on the other hand, it can be introduced through measurement or observation processes.
 Consequently, Klein et al.
 Klein et al.
-(2022) further explored the relationship among noise, degeneracy and certainty in biological networks and their specific meanings.
+(2022) further explored the relationship between noise, degeneracy, and certainty in biological networks and their specific meanings.
 For instance, in gene expression networks, highly deterministic relationships indicate that the expression of one gene almost invariably leads to the expression of another gene.
@@
 Klein et al.
-(2021) conducted an analysis of protein interaction networks across more than 1,800 species.
-They employed EI as a measure to assess the levels of noise and uncertainty in protein interactions.
+(2021) conducted an analysis of protein interaction networks across more than 1800 species.
+They employed EI as a measure for assessing the levels of noise and uncertainty in protein interactions.
 The findings revealed that the macro-scale network exhibited lower levels of noise and degeneracy.
-Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to nodes that were not part of the macro-scale interaction group.
-Through robust analysis, the authors demonstrate that eukaryotes exhibit a stronger degree of causal emergence compared to archaea.
-Additionally, to address the ’deterministic paradox,’ the authors introduce the concepts of the neutral process and the selection process in biological evolution.
-The neutral process operates at the micro scale, leveraging mutations to promote interactions and enhance species diversity.
-On the other hand, the selection process operates at the macro scale, effectively eliminating noise that hampers system operation and efficiency.
+Additionally, the nodes within the macro-scale interaction group demonstrated greater resilience compared to the nodes that were not part of the macro-scale interaction group.
+Through robust analysis, the authors demonstrated that eukaryotes exhibited a stronger degree of causal emergence compared to archaea.
+Additionally, to address the ’deterministic paradox,’ the authors introduced the concepts of the neutral process and the selection process in biological evolution.
+The neutral process operates at the micro-scale, leveraging mutations to promote interactions and enhance species diversity.
+On the other hand, the selection process operates at the macro-scale, effectively eliminating noise that hampers system operation and efficiency.
 Therefore, in order to adapt to the demands of evolution, it becomes essential for evolved biological systems to function across multiple scales.
@@
 Hoel and Levin (2020) conducted further research on the causal emergence within biological systems.
-The authors elaborated that macro and micro systems exist widely in biological systems, for example, the micro scale of a group of cells can be the potential ion channel changes, and the macro scale corresponds to the membrane potential changes of cells.
+The authors elaborated that macro- and micro-systems exist widely in biological systems.
+For example, the micro-scale of a group of cells can involve potential ion channel changes, whereas the macro-scale corresponds to the membrane potential changes of cells.
 Furthermore, the authors utilized EI to analyze gene regulatory networks, aiming to identify the most informative models for controlling mammalian heart development.
-By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it is revealed that informative macro-scale structures are prevalent across biological systems.
-Additionally, the authors emphasize the importance of evolutionary systems operating at multiple scales, as it offers significant advantages.
+By quantifying the causal emergence within the largest component of the gene network of Saccharomyces cerevisia, it was revealed that informative macro-scale structures were prevalent across biological systems.
+Additionally, the authors emphasized the importance of evolutionary systems operating at multiple scales due to the significant advantages they offer.
 Natural selection requires variation between populations, and degradation observed in biological systems serves as a crucial factor for evolution.
@@
 (2022) conducted an investigation into the impact of ant colony interaction history on task assignment and task switching.
-They employed effective information to examine how noise information propagates among ants, while also exploring the relationship between EI and the proportion of ants assigned to different tasks within ant colonies.
+They employed effective information to examine how noise information propagates among ants and explored the relationship between EI and the proportion of ants assigned to different tasks within ant colonies.
 The study revealed that the extent of historical interaction between ant colonies influenced task assignment.
@@
 For example, the interactions between foragers displayed significantly higher levels of noise when contrasted with interactions between nurses or cleaners.
-Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also play different roles in maintaining group cohesion.
-EI indicator and causal emergence theoretic framework can also be applied to artificial systems.
+Furthermore, even when ants switched functional groups, the cohesion within ant colonies ensured the stability of the overall colony, and different functional ant colonies also played different roles in maintaining group cohesion.
+The EI indicator and causal emergence theoretic framework can also be applied to artificial systems.
 For example, Marrow et al.
 Marrow et al.
-(2020) quantify and monitor the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer.
-By observing the changes of EI, including determinism(sensitivity) and degeneracy, throughout model training, the generalization ability of the model can be determined, thus helping us to better understand and explain the working principle of neural networks.
+(2020) quantified and monitored the changes in the causal structure of neural networks during training, in which EI was employed to evaluate the degree of causal influence of nodes and edges on downstream tasks at each layer.
+By observing the changes in EI, including determinism (sensitivity) and degeneracy, throughout model training, the generalization ability of the model could be determined, thus helping better understand and explain the working principle of neural networks.
 Varley et al.
 Varley (2020) attempted to apply the causal emergence framework to both discrete cellular automata and continuous Rossler systems.
-In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic and complex.
+In the case of cellular automata systems, the authors select 88 unique rules corresponding to four types: static, periodic, chaotic, and complex.
 By considering each state as a node, where each state determines the subsequent state, a directed state transition graph is constructed.
@@
 In addition, the authors also draw some quantitative conclusions, for example, among the 17 rules belonging to the third and fourth categories, 30% exist causal emergence, 70% show causal degradation, and the CE of cellular automata with the same rules remained relatively consistent across different sizes.
-Furthermore, the author employed the OPN algorithm to transform the continuous system into a discrete state transition graph for comparative analysis (see section 3.6.1).
-They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
-3.7 Critiques
-In the extensive literature on causality and emergence, Hoel’s theory has gained attention for its association of emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner.
-However, Dewhurst  Dewhurst (2021) provides some philosophical clarity of Hoel’s theory, arguing that it is epistemological rather than metaphysical.
-It suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving "genuinely novel causal powers." This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
-The computation of Hoel’s effective information relies on two premises: 1) knowledge of the system’s microscopic dynamics, and 2) knowledge of the coarse-graining scheme.
+Furthermore, the authors employed the OPN algorithm to transform the continuous system into a discrete-state transition graph for comparative analysis (see Section 3.5.5).
+They found that chaos dynamics demonstrated a correlation with low determinism, and the variations in the degeneracy and efficiency coefficients aligned with the changes observed in the determinism curve.
+3.5.8 Critiques
+In the extensive literature on causality and emergence, Hoel’s theory has attracted attention for linking emergence and causality through interventionism, introducing the concept of causal emergence in a quantitative manner.
+However, Dewhurst Dewhurst (2021) provided a philosophical clarification of Hoel’s theory, arguing that it was epistemological rather than metaphysical.
+This suggests that Hoel’s macroscopic causality is merely a causal explanation based on information theory, rather than involving “genuinely novel causal powers”.
+This also raises concerns about the assumption of a uniform distribution, as there is no empirical basis to favor it over other distributions.
+The computation of Hoel’s effective information relies on two premises: (1) knowledge of the system’s microscopic dynamics, and (2) knowledge of the coarse-graining scheme.
 However, in practice, it is rare for both conditions to be simultaneously satisfied, especially in observational studies where both are unknown.
 This limitation hinders the practical applicability of Hoel’s theory.
-It is pointed out that Hoel’s theory negleted the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022).
+It has been pointed out that Hoel’s theory neglects the constraints on coarse-graining methods, and certain coarse-graining methods can lead to ambiguity Eberhardt and Lee (2022).
 Additionally, the combination of some coarse-graining operations over states and coarse-graining operations over time does not exhibit commutativity.
-For instance, assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold.
+For instance, by assuming $A_{m\times n}$ is a coarse-graining operation over states (merging $n$ states into $m$ states), and ()×()($\cdot$)$\times$($\cdot$)( ⋅ ) × ( ⋅ ) is a coarse-graining operation over time (combining two time steps into one), the equation $A_{m\times n}(TPM_{n\times n})\times A_{m\times n}(TPM_{n\times n})=A_{m\times n}(TPM_{n\times n}\times TPM_{n\times n})$ does not always hold.
 This indicates that certain coarse-graining operations can result in a discrepancy between the evolution of macroscopic states and the coarse-grained states of the evolved microscopic systems.
 It implies the need for consistent constraints on coarse-graining strategies.
-That means solely maximizing EI may raise some problems, and further constraints must be added in the framework. We will discuss this problem in the section of 4.1.2.
-3.8 Rosas Fernando’s Quantifying Causal Emergence
+This means that solely maximizing EI may raise some problems, and further constraints must be added to the framework. We discuss this problem in Section 4.1.1.
+3.6 Fernando E. Rosas’s Quantification of Causal Emergence
 3.6.1 Basic Idea
 In Hoel’s framework, it is essential to find a coarse-graining strategy in order to determine the occurrence of causal emergence, and the outcome is influenced by the choice of the coarse-graining method.
-While it has been suggested by Hoel Hoel et al.
-(2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised as proposed by Dewhurst Dewhurst (2021).
-Third, EI is an global measure because it requires to intervene the input variable $X$ to be a uniform distribution on the whole domain Rosas et al.
+Although it has been suggested by Hoel Hoel et al.
+(2013); Hoel (2017) that an optimal strategy can be identified by maximizing EI, certain issues have been raised by Dewhurst Dewhurst (2021).
+EI is a global measure because it requires the input variable $X$ to be a uniform distribution over the whole domain Rosas et al.
 (2020).
-However, many regions are not be able to be observed from data.
-Consequently, there is a pressing need for an alternative theoretical framework of causal emergence that does not rely on a coarse-graining method.
+However, many regions are not observable from the data.
+Consequently, there is a pressing need for an alternative theoretical framework for causal emergence that does not rely on a coarse-graining method.
 In response, Fernando E.
-Rosas took an approach that did not require a coarse-graining strategy as prerequisite, attempting to break down excess entropy - the mutual information between a system’s past and future states - into non-overlapping parts to identify the information components most relevant to causal emergence.
-To accomplish this, he relied on the Partial Information Decomposition (PID) framework proposed by Williams and Beer, which provides a method for non-overlapping decomposition of joint mutual information Williams and Beer (2010).
-First, we will introduce Williams and Beer’s theoretic framework.
+Rosas took an approach that did not require a coarse-graining strategy as a prerequisite, attempting to break down excess entropy—the mutual information between a system’s past and future states—into non-overlapping parts to identify the information components most relevant to causal emergence.
+To accomplish this, he relied on the partial information decomposition (PID) framework proposed by Williams and Beer, which provides a method for the non-overlapping decomposition of joint mutual information Williams and Beer (2010).
+Below, we introduce Williams and Beer’s theoretical framework.
 3.6.2 Partial Information Decomposition
 The partial information decomposition (PID) framework investigates the general informational relationship between source variables and a target variable.
-To make the problem description simpler without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
-Figure 9: Venn diagram of PID.
-(a) Relationship of mutual information between Y and X1, X2.
-(b-d) Examples that joint mutual information is solely contributed by redundant, synergistic, and unique information.
-(e) Redundancy lattice representation of (a).
-The mutual information between the target variable and individual source variables,$I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibit a complex relationship and can not be intuitively converted to one another.
-Nonetheless, it’s an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
+To simplify the problem description without sacrificing generality, let us consider a system with two input variables ($X_{1}$, $X_{2}$) and one output variable ($Y$) as an example, as depicted in the Venn diagram below.
+The mutual information between the target variable and individual source variables, $I(X_{1};Y)$ and $I(X_{2};Y)$, as well as the mutual information between the target variable and the joint source variable, $I(X_{1},X_{2};Y)$, exhibits a complex relationship.
+Intuitively, one cannot be converted to another.
+Nonetheless, it is an intuitive notion that the overlapped circular plates of $I(X_{1};Y)$ and $I(X_{2};Y)$ divide the oval region of $I(X_{1},X_{2};Y)$ into four adjacent non-overlapping regions, representing the three types of information components of $I(X_{1},X_{2};Y)$:
 $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)+Un(X_{1};Y|X_{2})+Un(X_{2};Y|X_{1})+Syn(X_{1},X_{2};Y)$ (31)
@@
   • $Syn$: Synergistic information is the information held by all sources together, but not any individual one.
-If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates that they must be independent of one another.
-For a more intuitive understanding, let’s consider a few simple toy examples.
+If we can identify variable representations corresponding to these information components, the non-overlapping feature in the Venn diagram indicates their independence from one another.
+For a more intuitive understanding, let us consider a few simple toy examples.
 Case 1, $Y=X_{1}=X_{2}$: This is a scenario in which a single source variable can predict the target variable, and the addition of another source variable does not enhance the prediction of the target variable.
-In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, Fig 9(b).
+In this case, we have $I(X_{1},X_{2};Y)=Red(X_{1},X_{2};Y)$, as depicted in Figure 9b.
 Case 2, $Y=X_{1}\oplus X_{2},X_{1}\perp{X_{2}}$: In another scenario, neither of the source variables can predict the target variable individually, but together, they can predict the target variable synergistically.
-In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, Fig 9(c).
-Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, Fig 9(d).
-For general cases, Williams and Beer presented a method for calculating redundant information defined as $Red_{min}$.
-This method reflects the concept of redundancy, identifying it as the shared information across all sources, is equivalent to the minimum amount of information contributed by any single source:
+In this case, $I(X_{1},X_{2};Y)=Syn(X_{1},X_{2};Y)$, as depicted in Figure 9c.
+Case 3, $Y=X_{1},X_{1}\perp{X_{2}}$: In this scenario, the target variable can be predicted by one of the source variables but not the other, which implies $I(X_{1},X_{2};Y)=Un(X_{1};Y|X_{2})$, as depicted in Figure 9d.
+Figure 9: Venn diagram of PID.
+(a) Relationship of mutual information between Y and X1, X2.
+(bd) Examples showing that joint mutual information is solely contributed by redundant, synergistic, and unique information.
+(e) Redundancy lattice representation of (a).
+For general cases, Williams and Beer presented a method for calculating redundant information, defined as $Red_{min}$.
+This method reflects the concept of redundancy, identifying it as the shared information across all sources, which is equivalent to the minimum amount of information contributed by any single source:
 $Red_{min}(X_{1},X_{2};Y)\sum_{y\in\mathcal{Y}}P(Y=y)\min_{Xi}I(Y=y;X_{i})$ (32)
-Fig 9(e) presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010).
+Figure 9e presents an alternative way of visualizing the PID framework, known as the redundancy lattice Williams and Beer (2010).
 In this representation, {12} denotes synergistic information, $\left\{1\right\}$ and $\left\{2\right\}$ denote unique information, and $\left\{1\right\}$$\left\{2\right\}$ denotes redundant information.
-In the subsequent sections, we will also see that the $\phi$ID framework employs the notation of the redundancy lattice.
+In the subsequent sections, we demonstrate that the $\phi$ID framework employs the notation of the redundancy lattice.
 Although the PID framework remains compatible with scenarios involving more than two source variables, it should be noted that the corresponding Venn diagrams and redundancy lattices for these scenarios can become substantially more complicated and difficult to decipher, as discussed in Williams and Beer (2010).
@@
 To address this challenge, Fernando E.
-Rosas etc have developed Integrated Information Decomposition ($\phi$ID) Mediano et al.
+Rosas developed Integrated Information Decomposition ($\phi$ID) Mediano et al.
 (2019), which takes its name from Integrated Information Theory (IIT) Oizumi et al.
 (2014).
-This extension of PID provides a more comprehensive means for analyzing dynamic systems.
-To introduce Rosas’s framework clearly, we only consider a system with two variables. All the definitions and calculations can be generalized to the systems with more variables.
+This extension of PID provides a more comprehensive method for analyzing dynamic systems.
+To introduce Rosas’s framework clearly, we consider a system with only two variables. All the definitions and calculations can be generalized to systems with more variables.
 The objective of the $\phi$ID framework is to decompose excess entropy into non-overlapping information components.
-In a two-variable Markovian system, the excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states.
-From the causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, while $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent effects.
-Rosas first noted that mathematically, the PID framework is symmetric with respect to source and target variables.
+In a two-variable Markovian system, excess entropy is given by $E=I(X^{1}_{t},X^{2}_{t};X^{1}_{t+1},X^{2}_{t+1})$, where $X^{1}_{t}$ and $X^{2}_{t}$ represent the current states, and $X^{1}_{t+1}$, $X^{2}_{t+1}$ represent the future states.
+From a causation perspective, $X^{1}_{t}$ and $X^{2}_{t}$ represent causes, whereas $X^{1}_{t+1}$ and $X^{2}_{t+1}$ represent effects.
+Rosas first noted that mathematically, the PID framework is symmetric with respect to the source and target variables.
 There are two viewpoints for analyzing the above Markovian system with PID.
-One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the "cause" to elements), while the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the "effect" to elements).
-The redundancy lattices of these perspectives are illustrated in Fig 10 (a) and (b), and respectively, are referred to as "forward PID" and " PID".
-Figure 10: (a) Redundancy lattice of "forward PID", and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of "backward PID" and its target variable $X^{1}_{t}X^{2}_{t}$.
-Next, Rosas introduced the $\phi$ID framework, which consolidates the forward PID and backward PID into a single framework.
-In this framework, the one-to-many relationship in PID is expanded to include many-to-many relationships.
-He built full connections between the elements of redundancy lattices of the forward PID and the backward PID.
-This approach generates 16 relations between the source and target, as depicted by the colored lines in Fig 11 (a).
-(Each color corresponds to a specific source element of $X_{t}$.) These relations are referred to as "$\phi$ID atoms" and can be represented as vertices in a lattice, as shown in Fig 11 (b).
-This more complex lattice is referred to as the double-redundancy lattice because it is a "product" of two redundancy lattices (of forward PID and backward PID).
-Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID. (b) Double-redundancy lattice, whose vertices correspond to the edges in (a)
-In the double-redundancy lattice, each vertex is defined as a "$\phi$ID atom," denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, let $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$.
-For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is demoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$.
-$\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system.
+One viewpoint takes $X^{1}_{t+1}$ and $X^{2}_{t+1}$ together as the target variable (aiming to decompose the “cause” to elements), whereas the other viewpoint takes $X^{1}_{t}$ and $X^{2}_{t}$ together as the target variable (aiming to decompose the “effect” to elements).
+The redundancy lattices of these perspectives are illustrated in Figure 10a,b, and respectively, are referred to as “forward PID” and “ PID”.
+Figure 10: (a) Redundancy lattice of “forward PID” and its target variable $X^{1}_{t+1}X^{2}_{t+1}$. (b) Redundancy lattice of “backward PID” and its target variable $X^{1}_{t}X^{2}_{t}$.
+Next, Rosas introduced the $\phi$ID framework, consolidating forward PID and backward PID into a single framework.
+In this framework, the one-to-many relationship in PID was expanded to include many-to-many relationships.
+He built full connections between the elements of the redundancy lattices of forward PID and backward PID.
+This approach generates 16 relations between the source and target, as depicted by the colored lines in Figure 11a (each color corresponds to a specific source element of $X_{t}$).
+These relations are referred to as “$\phi$ID atoms” and can be represented as vertices in a lattice, as shown in Figure 11b.
+This more complex lattice is referred to as the double-redundancy lattice because it is a “product” of two redundancy lattices (of forward PID and backward PID).
+Figure 11: (a) Relationships between redundancy lattices of forward PID and backward PID.
+(b) Double-redundancy lattice, whose vertices correspond to the edges in (a).
+Distinct colors of dots and lines are employed to distinguish the source elements of $X_{t}$.
+In the double-redundancy lattice, each vertex is defined as a “$\phi$ID atom”, denoted as $I^{\alpha\rightarrow\beta}_{\partial},\alpha,\beta\in\mathcal{A}$, where $\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$.
+For instance, the $\phi$ID atom of the vertice ${\{12\}\rightarrow\{12\}}$ is denoted as $I^{\{12\}\rightarrow\{12\}}_{\partial}$.
+The $\phi$ID atom represents the broken-down information transmitted from the cause elements to the effect elements, and their cumulative sum amounts to the excess entropy of the entire system.
 Consequently, the formal definition of Integrated Information Decomposition ($\phi$ID) is:
 $\mathbb{E}=I(\mathbb{X_{t}};\mathbb{X_{t+1}})=\sum_{\alpha,\beta\in\mathcal{A}}{I^{\alpha\rightarrow\beta}_{\partial}}$ (33)
-The main advantage of $\phi$ID is that it can provide more fine-grained analysis of complex systems than PID by introducing the temporal dynamic.
+The main advantage of $\phi$ID is that it can provide a more fine-grained analysis of complex systems compared to PID by introducing the temporal dynamic.
 By decomposing the mutual information of the system at different times, $\phi$ID provides another approach for the quantitative study of causal emergence.
 3.6.4 Reconciling Emergences
-Based on the concept of synergistic information in the PID framework, Rosas proposed a quantitative definition of causal emergence and further classified it using the $\phi$ID framework, which addresses the problem of finding a coarse-graining strategy.
-The definition includes two aspects: the first is whether the system has the ability to produce causal emergence objectively.
-Then, on this basis, whether there is causal emergence happening with specific macroscopic features.
-In terms of the capacity to generate causal emergence for a system, this definition links causal emergence to synergistic relationships among variables in different times.
-Therefore, a system $X_{t}$ is said to have capacity for causally emergent features if and only if that:
+Building upon the concept of synergistic information in the PID framework, Rosas introduced a quantitative definition of causal emergence using the $\phi$ID framework to tackle the challenge of identifying an appropriate coarse-graining strategy.
+The definition includes two aspects: firstly, determining whether the system has the capacity to generate causal emergence, and secondly, assessing the occurrence of causal emergence given a specific macroscopic feature.
+Regarding a system’s capacity to exhibit causal emergence, this definition establishes a connection between causal emergence and synergistic relationships among variables across different time points.
+Consequently, a system denoted as $X_{t}$ is said to possess the capacity for causally emergent features if and only if:
 $Syn(X_{t};X_{t+1})>0$ (34)
-This means that causal emergence is regarded as the synergy effect among variables between the previous and subsequent moments in the system with Markovian dynamics.
-Then, Rosas further classify causal emergence by the $\phi$ID framework.
-Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ by the $\phi$ID, here are four information atoms corresponding to synergistic effect, which is regarded as causal emergence, are $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$.
-Since $\phi$ID includes forward and backward PID, he further divides causal emergence into downward causality and causal decoupling according to the different characteristics of those information atoms.
-The downward causality is defined by $\mathcal{D}(X_{t},X_{t+1})$, which encapsulating the information atoms that manifest a synergistic effect exclusively in the forward PID:
-$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (32)
-Conversely, the causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
+In this context, causal emergence is understood as the synergistic effect between variables at preceding and subsequent moments within a Markovian dynamics system.
+Then, Rosas further divided causal emergence into two parts in the $\phi$ID framework, downward causality and causal decoupling, based on the distinct characteristics of the information atoms.
+Among the sixteen $\phi$ID atoms obtained by decomposing the mutual information $I(X_{t};X_{t+1})$ using $\phi$ID, there are four information atoms corresponding to the synergistic effect, which is regarded as the composition of causal emergence.
+These atoms are denoted as $I^{\{12\}\rightarrow\alpha}_{\partial}(X_{t},X_{t+1})$, $\alpha\in\mathcal{A}=\{\{\{1\}\{2\}\},\{1\},\{2\},\{12\}\}$.
+Downward causality is denoted as $\mathcal{D}(X_{t},X_{t+1})$, encapsulating the information atoms that exclusively manifest a synergistic effect in forward PID, as shown in Figure 1213:
+$\mathcal{D}(X_{t},X_{t+1}):=\sum_{\begin{subarray}{c}\alpha\in A/\{12\}\end{subarray}}{I^{\{12\}\rightarrow\alpha}_{\partial}}(X_{t},X_{t+1})$ (35)
+While causal decoupling is denoted as $\mathcal{G}(X_{t},X_{t+1})$, pinpointing the specific information atom where both forward and backward PID exhibit synergy:
 $\mathcal{G}(X_{t},X_{t+1}):={I^{\{12\}\rightarrow\{12\}}_{\partial}}(X_{t},X_{t+1})$ (36)
-As the $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that
+Since $\phi$ID is a non-overlapping decomposition of all information, this classification takes into account all cases of causal emergence, that is,
 $Syn(X_{t};X_{t+1})=\mathcal{G}(X_{t},X_{t+1})+\mathcal{D}(X_{t},X_{t+1}).$ (37)
-Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) represented in double redundancy lattice.
-Rosas also gives out a version of causal emergence quantification depending on a macro-variable, i.e., a coarse-graining strategy.
-If a system has the capacity to generate causal emergence, then there can be some macroscopic features that exhibit causal emergence.
-A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ if the complete state $X$ of the system at time $t$ is known with perfect precision, which is equivalent to $V_{t}$ is statistically independent of $X_{t+1}$ when $X_{t}$ is given.
+Figure 12: Causal decoupling ($\mathcal{G}$) and downward causation ($\mathcal{D}$) are represented by a double-redundancy lattice.
+Distinct colors of dots are employed to distinguish the source elements of $X_{t}$, (colors are corresponding to Figure 11).
+In addition to that, Rosas also provided an approach to quantify the causal emergence of a specific macro-variable, i.e., a coarse-graining strategy.
+If a system has the capacity to generate causal emergence, there can be some macroscopic features that exhibit causal emergence.
+A feature variable $V$ is said to be supervenient on the underlying system if it does not provide any predictive power for future states at times $t+1$ when the complete state $X$ of the system at time $t$ is known with perfect precision.
+This is equivalent to $V_{t}$ being statistically independent of $X_{t+1}$ given $X_{t}$.
 Then, for a system described by $X_{t}$, a supervenient feature $V_{t}$ is said to exhibit causal causation if:
@@
 For this definition, the system’s capacity of causal emergence is required, where $Syn(X_{t};X_{t+1})$ > 0, since $Un(V_{t};X_{t+1}|X_{t})\leq Syn(X_{t};X_{t+1})$ holds for any supervenient feature $V_{t}$.
-Corresponding to the classification of system’s capacity, the Downward Causation exists of a feature variable $V$ when:
+Corresponding to the classification of the system’s capacity, the downward causation (as indicated by the red dotted line in Figure 13) of a feature variable $V$ exists when
 $Un(V_{t};$ $X_{t+1}^{1}|X_{t})>0$ (39)
@@
 $Un(V_{t};$ $X_{t+1}^{2}|X_{t})>0$
-and the Causal Decoupling exists when
+and causal decoupling (as indicated by the blue dotted line in Figure 13) exists when
 $Un(V_{t};V_{t+1}|X_{t},X_{t+1})>0,$ (40)
-which also needs to be premised on the capacity of the system.
+which also depends on the capacity of the system.
 Furthermore, $V_{t}$ is said to have pure causal decoupling if $Un(V_{t};X_{t+1}^{\alpha}|X_{t})=0$ and $Un(V_{t};X_{t+1}|X_{t})>0$.
-If all the emergent features exhibit pure causal decoupling, then the system is said to be perfectly decoupled.
+If all the emergent features exhibit pure causal decoupling, the system is said to be perfectly decoupled.
 Figure 13: Synergistic causes can be further divided into causal decoupling and downward causation.
-Although a rigorous quantitative definition for causal emergence is proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems.
-Also, there are many calculation methods in the PID framework, which makes the quantitative definition determined by the specific quantitative method when applied.
-Therefore, to solve those problems, Rosas also relaxed the calculation methods of causal emergence, thus obtaining sufficient conditions for the two phenomena as identification criteria.
-Specifically, in order to avoid exploring the specific quantitative relationship of synergistic and redundant information, this criteria repeatedly subtracts redundant information to make positive results a sufficient condition for emergent phenomena.
-Although this makes the quantitative method lose some of its generalizability, it improves the reliability of the results.
+Although a rigorous quantitative definition for causal emergence was proposed, the mathematical formulations used in $\phi$ID can be complex and computationally demanding, making it difficult to apply the method to real-world systems.
+Also, In addition, the inconsistency in PID calculation causes the definition of causal emergence to depend on a specific PID calculation.
+To address these problems, Rosas relaxed the calculation of causal emergence and established a identification criteria based one the sufficient conditions of causal decoupling and downward causation.
+Specifically, to avoid exploring the specific quantitative methods of synergistic and redundant information, this criteria repeatedly subtracts redundant information, making the result a sufficient condition for causal emergence, which lose some of its generality but improves the reliability.
 The three indicators to be used are:
 $\Psi_{t,t+1}(V):={I(V_{t};V_{t+1})}-\sum_{\begin{subarray}{c}j\end{subarray}}I(X_{t}^{j};V_{t+1}),$ (41)
-which measures the mutual information between macro-variables across two time steps minus the information between micro-state to macro-state, where $V$ is a predefined macro-variable, but how to find such macro-variable is never discussed in Rosas et al.
+which measures the mutual information between macro-variables across two time steps minus the mutual information between each micro-state and macro-state.
+$\Delta_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}\left({I(V_{t};X_{t+1}^{j})}-\sum_{\begin{subarray}{c}i\end{subarray}}I(X_{t}^{i};X_{t+1}^{j}))\right),$ (42)
+which is the maximum of the difference between mutual information of $V_{t}$ and $X^{j}_{t+1}$ and the summation of the mutual information of $X_{t}^{i}$ and $X_{t+1}^{j}$.
+$\Gamma_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}{I(V_{t};X_{t+1}^{j})},$ (43)
+which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$.
+For the above indicators, $V$ is a predefined macro-variable, but the specific method for finding such a macro-variable is not discussed in Rosas et al.
 (2020).
-$\Delta_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}\left({I(V_{t};X_{t+1}^{j})}-\sum_{\begin{subarray}{c}i\end{subarray}}I(X_{t}^{i};X_{t+1}^{j}))\right),$ (39)
-which is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$ in next step minus all the summation of mutual information between $X_{t}^{i}$ and $X_{t+1}^{j}$.
-$\Gamma_{t,t+1}(V):=\max_{\begin{subarray}{c}j\end{subarray}}{I(V_{t};X_{t+1}^{j})},$ (40)
-this is the maximum mutual information between $V_{t}$ and $X^{j}_{t+1}$
-The specific functions of the indicators are as follows: $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$.
-Similarly, $\Delta_{t,t+1}(V)>0$ is as sufficient condition for $V_{t}$ to show downward causation.
-Finally, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ is a sufficient condition for causal decoupling to occur.
-In summary, Rosas proposes an approach to quantitatively describe causal emergence by establishing the relationship between causal emergence and the synergistic effect of variables in different times, and further classified causal emergence by $\phi$ID.
-The definition not only objectively judges whether the system has the capacity to produce causal emergence, but can also measure the characteristics of causal emergence of specific macro features.
-The most notable contributions of his work were to bring the study of emergence closer to quantitative and empirical scientific research, to classify types of causal emergence, and to complement philosophical discussions about emergence.
-There are some applications of PID or $\phi$ID.
-Recently, Luppi et al.
-apply $\phi$ID method on brain BOLD signals Luppi et al.
-(2022) to identify the synergistic core of the brain.
-They found synergistic information is for the integration of different parts to a whole, and redundancy is for robustness.
-With the development of partial information decomposition techniques, we can further decompose the mutual information between two variables.
-This may help us to reveal the properties of systems with more aspects.
-For example, in  Varley and Hoel (2022), the authors first decompose the mutual information of a system, and calculate an indicator of synergy bias according to partial information decomposition method Williams and Beer (2010) to characterize how synergistic information distribute on different levels of the system.
-Higher synergy bias indicates that more partial information is in synergistic relationships.
-Then, the authors found that in some systems with occurrence of causal emergence, if the system is coarsen to a reduced system, the synergy bias will increase.
-Therefore, partial information is converted from redundancy to synergy as we coarse-grain the system.
-The conclusion is emergence is a kind of information conversion.
+The specific usage of the indicators are as follows: Firstly, $\Psi_{t,t+1}(V)>0$ is a sufficient condition for the causal emergence of $V_{t}$.
+Secondly, $\Delta_{t,t+1}(V)>0$ is a sufficient condition for $V_{t}$ to show downward causation.
+Thirdly, $\Psi_{t,t+1}(V)>0$ and $\Gamma_{t,t+1}(V)=0$ together constitute a sufficient condition for causal decoupling.
+In summary, Rosas proposed an approach to quantitatively characterize and classified causal emergence based on $\phi$ID by establishing the relationship between causal emergence and the synergistic effect of variables at different time points and further categorized causal emergence.
+The definition not only provides an objective assessment of a system’s capacity for causal emergence but also enables the measurement the causal emergence associated with a specific macro-feature.
+His significant contributions include bridging the gap between the study of causal emergence and quantitative empirical research, classifying the different types of causal emergence, and complementing philosophical discussions on the topic.
+The PID and $\phi$ID frameworks hold the potential for explaining data in various applications.
+For instance, Luppi et al.
+recently employed the $\phi$ID method to analyze brain BOLD signals Luppi et al.
+(2022), aiming to identify the synergistic core of the brain.
+Their findings revealed that synergistic information facilitates the integration of different brain regions, whereas redundancy contributes to robustness.
+This study offers valuable insights into understanding the underlying mechanisms of the brain.
+The advancement of partial information decomposition techniques allows for further analysis of the mutual information between variables, enabling a deeper understanding of system properties from multiple perspectives.
+In a study by Varley et al.
+Varley and Hoel (2022), the authors applied partial information decomposition to decompose the mutual information of a system.
+They calculated an indicator of synergy bias to assess how synergistic information is distributed across different levels of the system using the method proposed by Williams and Beer Williams and Beer (2010).
+A higher synergy bias indicated a greater amount of partial information involved in synergistic relationships.
+Subsequently, the authors observed that in certain systems exhibiting causal emergence, when the system is simplified or reduced, the synergy bias increases.
+This suggests that as we coarse-grain the system, partial information undergoes a transformation from redundancy to synergy.
+The overall conclusion drawn was that emergence can be understood as a form of information conversion.
+3.6.5 Causal Emergence Identification from Data
+In the previous sections, we introduced several works on quantifying emergence through causality and other information-theoretic concepts.
+All of these works tried to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence.
+However, in practical applications, for a theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time-series data of dynamic systems, and provide explanations for the results.
+The first method for causal emergence identification was introduced in Rosas’s paper on causal emergence Rosas et al.
+(2020) and involves the three indicators defined by Equations (41)–(43), as mentioned in Section 3.6.4.
+The identification criteria were exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
+Although these three indicators avoid the problem of redundant information calculation, it is important to note that they serve as sufficient conditions, rather than as definitive proof of emergence.
+In other words, indicators greater than 0 can suggest the presence of emergence, but indicators less than 0 do not necessarily imply the absence of emergence.
+The construction of this indicator faces challenges in identifying emergence in systems with significant redundant information or a large number of variables, which is often the case in many real-world systems.
+Additionally, a limitation of this method is the requirement of a predetermined coarse-grained variable $V$, and different choices of this variable can significantly impact the results.
+As a result, the development of an automatic coarse-graining strategy based on data remains an unresolved issue.
+As a result, the current theoretical frameworks for causal emergence lack a practical and effective identification algorithm.
+Although previous studies have proposed methods based on static network structures and approximations for information decomposition, there is still a need for a comprehensive approach that can be applied to general Markovian dynamic systems.
+One of the main challenges is the necessity to search for all possible functions of coarse-graining or decomposing subsets within the data to identify causal emergence in complex systems.
+However, conventional numerical methods are unable to handle the computational costs associated with such an extensive search in a vast functional space.
+Therefore, the development of new methods is imperative to address these issues.
+In Section 4, we explore the application of machine learning techniques to address the challenge of identifying causal emergence within time-series data.
 3.6.6 Comparison with Hoel’s Framework
-When comparing Hoel’s framework to Rosas’ quantification framework of causal emergence, several clear advantages can be observed in the latter.
-Firstly, Rosas’ theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal.
+When comparing Hoel’s framework to Rosas’s quantification framework for causal emergence, several clear advantages can be observed in the latter.
+Firstly, Rosas’s theory does not require a predetermined coarse-graining method, making it more mathematically rigorous and formal.
 Secondly, it offers a detailed decomposition of causal emergence, specifically downward causation and causal decoupling.
@@
 Firstly, in order to obtain the full information lattice, a systematic iteration of all variable compositions is required.
-Additionally, despite the use of formula 35, it is necessary to define a macro-variable.
+Additionally, despite the use of Formula (38), it is necessary to define a macro-variable.
 Unfortunately, the authors do not provide any method to identify such a variable.
@@
 It is crucial to discuss how to incorporate causal elements, such as intervention and counterfactuals, into the framework.
-Finally, the issue of identifying whether causal emergence occurs in a system based on the given time series data of its behaviors has not been addressed in the preceding discussion.
-To tackle this problem, the application of emerging technologies in machine learning and artificial intelligence is required.
+Finally, the issue of identifying whether causal emergence occurs in a system based on the given time-series data of its behaviors has not been addressed in the preceding discussion.
+To address this problem, the application of emerging technologies in machine learning and artificial intelligence is required.
 These technologies can provide valuable tools and techniques for detecting and analyzing causal emergence.
-4 Causal Emergence and Artificial Intelligence
-Recently, newly emerging artificial intelligence technologies have made significant breakthroughs in tackling a range of important and challenging problems.
+4 Causal Emergence and Machine Learning
+Recently, newly emerging machine learning technologies have made significant breakthroughs in addressing a range of important and challenging problems Silver et al.
+(2016); Dabney et al.
+(2020); Senior et al.
+(2020); LeCun et al.
+(2015); Wang et al.
+(2018); Mikolov et al.
+(2013); Devlin et al.
+(2018); Vaswani et al.
+(2017); Mnih et al.
+(2015).
 Examples include defeating human champions in complex games like Go Silver et al.
 (2016); Dabney et al.
-(2020), and predicting the intricate structures of protein folding Senior et al.
+(2020), predicting the intricate structures of protein folding Senior et al.
+(2020), and generating human language using large language models Fan et al.
 (2023).
-These achievements have been made possible through the application of machine learning methods, which leverages well-designed neural network architectures and automatic differentiation techniques.
-Machine learning methods, with their diverse and well-designed neural network structures, along with automatic differentiation techniques, have demonstrated the ability to efficiently approximate complex functions within a vast functional space.
-Given this capability, it raises the question of whether we can leverage these emerging techniques to address the challenging problem of finding effective coarse-graining strategies or optimal decomposition methods for complex systems using time series data?
-The answer is yes.
-We will introduce two latest works on causal emergence identification from data based on neural networks.
-Interestingly, these works indicate that causal emergence identification problem can also provide us with insight into connections with another related field, causal representation (machine) learning although causal emergence identification and causal representation learning are totally different research areas for distinct problems from appearance, they share commonalities.
-Furthermore, we also review the works on causal model abstractions and reinforcement learning based on world models as extensions because similar machine learning architecture are shared by these machine learning frameworks.
-4.1 Causal Emergence Identification
-In the previous section we have introduced several works on quantifying emergence by causality and other information theoretic conceptions.
-Among which, all works try to propose quantitative measures, conceptual frameworks, and numeric examples based on Markovian dynamics for causal emergence.
-However, in practical applications, in order for the theoretical framework to be implemented, we need to automatically identify causal emergence from real data, especially time series data of dynamic systems, and give explanations for the results.
-4.1.1 Rosas’ Identification Method
-The first method for causal emergence identification is given by Rosas’ paper in causal emergence Rosas et al.
+These achievements have been made possible through the application of machine learning methods, which leverage well-designed neural network architectures and automatic differentiation techniques.
+However, there is a crucial limitation of conventional machine learning: it can only capture the information or associations within the data, without being able to uncover the underlying causal relationships.
+According to the theory of causal hierarchy by Judea Pearl Pearl et al.
+(2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship that is invariant in different environments, whereas the latter may be more dependent on contexts and reflects the limitation of the data.
+Therefore, it is necessary to develop new machine learning frameworks to incorporate the consideration of causalitySchölkopf et al.
+(2021); Kaddour et al.
+(2022); Cui and Athey (2022).
+Numerous studies have provided evidence to show that incorporating aspects of causality can lead to improved performance in machine learning tasks, including out-of-distribution generalization, adaptation to diverse environments, and handling interventions Peters et al.
+(2016); Cui and Athey (2022); Kuang et al.
 (2018).
-Actually, the three indicators, the Equations 38, 39 and 40 as mentioned in section 3.8.4.
-The identification criteria have been exemplified in three case studies, leading to the following conclusions: particle collisions emerge as a distinctive feature within Conway’s Game of Life, flock dynamics emerge as a characteristic feature in simulated bird behavior, and the representation of motor behavior in the cortex emerges from neural activity.
-Although these three indicators avoid the problem of redundant information calculation, they are all sufficient conditions, that is, indicators greater than 0 can indicate emergence.
-The construction of this indicator repeatedly cuts out redundant information, which makes it almost impossible to identify the emergence of this indicator in systems with large redundant information or systems with many variables, and this is the case for most of the systems we face.
-Another defect of this method is that the coarse-grained variable $V$ must be given in prior, and different $V$ affects the results in a large degree.
-Therefore, the automatic coarse-graining strategy design according to data is still remain unsolved.
-New methods are required.
-Therefore, the current theoretical frameworks of causal emergence do not give a good workable identification algorithm.
-Although existing works have tried to propose some workable methods based on static network structures and approximation for information decomposition, a general method for general Markovian dynamic systems is still lacking.
-One of the important reasons is to identify causal emergence within data needs to search for all the functions of coarse-graining or decomposing subsets for a complex system.
-However, the conventional numeric methods can not afford the computational cost for such search in a huge functional space.
-New methods are required.
+Machine learning and causal inference are also connected to causal emergence in two aspects, which we call “causal emergence with machine learning” and “causal emergence for machine learning”.
+On one hand, machine learning can be employed to address the challenge of identifying causal emergence from data, referred to as “causal emergence with machine learning”.
+On the other hand, causal emergence theory and effective information (EI) measures have potential applications in the field of machine learning, known as “causal emergence for machine learning”.
+However, the extension depends on a deeper understanding of how causal emergence theory is related to causality.
+4.1 Causal Emergence with Machine Learning
+In Section 3.6.5, we discussed the challenge of identifying causal emergence.
+Although Rosas et al.
+presented a method to address this issue by calculating an approximate measure of causal emergence based on their theoretical framework (Equation (41)), the method requires the availability of a macro-variable, denoted as $V_{t}$, which needs to be defined in advance.
+However, defining this variable in a general sense when confronted with data remains challenging.
+Hence, the question arises of whether we can leverage machine learning to automatically learn the macro-variable.
+In this section, we explore two recent approaches that utilize machine learning and neural network techniques to address the problem of causal emergence identification from the perspective of Heol et al.’s theoretical framework for EI maximization.
+These methods aim to learn the coarse-graining strategy, macro-variable, and macro-dynamics directly from the data, eliminating the need for explicit definitions.
+By leveraging the power of machine learning, we can potentially overcome the limitations of manual variable definition and enable a more automated and data-driven approach to identifying causal emergence.
 4.1.1 Neural Information Squeezer (NIS)
-Here we introduce one of the latest works on this direction which is just published in this special issue.
-In this paper, the authors formulate the problem of causal emergence identification on continuous Markov dynamics.
-Suppose the time series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
+The first work we introduce has just been published in this special issue.
+In this paper, the authors formulated the problem of causal emergence identification in continuous Markov dynamics.
+Suppose the time-series data $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\dots,\boldsymbol{x}_{T}$ are generated by a stochastic dynamical system continuously, which can be described by a differential equation,
 $\frac{d\boldsymbol{x}}{dt}=g\left(\boldsymbol{x}(t),\xi\right),$ (44)
 where $\xi$ is a Gaussian noise.
-The time series $\{\boldsymbol{x}_{t}\}$ are the observations of the micro-states of the system, and the problem that we confront is identifying if the causal emergence occurs for the original system.
-And we also want to know the appropriate coarse-graining strategies, the macro dynamics which correspond to the coarse-graining method, and the dimension of the phase space on macro-level if the causal emergence occurs.
-To solve the problem, the authors propose a mathematical framework by converting the original issue into an optimization problem for functions.
-That is, we can find the appropriate coarse-graining strategy $\phi_{q}$, which is defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where, $q<p$ is the dimension of the macro-states, and the macro-level Markov dynamics $f$, which is defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$, the dimension averaged EI of the macro-dynamics (see Equation 24), that is,
+The time series $\{\boldsymbol{x}_{t}\}$ represents the observations of the micro-states of the system.
+The problem we confront is identifying whether causal emergence occurs for the original system.
+We also want to know the appropriate coarse-graining strategies, the macro-dynamics corresponding to the chosen coarse-graining method, and the dimension of the phase space at the macro-level when the causal emergence occurs.
+To address this problem, the authors proposed a mathematical framework by converting the original issue into an optimization problem for functions.
+Specifically, the goal is to find the appropriate coarse-graining strategy $\phi_{q}$, defined on $\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$, where $q<p$ is the dimension of the macro-states.
+Additionally, the aim is to determine the macro-level Markov dynamics $f$, defined on $\mathcal{R}^{q}\rightarrow\mathcal{R}^{q}$ with the random noise $\zeta$, by maximizing $\mathcal{J}$—the dimension-averaged EI of the macro-dynamics (see Equation (27)).
+This is expressed as,
 $\max\mathcal{J}(\hat{f}_{\phi_{q}}),\qquad$ (45)
-However, this problem has a trivial solution, e.g.
-$\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity.
-Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy within all $q$ dimensional functions.
+However, this problem has a trivial solution, e.g., $\phi$ is a constant map for all micro-states, and the macro-dynamic is an identity.
+Surprisingly, this solution has a large EI because the identity map has the largest determinism and lowest degeneracy among all $q$-dimensional functions.
 This is trivial because too much information has been abandoned by the coarse-graining strategy such that the macro-dynamic is useless.
-To conquer the problem, the authors add a reasonable constraint to the original optimization framework, which is called effectiveness requirement.
-A coarse-graining strategy and the corresponding macro-dynamic are effective, which means there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state instead of the macro-state in the next time step according to the state in the previous step, the constraint can be written with,
+To address the problem, the authors added a reasonable constraint to the original optimization framework, called the effectiveness requirement.
+A coarse-graining strategy and its corresponding macro-dynamics are considered effective if there exists another decoarse-graining function $\phi^{\dagger}$ such that the functions $\phi$, $f$, and $\phi^{\dagger}$ can predict the micro-state, instead of the macro-state, in the next time step based on the state in the previous step.
+The constraint can be written as,
 $||\phi^{\dagger}(f(\phi(\boldsymbol{x}_{t})))-\boldsymbol{x}_{t+1}||<\epsilon,$ (46)
-for all time step $t$, where, $\epsilon$ is a given constant.
-In this way, an effective coarse-graining strategy and the macro-dynamic can not only maximize the EI but also reproduce the original micro-dynamic as much as it can.
-This constraint is an important complimentary for the framework of EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states which can solve the problem of ambiguity of macro-states.
-This constraint is also a guarantee of commutativity.
-To see this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ on both sides of Equation 43 to obtain $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
+for all time steps $t$, where $\epsilon$ is a given constant.
+This way, an effective coarse-graining strategy and macro-dynamics not only maximize the EI but also reproduce the original micro-dynamics as much as possible.
+This constraint is an important complement to the framework for EI maximization because the decoarse-graining function $\phi^{{\dagger}}$ can map macro-states back to micro-states, addressing the problem of the ambiguity of macro-states.
+This constraint also ensures commutativity.
+To demonstrate this, we can apply $(\phi^{{\dagger}})^{-1}\approx\phi$ to both sides of Equation (46), resulting in $||f(\phi(x_{t}))-\phi(x_{t+1})||<\phi(\epsilon)$.
 However, the problem is still hard to solve.
-The paper Zhang and Liu (2022) proposes a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step, and maximize EI for different $q$ in the second step.
-The first step can be solved by training a neural network easily, and the second step converts the complex functional optimization problem into a line search in one dimensional space.
+In Zhang and Liu (2022), the authors proposed a two-stage method, which can minimize the prediction error under a given dimension $q$ of macro-states in the first step and maximize EI for different values of $q$ in the second step.
+The first step can be solved by training a neural network, and the second step converts the complex functional optimization problem into a line search in one-dimensional space.
+To realize the abstract mathematical framework, the authors proposed an encoder-dynamics learner-decoder framework for neural networks, named the Neural Information Squeezer (NIS), vividly describing the basic working principle of the framework as depicted in Figure 14.
 Figure 14: The architecture and workflow of the Neural Information Squeezer (NIS).
-NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$.
+The NIS consists of three main components: an encoder $\phi$, a dynamics learner $f$, and a decoder $\phi^{{\dagger}}$.
 The encoder $\phi$ comprises an invertible neural network $\psi$ and a projection operator $\chi$, which simply discards fixed dimensions output by $\phi$.
@@
 In cases where the input is incomplete, it can be complemented using Gaussian random noise.
-To realize the abstract mathematical framework, the authors propose an encoder-dynamics learner-decoder framework of neural network, named neural information squeezer (NIS), which describes the basic working principle of the framework vividly.
-To reduce the complexity of the encoder-decoder which corresponds to $\phi$ and $\phi^{\dagger}$, respectively, the authors use an invertible neural network to implement the encoder, and invert the input-output workflow of the same neural network to realize the computation of the decoder.
-Concretely, the encoding process is decomposed into two step computation, information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on $p$ dimensional input space, and information discarding with the simple projection operation, that is:
+In order to reduce the complexity of the encoder and decoder, corresponding to $\phi$ and $\phi^{\dagger}$, respectively, the authors used an invertible neural network to implement the encoder and inverted the input-output workflow of the same neural network to implement the decoder.
+Concretely, the encoding process was decomposed into two computation steps: information conversion with the invertible neural network $\psi:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ on a $p$-dimensional input space, and information discarding with a simple projection operation.
+This is expressed as:
 $\phi(\boldsymbol{x})=\chi_{q}(\psi(\boldsymbol{x})),$ (47)
-where $\chi_{q}$ is the projection operator to project a $p>q$ dimensional vector into a $q$ dimensional vector with the dimensions from $1$ to $q$ left.
-The decoding process also contains two stages, random sampling a Gaussian noise to complete the details of macro states, and information conversion but in an inverted way, that is:
+where $\chi_{q}$ is the projection operator that projects a $p>q$-dimensional vector into a $q$-dimensional vector, retaining dimensions from $1$ to $q$.
+The decoding process also consists of two stages: randomly sampling Gaussian noise to complete the details of macro-states, and information conversion but in an inverted way, that is,
 $\phi^{{\dagger}}(\boldsymbol{x})=\psi^{-1}(\boldsymbol{x}\bigoplus\zeta),$ (48)
-where, $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$ dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.
+where $\zeta\sim\mathcal{N}(0,I_{p-q})$ is a $p-q$-dimensional standard Gaussian random vector, and $\bigoplus$ is the operation of vector concatenation.
 The adoption of an invertible neural network in the framework serves two main purposes.
@@
 These properties enable researchers to gain insights and understand the characteristics of the framework more easily.
-This architectural design with invertible neural network can not only reduce the complexity of computation, but can also separate the input information from the micro-states into noise and EI for prediction clearly.
-The authors proved a mathematical theorem stating that, as the neural network is converged by training, the mutual information on the adjacent macro states approaches to the mutual information of the two adjacent-step micro states in the data.
-This means, all the information dropped by the encoder is almost irrelevant with the prediction.
-All the useful information is squeezed into the information bottleneck of the dynamics learner at the macro-level.
-And another theorem has proved that the degree of information squeezing for micro-states is increased as we reduce the dimension of the macro states, and thus more and more useless information in the original micro-state data is thrown out by the narrower and narrower information bottleneck, the macro-state dynamics learner.
-To solve the problem of causal emergence identification, we first optimize the neural networks, and the dimension averaged EI, $\mathcal{J}_{q}$ can be obtained for different $q$.
-When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension averaged EI for the micro-dynamics.
+This architectural design with an invertible neural network not only reduces the complexity of computation but also separates the input information from the micro-states into noise and EI for prediction.
+The authors proved a mathematical theorem stating that, as the neural network converges through training, the mutual information between adjacent macro-states tends to approach the mutual information of two adjacent-step micro-states in the data.
+This means that all the information discarded by the encoder is almost irrelevant to the prediction.
+All the useful information converges into the information bottleneck within the macro-level dynamics learner.
+Another theorem demonstrates that the degree of information compression for micro-states increases as we reduce the dimension of the macro-states.
+Thus, the macro-state dynamics learner, acting as a narrower information bottleneck, progressively discards more and more extraneous information from the original micro-state data.
+To address the problem of causal emergence identification, we first optimize the neural networks, and the dimension-averaged EI, $\mathcal{J}_{q}$, can be obtained for different $q$.
+When $q=p$, the derived $\mathcal{J}_{p}$ is the dimension-averaged EI for the micro-dynamics.
 Finally, the calculation of
 $\Delta\mathcal{J}=\mathcal{J}_{q}-\mathcal{J}_{p}$ (49)
-is the measure of dimension averaged causal emergence (see Equation 25. If $\Delta\mathcal{J}>0$, we can judge that causal emergence occurs.
-Numeric examples are given to show the effectiveness of the whole framework.
-Interestingly, the neural information squeezer can work on the time series data with discrete micro-state although the whole framework is built based on continuous dynamical systems.
-Particularly, the model can work out the same coarse-graining strategy and macro-level dynamic as in the example of boolean network with four nodes in Hoel’s original paper, although any information about the node grouping and state mapping is not given to the framework.
-It is also interesting to compare NIS with the mentioned frameworks or models in previous sections.
-Compare to the theory of computational mechanics, NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as the states, and when the whole framework is well-trained such that it can predict the future micro-states precisely, the encoded macro-states converge to effective states which can be treated as causal states in computational mechanics.
-However, the objective that maximize the effective information (EI) in NIS has no correspondence.
-NIS also has similarity with G-emergence because it also adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step.
+is the measure of dimension-averaged causal emergence (see Equation (28). If $\Delta\mathcal{J}>0$, we can state that causal emergence occurs.
+Numeric examples are provided to show the effectiveness of the whole framework.
+Interestingly, the Neural Information Squeezer can handle time-series data with discrete micro-states, even though the whole framework was initially designed for continuous dynamical systems.
+Particularly, the model determines the same coarse-graining strategy and macro-level dynamic as illustrated in the example of a boolean network with four nodes in Hoel’s original paper, although no information about node grouping and state mapping was provided in the framework.
+It is also interesting to compare the NIS with the frameworks and models mentioned in previous sections.
+Compared to the theory of computational mechanics, the NIS can be treated as a kind of $\epsilon$-machine because any encoded macro-states can be understood as states.
+When the whole framework is well trained, such that it can precisely predict future micro-states, the encoded macro-states converge to effective states, which can be treated as causal states in computational mechanics.
+However, the objective that maximizes the effective information (EI) in the NIS has no correspondence.
+The NIS also shares a similarity with G-emergence because it adopts the ideas of Granger causality: the effective macro-state is optimized by predicting the micro-state in the next time step.
 However, there are several obvious differences between the two frameworks.
-In theory of G-emergence, the macro-state must be selected by hand at first, however, it is optimized automatically.
-Second, NIS uses neural networks to predict the future states, while, G-emergence uses auto-regressive techniques to fit the data.
+In the theory of G-emergence, the macro-state must initially be selected manually, although it is optimized automatically.
+In addition, the NIS uses neural networks to predict future states, whereas G-emergence uses auto-regressive techniques to fit the data.
 Due to the universal approximate theorem (see Hornik (1991); Teshima et al.
-(2020)), neural networks are more advantage.
-NIS provides a solution for causal emergence identification from data, and it is the first time to solve the problem of the automatic construction of coarse-graining strategy and the macro-dynamics from data for causal emergence identification.
-Another important benefit of NIS which is to be noticed is that, what NIS has solved actually is a generalized causal emergence identification problem but not limited on a justification.
-Because after training, NIS can not only identify if causal emergence occurs, but also can obtain a neural network for coarse-graining the data and a neural network to simulate macro-dynamics.
-This may provide us richer information than just a justification.
+(2020)), neural networks are more advantageous.
+The NIS provides a solution for causal emergence identification from data, and it is the first solution for addressing the problem of the automatic construction of the coarse-graining strategy and macro-dynamics from data for causal emergence identification.
+Another important benefit of the NIS is that it essentially addresses a generalized causal emergence identification problem, extending beyond mere justification.
+This is because following training, the NIS not only identifies whether causal emergence occurs but also obtains a neural network for coarse-graining the data and another for simulating the macro-dynamics.
+This can provide us with richer information compared to a mere justification.
 4.1.2 Neural Information Squeezer Plus (NIS+)
-One of the biggest problem of NIS is that it does not directly solve the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner).
-Instead, it searchs for the optimal dimension of the macro-state space $q$ such that EI can be maximized.
-Therefore, for given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimize Equation 43) but not for EI.
-Second, NIS can not support complex computations such as the coarse-graining functions containing multiple steps of information conversion and discarding.
-To tackle these problems, NIS+ is proposed in ref Yang et al.
+One of the biggest problems of the NIS is that it does not directly address the EI maximization problem by optimizing the neural networks (encoder, decoder, and dynamics learner).
+Instead, it searches for the optimal dimension of the macro-state space $q$ such that EI can be maximized.
+Therefore, for a given $q$, the neural networks are optimal for reconstructing the micro-dynamics (minimizing Equation (46)) but not for EI.
+In addition, the NIS cannot support complex computations such as coarse-graining functions involving multiple steps of information conversion and discarding.
+To address these problems, the NIS+ was proposed, as discussed in Yang et al.
 (2023).
-Mathematically, the problem of EI maximization can be converted into a machine learning problem according to the definition of EI and the variational inequality and probability reweighting technique.
-After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},g_{\theta^{\prime}}$ with parameters $\omega,\theta$ and $\theta^{\prime}$, respectively.
-Formally, the problem of minimization of loss function problem without constraints can be written:
+Mathematically, the problem of EI maximization can be transformed into a machine learning problem based on the definition of EI and the variational inequality and probability reweighting technique.
+After this conversion, the minimization problem can be solved by training three neural networks $\psi_{\omega},f_{\theta},$ and $g_{\theta^{\prime}}$ with parameters $\omega,\theta$, and $\theta^{\prime}$, respectively.
+Formally, the minimization problem without constraints can be written as:
 $\min_{\omega,\theta,\theta^{\prime}}\sum_{t=1}^{T-1}w(\boldsymbol{x}_{t})||\boldsymbol{y}_{t}-g_{\theta^{\prime}}(\boldsymbol{y}_{t+1})||+\lambda||\hat{\boldsymbol{x}}_{t+1}-\boldsymbol{x}_{t+1}||,$ (50)
 where $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t}))$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})=Proj_{q}(\psi_{\omega}(\boldsymbol{x}_{t+1}))$ are the macro-states.
-$\lambda$ is a Lagrangian multiplier which will be taken as a hyper-parameter in experiments.
-$w(\boldsymbol{x}_{t})$ is the inverse probability weights which is defined as:
+$\lambda$ is a Lagrangian multiplier, which is regarded as a hyperparameter in experiments.
+$w(\boldsymbol{x}_{t})$ is the inverse probability weight, which is defined as:
 $w(\boldsymbol{x}_{t})=\frac{\tilde{p}(\boldsymbol{y}_{t})}{p(\boldsymbol{y}_{t})}=\frac{\tilde{p}(\phi(\boldsymbol{x}_{t}))}{p(\phi(\boldsymbol{x}_{t}))},$ (51)
@@
 Consequently, the weight $w$ is computed as the ratio of these two distributions.
-Theorems are proved to guarantee the correctness of this conversion.
-The overall framework of NIS+ can be depicted in Figure 15.
-Figure 15: The workflow and architecture of our proposed framework, Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data.
+Theorems have been proven to ensure the correctness of this conversion.
+The overall framework for the NIS+ is depicted in Figure 15.
+Additionally, to apply the NIS+ to data generated by complex systems like multi-agent systems and cellular automata, the structures of the encoder can be extended by defining two kinds of combinations: stacking the basic encoder and concatenating the basic encoder.
+Here, a basic encoder is the composition of an invertible neural network and a projection operation.
+By inverting the directions of all the basic encoders and supplementing empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
+Serious numerical experiments have demonstrated that the NIS+ can automatically learn emergent dynamics and the coarse-graining strategy from data and can integrate these results to identify causal emergence.
+For instance, in experiments utilizing training data generated by introducing Gaussian noise to SIR dynamics, the NIS+ demonstrated superior performance compared to the NIS in recovering the vector field of derivatives that aligned with the actual macro-level SIR dynamics.
+To assess the ability of the NIS+ to learn emergent dynamics, the authors generated data based on macro-states using the SIR dynamics as the ground truth.
+However, the NIS+ cannot directly utilize this data; only micro-level noisy data can be employed.
+These micro-level data were generated by introducing Gaussian noise to the macro-states.
+In the experiments conducted with the Boid model, training data were generated using the renowned Boid model, which simulates the herding behavior of birds.
+Both the NIS+ and NIS demonstrated the ability to automatically learn macro-level states and dynamics.
+In this simulation, the authors divided all artificial birds (boids) into two distinct groups.
+Interestingly, the learned macro-states could also be categorized into two groups.
+Within each group, the predicted trajectories generated by the learned macro-dynamics of this group accurately followed the center of the boid groups.
+Figure 15: The workflow and architecture of our proposed framework, the Neural Information Squeezer Plus (NIS+), for discovering causal emergence within data.
 (a) Various forms of input data.
-(b) The framework of our proposed model, NIS+, which incorporates our previous model, NIS.
-The boxes represent functions or neural networks, and the arrow pointing to a cross represents the operation of information discarding.
-$\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, while $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state.
+(b) The framework for our proposed model, the NIS+, incorporates our previous model, the NIS.
+The boxes represent functions or neural networks, and an arrow pointing to a cross represents the operation of information discarding.
+$\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t+1}$ represent the observational data of micro-states, whereas $\hat{\boldsymbol{x}}_{t+1}$ represents the predicted micro-state.
 $\boldsymbol{y}_{t}=\phi(\boldsymbol{x}_{t})$ and $\boldsymbol{y}_{t+1}=\phi(\boldsymbol{x}_{t+1})$ represent the macro-states obtained by encoding the micro-states using the encoder.
@@
 The mathematical problems that each framework aims to solve are also illustrated in the figure.
-(c) The various output forms of NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
-Additionally, to apply NIS+ to the data generated by the complex systems of multi-agent systems and cellular automata, the structures of encoder can be extended by defining two kinds of combinations: one is to stack the basic encoder, the other is to concatenate the basic encoder.
-Here, a basic encoder is the composition of an invertible neural network and a projection operation.
-By inverting the directions of all the basic encoder, and supplements the empty inputs with Gaussian noise vectors, we can obtain the corresponding decoder.
-Numerical experiments demonstrate the effectiveness of NIS+ in capturing emergent behaviors and finding causal emergence in various simulations, including SIR dynamics, bird flocking (Boid model), and the Game of Life.
-Additionally, NIS+ is applied to find emergent features in real brain data from 830 subjects while watching the same movie.
-Experiments also demonstrate that the learned dynamical model exhibits improved performance in terms of out-of-distribution generalization compared to other models.
-4.2 Implications in Causal Representation Learning and Out-of Distribution Generalization
-We found that, as a framework of machine learning, neural information squeezer and plus can be regarded as a kind of framework for causal feature learning Chalupka et al.
-(2017); Spirtes et al.
-(2000); Chalupka et al.
-(2015) or causal representation learning Schölkopf et al.
+(c) The various output forms of the NIS+, which include the learned macro-dynamics, captured emergent patterns, functions of coarse-graining, and the degree of causal emergence.
+The ability to capture emergent patterns was verified through experiments conducted in the “Game of Life”.
+The NIS+ demonstrated the capability to automatically discover both static and dynamic patterns, such as the “glider”, within the learned latent space.
+Additionally, when clear emergent patterns appeared in the data, larger effective information (EI) was observed for the learned emergent macro-dynamics.
+Across all the training data generated through simulations, the NIS+ outperformed comparative models, including the NIS, variational autoencoders, and feed-forward neural networks, in experiments testing the model on data distributions that differed from the training data.
+The paper also presented results on real data.
+Two sets of real fMRI time-series data from human subjects were used.
+The first dataset, AOMIC ID1000 Snoek et al.
+(2021), contained fMRI scanning data collected while subjects watched the same movie clip.
+As a comparison, another fMRI dataset, AOMIC PIOP2 Snoek et al.
+(2021), containing resting-state data from 50 subjects, was also utilized.
+The primary training objective was to predict the fMRI signal at the next time step using the learned macro-dynamics and decoder while simultaneously maximizing the effective information (EI) of the macro-level dynamics.
+Ultimately, a one-dimensional macro-state was found to represent all 100-dimensional micro-state data.
+Through attribution analysis, the authors found that micro-level signals with larger weights contributing to the one-dimensional macro-state were located in areas associated with visual tasks.
+However, in the comparison dataset with resting-state data, seven-dimensional macro-states were found to represent the micro-states, and attribution experiments revealed that micro-states with larger weights were distributed across different brain areas.
+After the macro-dynamics and coarse-graining strategy were learned, the EIs at both the macro-level and micro-level could be calculated so that causal emergence could be quantified.
+The experiments showed that the measure of causal emergence depends on different types of noise.
+If noise was added to the observational data (micro-state), the degree of causal emergence increased with the level of noise.
+In contrast, if the noise was from the dynamical mechanism, i.e., added to macro-states, the degree of causal emergence decreased with the level of noise.
+This implies that coarse-graining can learn to eliminate noise in raw data so that better macro-dynamics with larger EI can emerge.
+However, if the noise is from the internal dynamics, it cannot be removed through coarse-graining.
+4.2 Why is Causal Emergence “Causal”?
+In the previous sections, we discussed two frameworks that utilize machine learning and neural network techniques to address the challenge of causal emergence identification.
+Both frameworks place significant emphasis on maximizing EI to learn the coarse-graining strategy and macro-dynamics.
+Notably, the NIS+ framework extends the NIS approach to achieve real EI maximization, whereas the NIS relies on adjusting the hyperparameter $q$ to optimize EI.
+Moreover, the NIS+ has demonstrated an additional capability to generalize data distributions that differ from the training data.
+This suggests that the NIS+ can learn an invariant causal mechanism that is independent of the input distribution shifts through EI maximization.
+However, before extending this method to other machine learning scenarios, an essential question needs to be addressed: Is there a genuine relationship between causal emergence, EI maximization, and causality itself?
+In the previous sections, we introduced two theoretical frameworks of causal emergence: Hoel’s framework and Rosas’s framework.
+In both frameworks, the measurement of causal emergence heavily relies on mutual information.
+However, it is crucial to acknowledge that mutual information primarily quantifies correlation rather than causality.
+So why do the authors assert that their theories pertain to “causal” emergence?
+The key point lies in the research objectives of these frameworks, which focus on a Markov dynamical system.
+In this context, the dynamical mechanism is assumed to be known, and confounding variables are not considered.
+Under this assumption, measures related to correlation can effectively indicate causality.
+While mutual information itself is not a direct measure of causality, it can serve as a proxy for causality when the underlying dynamics are well understood and confounding factors are absent.
+Therefore, in the context of these frameworks, the use of mutual information as a measure for causal emergence is justified because it captures the relationships between variables in a way that aligns with the assumed known dynamics of the system.
+When comparing the two frameworks, Hoel’s framework can be considered more “causal” than Rosas’s framework.
+This is because Hoel’s framework introduces the do operator in the definition of effective information (EI), as expressed in Equation (7).
+The do operator allows for interventions and represents a causal manipulation of variables, which strengthens the causal interpretation of the framework.
+Furthermore, the significance of EI lies in its deep connections with other measures of causation proposed in various fields Comolatti and Hoel (2022).
+According to the information in Section 3.5, these measures aim to quantify the causal impact of an intervention or treatment on an outcome.
+By incorporating the do operator, EI aligns with these causal effect measures and provides a valuable tool for assessing causal relationships.
+Furthermore, let us delve into the application background of the NIS and NIS+ to understand why these frameworks, based on EI maximization, are considered causal.
+Firstly, the input data in both frameworks, denoted as $x_{t}$, are generated by a Markovian dynamical system, with no other confounding factors considered, apart from the system itself.
+As a result, the state of two consecutive time steps, represented as $x_{t}$ and $x_{t+1}$, without loss of generality, can be modeled by a simple causal diagram $x_{t}\rightarrow x_{t+1}$ due to the Markovian property.
+This causal diagram naturally adheres to the exogeneity assumption in causal inference Pearl (2009), indicating the absence of any confounders between the two states.
+Consequently, the “do” operator in EI can be replaced with conditional probability, which can be estimated from observations.
+Taking all these aspects into consideration, we conclude that the theoretical frameworks of causal emergence and the machine frameworks for causal emergence identification, namely the NIS and NIS+, are indeed related to causality.
+Moreover, they can be extended to other scenarios within the realm of machine learning.
+We further explore and discuss this point in the subsequent sections.
+4.3 Causal Emergence for Machine Learning
+In this sub-section, we introduce how causal emergence can enhance machine learning in out-of-distribution scenarios.
+It turns out that the do intervention introduced in EI captures the causal dependency from the data generation process and thus complements correlation-based machine learning algorithms.
+4.3.1 Out-of-Distribution Generalization
+As demonstrated in the previous subsection, the measurement of causal emergence by vanilla effective information Hoel et al.
+(2013) is defined for stationary dynamics systems in a self-supervised fashion since the input (cause) space and the output (effect) space share the support.
+However, Hoel generalized effective information to general input-output systems Chvykov and Hoel (2020) (see Section 3.5.5), where the cause and the effect space take arbitrary support, which could either be discrete or continuous.
+Empowered by the general notation of EI, causal emergence could be applied in supervised machine learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or its learned representation) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause (feature) to the effect (target).
+It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffices for common prediction tasks with an i.i.d.
+assumption, implying that the training and test data are independently and identically distributed.
+However, if samples are drawn from outside the training distribution, it is essential to learn a representation space that generalizes from the training to the test environment, known as out-of-distribution (OOD) generalization.
+Due to the common belief that causal relations generalize better than statistical correlations Arjovsky et al.
+(2019), causal emergence theory could serve as a criterion for the causation embedded in the representation space.
+The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regard to out-of-distribution generalization.
+Introduction to OOD
+Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic.
+Out-of-distribution generalization toward arbitrary distributional shifts is generally considered impossible Ye et al.
 (2021).
-Causal feature learning aims that to find a high-level causal relations from the micro-level data.
-In causal representation learning the macro states are causal variables or causal representations Schölkopf et al.
-(2021), macro-dynamics are causal mechanisms, and the learned coarse-graining strategy is the representation or encoding process.
-In this sub-section, we will introduce causal representation learning in details at first, and compare the two emerging fields.
-4.2.1 Introduction of Causal Representation Learning
-Causal representation learning is an emerging field in artificial intelligence, it tries to combine two important fields in machine learning: representation learning and causal inference, together.
+Therefore, it is commonly assumed that there exists an underlying causal mechanism governing the data generation process, which remains invariant to distributional shifts.
+Formally, according to Pearl’s Structural Equation Model Pearl (2009), which causally formulates data generation, the target $Y$ is supposed to be generated from a part of the feature $X$, i.e., $Y=f(X_{c},\epsilon)$, where $X_{c}$ represents the selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$.
+Obviously, $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments.
+On the other hand, the raw feature $X$ might comprise a mixture of causal variables and non-causal variables.
+The latter could arise from confounding, selection bias, or anti-causal mechanisms, all of which introduce spurious correlations between the target and non-causal variables.
+The strength and direction of these spurious correlations vary across environments, resulting in distributional shifts that hamper the out-of-distribution performance of machine learning techniques.
+As a remedy, a prominent branch of OOD algorithms Arjovsky et al.
+(2019); Huh and Baidya (2022); Zhao et al.
+(2019); Ganin et al.
+(2016) learns an invariant representation space distilling the causal variables from the raw feature space.
+Typically, a metric of the representation is devised to verify the stability of the correlation between the learned representation space and the target and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features.
+The paradigm for invariant representation learning can be expressed as Wang et al.
+(2022):
+$\min_{f,\phi}\mathcal{R}(f(\phi(X)),Y)+\lambda\ell_{reg}(\phi),$ (52)
+where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function, and $\ell_{reg}$ is the regularization.
+Representation learning-based OOD algorithms are diversified in their constraints on the representation space characterized by the regularizers.
+Each metric of the representation fits a specific data generation pattern and does not guarantee to filter out the causal variables when the setting is violated.
+However, causal emergence potentially offers a united metric of representation for OOD generalization rooted in causality theory.
+Before diving into the application of causal emergence, we first discuss the representative regularizers of representation learning and their limitations.
+Domain-invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments.
+The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al.
+(2018) or an adversarial classifier for the environment index Ganin et al.
+(2016).
+However, the alignment constraint provides a generalization guarantee only for a covariate shift Zhao et al.
+(2019), where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$.
+On the other hand, invariant risk minimization Arjovsky et al.
+(2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$.
+Invariant risk minimization has been proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target but is reported to fail under a covariate shift Ahuja et al.
+(2021).
+Further, conditional invariant representation Li et al.
+(2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect.
+The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
+It has been theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models, and it does not generalize across all data generation processes Kaur et al.
+(2022).
+As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, hindering the general applicability of invariant representation learning.
+Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedures, increasing the data collection and labeling expenses.
+To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
+Causal Emergence and OOD
+The link between OOD generalization and causal emergence can be attributed to the do intervention of effective information.
+By definition Pearl (2009), intervening in the feature distribution with $do(X)$ modifies the structural equations of all the variables in the feature space to a constant value, whereas the equation for the target remains intact as $Y=f(X_{c})$.
+On one hand, the assumption of the causal mechanism’s invariance to distributional shift is respected since the causal correlation between causal variables and the target is reserved.
+On the other hand, the spurious correlation between non-causal variables and the target is removed by the intervention.
+For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation with $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$.
+As a result, $X_{n}$ is no longer caused by $Y$, and the spurious correlation vanishes.
+In another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$.
+Then, the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist.
+In summary, the intervention introduced by EI captures the causal correlation between features and the target while suppressing the spurious correlation.
+It renders EI an ideal metric to assess the causality contained in the representation space and, thus, the generalization of the induced algorithm.
+Recall that an increment in effective information implies the occurrence of causal emergence.
+Therefore, we propose a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning.
+Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e., $I(g(X);Y)\leq I(X;Y)$, as per the data processing theorem.
+Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space.
+In contrast, EI (or normalized EI) would possibly peak at a medium stage of abstraction of raw features, coinciding with the philosophy of OOD generalization that less could be more.
+Ideally, at the peak of EI (or normalized EI) where causal emergence occurs, all non-causal features are excluded, and the causal features are revealed, resulting in the most informative representation while remaining invariant to distributional shifts.
+The removal of spurious correlations and recovery of causal mechanisms by manipulating feature distributions are principles shared by the reweighting technique widely adopted in the OOD literature.
+For instance, stable learning Shen et al.
+(2020a, b); Zhang et al.
+(2021) is designed for scenarios where collinearity exists in input variables, causing spurious correlations between non-causal variables and the target.
+It achieves this by reweighting training samples to decorrelate the features, thus reducing collinearity and eliminating spurious correlations.
+Similar to EI, stable learning is free of environment index labeling.
+Further, both sample reweighting and feature decorrelation share the philosophy of distribution intervention.
+In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
+5 Discussion and Perspectives
+We have presented several quantitative theoretical frameworks on causal emergence.
+However, there are numerous unexplored problems and implications that warrant further discussion.
+In this section, we address four key topics for future research: causal emergence and causal representation learning, ontological and epidemiological causality and emergence, potential applications in complex systems, and understanding complex systems from the perspective of causal emergence.
+5.1 Causal Emergence and Causal Representation Learning
+Causal representation learning (CRL) is an emerging field in artificial intelligence (AI) Schölkopf et al.
+(2021) that combines two important fields in AI: representation learning and causal inference.
 Representation learning aims to extract important features (or representations) hidden in data to make predictions automatically Bengio et al.
 (2013).
-Deep neural networks are always adopted, important features in data are represented by the activation modes of neurons LeCun et al.
-(2015); Goodfellow et al.
-(2016).
-The modes in different layers usually represent significant features in various scales, and deeper layers always correspond to larger scales Alzubaidi et al.
-(2021).
-The features in different layers are combined together to make predictions or classifications.
-For example, a deep neural network for human face recognition may have two components.
-The first one is a deep neural network with stacked convolution units LeCun et al.
-(1998), and it tries to extract features of faces on different scales.
-The second component is always a common feed-forward neural network which can accept the input features of faces and output a label to identify the character represented by the input image .
-The first component is for representation, and the second one is for classification.
-The neural network as a whole must learn to extract important features from images to make high precision classification.
-Although representation learning has achieved remarkable success in various domains such as image classification LeCun et al.
+It can be regarded as a typical application of deep learning and has achieved remarkable success in various domains such as image classification LeCun et al.
 (2015), face recognition Wang et al.
@@
 (2015); Dabney et al.
-(2020), it suffers from a critical limitation: it can only capture the information or associations within the data, but not the underlying causal relationships.
-According to the theory of causal hierarchy by Judea Pearl Pearl et al.
-(2000); Pearl (2009); Cui and Athey (2022), causality is very different from association because the former always represents a more stable relationship which is invariant in different environments, while the latter may be more dependent on contexts and reflect the limitation of the data.
-For instance, in the example of face recognition mentioned previously, if all the images in the training data have outdoor backgrounds, e.g.
-grasslands, lakes, trees, etc., the neural network may not work well on the images with indoor backgrounds Cui and Athey (2022); Zhang et al.
 (2020).
-This is because the neural network may learn the fake association between backgrounds but not foregrounds and labels.
-However, the association is only the reflection of the bias in the training data, but not the real causal relationship.
-In causal inference, people use graphs to represent causal relationships.
-Nodes represent causal variables, and edges which are always combined with other information like structural equations or conditional probabilities to represent causal relations Pearl et al.
+However, conventional representation learning suffers from a critical limitation: it can only capture the information of associations within the data but not the underlying causal relationships.
+Therefore, it is important to consider causality in representation learning.
+To address the problem, CRL tries to combine the advantages of the two sub-fields: representation learning and causal inference to extract the important features and the relationships with causation behind data automatically Schölkopf et al.
 (2021).
-In conventional causal inference, while the parameters in a causal model can be adjusted based on data, all the variables and causal relationships are typically extracted manually rather than directly from data itself.
-To conquer the problem, causal representation learning tries to combine the advantages of the two sub-fields to extract the important features and the relationships with causality behind data automatically Schölkopf et al.
+A typical CRL scenario is depicted in Figure 16.
+Suppose there is a set of variables and causal mechanisms (including causal graphs and structural equations) that describes how the world of robotic arms and colorful blocks works.
+However, these mechanisms are not directly observable for the agent undergoing causal representation learning.
+Instead, what the agent can observe is a set of images generated by robotic arms and blocks.
+The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks, for example, causal relationship discovery Monti et al.
+(2020) and predictions in different environments Peters et al.
+(2016) or domains Cui and Athey (2022); Kuang et al.
 (2018).
-Figure 16: Illustration for the workflow of a typical causal representation learning agent.
-In the top panel, we have a depiction of a real block world with a robotic arm, along with the corresponding causal graph representing the relationships between the variables in the real world, nevertheless, it is totally unknown for the agent.
-Additionally, we have the data in the form of images generated by the block world.
-In the bottom panel, we see the structure of the causal representation learning agent.
+Figure 16: Illustration of the workflow of a typical causal representation learning (CRL) agent.
+The top panel depicts a real block world with a robotic arm and the corresponding causal graph, representing relationships between variables in the real world, which is totally unknown to the agent.
+Additionally, the data are presented as images generated by the block world.
+The bottom panel depicts the structure of the causal representation learning agent.
 It utilizes a convolutional neural network (CNN) to extract meaningful representations from the data and perform downward tasks based on these representations.
 The encoder plays a crucial role in capturing relevant features and patterns from the input data, enabling the agent to learn and understand causal relationships.
-A typical scenario of causal representation learning can be depicted by Figure 16.
-Suppose there are a set of variables and causal mechanisms (including causal graphs and structural equations) between them which can describe how the world of robotic arms and colorful blocks works, whereas, they are not observable for the agent of causal representation learning.
-What the agent can observe is a set of images which are generated by robotic arms and blocks.
-The objective of a causal learning framework is to extract the causal variables and mechanisms from the observed images, and the variables and mechanisms can be used to implement other downstream tasks.
-Formally, if the input data can be described by $X=(\boldsymbol{x}_{1},\boldsymbol{x}_{2},...,\boldsymbol{x}_{p})$, causal representation learning will construct two functions and a directed acyclic graph, which include a causal representation function: $\tilde{\phi}:\mathcal{R}^{q}\rightarrow\mathcal{R}^{p}$, such that:
-$X=\tilde{\phi}(Y),$ (49)
-where, $Y=(\boldsymbol{y}_{1},\boldsymbol{y}_{2},...,\boldsymbol{y}_{q})$ is a set of causal variables and $q\ll p$; a directed acyclic causal graph $\mathcal{G}_{S}\equiv(Y,E)$, on which the node set is $Y$, and the directed links in $E$ represent causal relations; as well as a set of causal mechanisms:
-$\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i}),$ (50)
-for $i=1,2,…,q$, where $f_{i}$ is the structural equation for the $i$th causal variable $\boldsymbol{y}_{i}$, which describes how the parent causal variables $PA_{i}\equiv\{j|j\rightarrow i\in E\}\subset\{\boldsymbol{y}_{i}\}$ and the external unknown variable $U_{i}$ determine the value of $\boldsymbol{y}_{i}$.
-And there is a requirement for $U_{i}$ for all $i,j$, that is:
-$P(U_{i},U_{j})=P(U_{i})P(U_{j}).$ (51)
-Thus, any two external unknown variables need to be independent each other.
-However, learning the three elements including the two functions and a graph in the same time is very hard, we can convert the problem to reduce the difficulty.
-This is to learn disentanglement representations, in which the representations are independent(disentangled) each other Wang et al.
-(2022).
-We will show how we can derive disentanglement representations from Equations 49 to 51.
-Notice that if we combine the equations $\boldsymbol{y}_{i}=f_{i}(PA_{i},U_{i})$ for all $i\in[1,q]$ together, we have $Y=F(Y,U)$, where $U=(U_{1},U_{2},...,U_{n})$, and $F=(f_{1},f_{2},...,f_{q})$.
-That means, the causal mechanisms actually describe how the unknown variables $U$ determine the causal variables $Y$.
-Suppose we can solve this equation to obtain $Y$ expressed by $U$, and we define $\tilde{F}$ such that: $Y=\tilde{F}(U)$, according to Equation 49,
-$X=(\tilde{\phi}\circ\tilde{F})(U)$ (52)
-if we define $\phi\equiv\tilde{\phi}\circ\tilde{F}$, then $\phi$ can be understood as a decoder which can decode a disentanglement representation to data.
-Further, if we could construct a encoding function $\phi^{{\dagger}}:\mathcal{R}^{p}\rightarrow\mathcal{R}^{q}$ which takes $X$ to $U$, and we require that:
-$\phi\circ\phi^{{\dagger}}\sim I,$ (53)
-where, $I$ is the identity mapping, thus, $\phi\circ\phi^{{\dagger}}$ is an auto-encoder, which can encode the input data $X$ into a latent space $U$ and reconstruct $X$ according to $U$.
-While, according to Equation 51, the latent variables are independent each other.
-That is, we represent the input data by a set of independent or disentanglement latent variables, and we can reconstruct the input data according to these independent variables.
-Therefore, this is called disentanglement representation learning.
-In practice, independent component decomposition method can be use to construct the decoder Hyvarinen and Morioka (2016); Khemakhem et al.
-(2020).
-To reduce the computational complexity, we can also use invertible neural network to model $\phi$ Dinh et al.
-(2014).
-With the disentangled representation variables $U_{i}$, we can effectively perform downstream tasks.
-One prominent application of disentangled representation learning is causal discovery.
-By leveraging the disentangled representations, we can uncover causal relationships between variables and gain insights into the underlying causal structure of a system Monti et al.
-(2020).
-Causal discovery allows us to identify cause-and-effect relationships and understand how different factors interact and influence each other.
-This knowledge can be invaluable in various domains, such as understanding complex systems, making informed decisions, and designing interventions or policies.
-To conclude, a typical causal machine learning or disentanglement representation learning framework is always composed of a module for causal or disentanglement representations and a module which contains sub-units for different downstream tasks.
-The representation module, which contains an encoder and a decoder, tries to extract abstract representations of the data, and the representations are not for prediction but for causal inference.
-The downstream tasks may be variant.
-For example, the transferable mechanism learning tries to look for independent causal mechanisms to complete prediction tasks in very different environments Peters et al.
-(2016).
-And existing studies have shown that learning causal models may help in transferring modules across different domains Cui and Athey (2022); Kuang et al.
-(2018).
-Causal representation learning has been applied to various problems.
-4.2.2 Comparisons between Causal Emergence Identification and Causal Representation Learning
-There are similarities and differences between causal emergence and causal representation learning (CRL). We will introduce them in detail separately.
-Similarities
-It is interesting to compare the tasks of causal emergence identification under the theoretic framework of Erik Hoel and causal representation learning.
-Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro variables are the causal variables, the macro dynamics are the causal mechanisms, the coarse-graining strategy is an encoding process from the original data to the representations.
-Under this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism.
-To identify causal emergence, we need to learn the appropriate causal variables to represent data and to discover causal mechanism on the macro-level (on representational latent space) such that the EI of the learned causal mechanism in latent space is larger than the original data.
-We use Table 1 to clarify the comparisons item by item.
-Table 1: Comparison between causal emergence identification and causal representation learning.
+When comparing Figures 14 and 15 with Figure 16, we found that as a framework for machine learning, Neural Information Squeezer Plus can be regarded as a kind of CRL framework.
+Thus, with the NIS and NIS+, we effectively employ CRL to address the problem of identifying causal emergence (CE).
+Both CE and CRL frameworks feature an encoder and decoder designed to represent raw data as causal variables.
+Latent causal mechanisms can be learned in both frameworks.
+In this section, we compare the similarities and differences between these two emerging fields.
+5.1.1 Similarities
+It is interesting to compare the tasks of causal emergence identification within the theoretical framework of Erik Hoel and causal representation learning.
+Actually, the task of causal emergence identification can be understood as a task of causal representation learning, where the macro-variables are the causal variables, the macro-dynamics are the causal mechanisms, and the coarse-graining strategy is an encoding process that transforms the original data into representations.
+Within this framework, EI can be understood as a measure of the strength of the causal effect on the mechanism.
+To identify causal emergence, we need to learn the appropriate causal variables to represent data and discover causal mechanisms at the macro-level (in the representational latent space) to ensure that the EI of the learned causal mechanism in the latent space is larger than that of the original data.
+Table 1 shows a detailed comparison of causal emergence identification and causal representation learning.
+Table 1: Comparison of causal emergence identification and causal representation learning.
 Causal Emergence Identification Causal Representation Learning
-Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real
+Data Observations (time series) of micro-states Raw data generated by some causal mechanism in real life
 Latent variables Macro-states Causal representations
@@
 Optimization for causality EI maximization Prediction loss, disentanglement
-Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has stronger causal effect Finding an optimal representation of the raw data such that the independent causal mechanism can be realized by the representations.
+Objective Finding an optimal coarse-graining strategy and a macro-dynamic that has a stronger causal effect Finding an optimal representation of the raw data to ensure that the independent causal mechanism can be realized by the representations
 With these similarities, the technologies and concepts from both fields can learn from each other.
@@
 On the other hand, the learned abstract causal representations can be viewed as macro-states, which enhances the explainability of causal representation learning.
-This perspective allows us to consider that CRL is actually uncovering the hidden causal emergent features within the data.
+This perspective suggests that CRL essentially uncovers hidden causal emergent features within the data.
 Moreover, these similarities between emergent phenomena and CRL contribute to a deeper understanding of emergence itself.
-By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract a more profound causal structures.
+By applying CRL frameworks to data generated by dynamical systems exhibiting emergent phenomena, we can extract more profound causal structures.
 These profound causal structures may serve as the origins of the mysteries surrounding emergence.
-By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we gain insights into the mechanisms behind emergent phenomena.
-These causal structures provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components.
-Exploring these intricate causal structures brings us closer to unraveling the enigma of emergence and sheds light on the fundamental principles governing complex systems.
-Consequently, the identification of emergent phenomena requires an observer, whether it be a human or a machine, to actively search for them.
-By employing a machine observer, it is possible to potentially resolve the longstanding debates regarding whether emergent causality is epistemic or ontological.
-This is because the machine observer can provide a relatively objective outcome, although it is important to note that the result is still influenced by the observer.
-Furthermore, an intriguing area for future exploration involves investigating how the structure and capabilities of the observer can impact the emergence results that are identified by modifying the parameters of machine learning.
-This highlights the potential influence of the observer’s characteristics on the outcomes obtained through the rewiring of machine learning parameters.
-Differences
-However, there are several differences between the CRL and causal emergence in theory.
-The biggest gap is that CRL assumes that there is a real causal mechanism behind it, and the data is generated by this causal mechanism Schölkopf et al.
+By delving into the depths of CRL and uncovering the hidden causal relationships within complex systems, we can gain insights into the mechanisms behind emergent phenomena.
+These causal structures may provide a foundation for understanding the emergence of novel properties and behaviors that arise from the interactions of simpler components.
+5.1.2 Differences
+However, there are several theoretical differences between CRL and causal emergence.
+The biggest difference is that CRL assumes that there is a real causal mechanism behind the data, and the data are generated by this causal mechanism Schölkopf et al.
 (2021).
-While for causal emergence identification, the emergent variables and mechanisms on macro-level are just handy ways to observe and understand data, "real causality" may not exist.
-The difference can be illustrated by Figure 17
-Figure 17: The Comparison between Causal Emergence Identification and Causal Representation Learning.
+In contrast, for causal emergence identification, the emergent variables and mechanisms at the macro-level are just handy ways to observe and understand data, and “real causality” may not exist.
+The differences are illustrated in Figure 17.
+Figure 17: Comparison of causal emergence identification and causal representation learning.
 All the potential co-founders are ignored.
-The black arrows represent causal relations, while the blue arrows represent interactions between macro-variables.
-The dashed directed lines represent the information flows for coarse-graining, representation or data generation.
-While, if epistemological perspective is adopted, this difference disappears because all they do is to extract meaningful information from observational data to obtain representations with stronger causal effects.
+The black arrows represent causal relations, whereas the blue arrows represent interactions between macro-variables.
+The dashed lines with arrows represent the information flows for coarse-graining, representation, or data generation.
+However, if an epistemological perspective is adopted, this difference disappears because both approaches extract meaningful information from observational data to obtain representations with stronger causal effects.
 Another major difference is that the causal mechanism for macro-states in causal emergence is a dynamical system.
-Therefore, if the dynamics in macro-state space is on a network, circular structures may exist because the state variables will be iterated in the dynamical system.
-As a result, feed-backs and circular indirect interactions are allowed in such model.
+Therefore, if the dynamics in the macro-state space are on a network, circular structures may exist because the state variables will be iterated in the dynamical system.
+As a result, feedback and circular indirect interactions are allowed in such a model.
 However, the causal mechanism is always represented by a structural causal model in CRL.
-While, these models are always directed acyclic graphs, therefore, loops or circular structures are not allowed.
-The problem is that the meaning of the links in a network dynamics model is different from the links in structural causal model.
-Hence, the conflict can be resolved if we comprehend that the links in a structural causal model depict an average (net) causal effect between variables when they are in a steady state.
-On the other hand, the links in network dynamical models within causal emergence represent instantaneous interactions.
-Finally, unknown co-founders are always ignored in causal emergence, while, they play very important roles in causal structural models.
-4.2.3 Out-of-distribution Generalization
-Causal representation learning is established in a self-supervised manner, in the sense that it reveals the intrinsic causal structure of data.
-Thus, the learned causal representations could not be exploited directly for supervised prediction task, where the causation between data and a specified target must be characterized.
-The measurement of causal emergence by vanilla Effective Information Hoel et al.
-(2013) is also defined for stationary dynamics systems in a self-supervised fashion, since the input(cause) space and the output(effect) space share the support.
-However, Hoel generalizes Effective Information to general input-output systems Chvykov and Hoel (2020)(see section 3.6.1 where the cause and the effect space take arbitrary support, which could either be discrete or continuous.
-Empowered by the general notation of EI, causal emergence could be applied in supervised learning to evaluate the strength of causation between the feature space $\mathcal{X}$ (or a representation learned upon which) and the target space $\mathcal{Y}$, thus enhancing the prediction from the cause(feature) to the effect(target).
-It is worth noting that a direct fitting from $\mathcal{X}$ to $\mathcal{Y}$ on the observations suffice for common prediction tasks with an i.i.d.
-assumption, implying that the training and test data are independently and identically distributed.
-However, if samples are drawn out of the training distribution, it is essential to learn a representation space that generalizes from training to test environment, known as out-of-distribution (OOD) generalization.
-Due to the common belief that causal relations generalize better than statistical correlation Arjovsky et al.
-(2019), causal emergence theory could serve as a criteria for the causation embedded in the representation space.
-The occurrence of causal emergence indicates the revelation of potential causal factors of the target, thus producing a robust representation space with regarding to out-of-distribution generalization.
-Introduction of OOD
-Before the detailed discussion of the connection between causal emergence and OOD generalization, we begin with a brief introduction to the latter topic.
-Out-of-distribution generalization towards arbitrary distributional shifts is generally impossible Ye et al.
+Although these models are always directed acyclic graphs, loops or circular structures are not allowed.
+However, these differences are not important because Markovian dynamics can always be converted to causal models, as mentioned in Section 2.1.
+Finally, unknown co-founders are always ignored in causal emergence, whereas they play very important roles in causal structural models.
+In summary, there are deep connections between causal emergence and CRL.
+On one hand, the machine learning framework for causal emergence identification exhibits similar structures to the frameworks for CRL.
+And even the concepts can establish one-to-one correspondence, as shown in Table 1.
+Therefore, the ideas and techniques of both sides can learn from each other.
+On the other hand, studies of causal emergence can provide new perspectives and insights for causal representation learning algorithms.
+For example, we can interpret the learned representations and causal mechanisms as emergent causal laws.
+The maximization of EI may improve the efficiency of learning the causalities for both CRL and reinforcement learning agents.
+5.2 Ontological and Epistemological Causality and Emergence
+Although machine learning techniques have facilitated the learning of causal structures and models, as well as the exploration of emergent properties and causation, it is important to consider whether the results obtained through machine learning reflect ontological causality and emergence or if they are merely epistemological phenomena.
+Here, ontological causality refers to the causal relationships that exist in objective reality, independent of our knowledge or understanding.
+Ontological causality explores the fundamental mechanisms and interactions that give rise to causal effects in the world.
+Similarly, ontological emergence is concerned with the objective existence of emergent properties and their underlying mechanisms.
+Epistemological causality, on the other hand, focuses on our knowledge and understanding of causality.
+It deals with how we perceive, model, and explain causal relationships based on our observations and experiences, but the causality may not exist in the real world.
+Epistemological emergence, akin to epistemological causality, focuses on our comprehension and elucidation of emergent phenomena.
+In other words, epistemological causality and emergence depend on an observer, and different observers may have distinct perspectives about causality and emergence for a particular objective phenomenon.
+There has been a longstanding debate regarding the ontological and epistemological aspects of causality and emergence throughout history Kim (1992); Crutchfield and Young (1989); Bonabeau and Dessalles (2011); Yurchenko (2023).
+The authors of Yurchenko (2023) highlighted that the concept of “causation” in the literature is often vague and should be differentiated into “cause” and “reason”, aligning with ontological and epistemological causality.
+“Cause” means genuine cause that sufficiently causes the effects (causal closure principle Clayton and Davies (2006) and exclusion principle Kim (2007)), whereas “reason” serves as a mere explanation for individuals to comprehend the effects.
+Reason may not possess the same level of rigor as genuine cause, but it does offer a certain degree of predictability that individuals find valuable in certain circumstances.
+However, distinguishing between them, particularly when addressing specific phenomena, remains a challenging task.
+One such controversial concept is downward causation, which has sparked extensive discussions.
+The question of whether downward causation exists objectively or not remains open.
+In Rosas et al.
+(2020), Rosas argued that downward causation not only exists independently of particular observers but also provides a characterization of the phenomenon through quantified measures using information decomposition.
+However, Yurchenko proposed that it is important to differentiate between two separate concepts: causality and reasoning.
+According to this perspective, downward causation falls under the category of “reason” rather than causation.
+In this context, causality and reasoning can be seen as representing ontological and epistemological causality, respectively.
+Similarly, debates persist regarding the nature of causal emergence.
+The question arises as to whether causal emergence is a genuine phenomenon that exists independently of particular observers Dewhurst (2021).
+In Section 3.5.4, we highlighted that different coarse-graining strategies can lead to distinct macro-dynamic mechanisms with varying measures of EI.
+In essence, different coarse-graining strategies can be seen as representing different observers.
+However, Hoel’s theory proposes a criterion to differentiate between coarse-graining methods, namely EI maximization.
+Consequently, for a given set of Markovian dynamics, only the coarse-graining strategy and corresponding macro-dynamics that maximize the EI measure can be considered as an objective outcome.
+Nevertheless, a challenge arises when multiple solutions exist that maximize EI, introducing a degree of subjectivity Dewhurst (2021).
+The authors of Zhang and Liu (2022); Yang et al.
+(2023) addressed this issue by introducing constraints on prediction errors in micro-states, which partially alleviates the problem.
+However, the question of whether optimal results have unique solutions remains an open problem that requires further investigation.
+Similarly, although the incorporation of machine learning cannot resolve debates surrounding ontological and epistemological causality and emergence, it can provide objective standards that help mitigate subjectivity.
+This is because machine learning algorithms strive to optimize objective functions.
+Thus, machine learning agents can be viewed as “objective” observers that make judgments on causality and emergence.
+This represents an additional advantage of incorporating machine learning.
+However, the issue of unique solutions is crucial in this approach and warrants further attention.
+Are the machine learning results ontological or epistemological?
+The answer is that the results are epistemological and depend on machine learning algorithms.
+However, this does not mean that all the results of machine learning are meaningless since if the learning agents are well trained and the defined mathematical objectives are effectively optimized, the results can also be considered objective and independent of the algorithms.
+It is intriguing to pose the same question regarding human cognition.
+Do all cognitive results in the human brain, encompassing causality, coarse-graining strategies, and macro-dynamics, reflect objective reality, or are they purely epistemological?
+Firstly, answering this question is exceedingly challenging.
+Secondly, the cognitive results specific to a particular human brain must objectively manifest within the neural network structures.
+Thus, for that brain, these results are ontological.
+Similar reasons hold for machine learning.
+The cognitive results of machine learning algorithms are recorded by the structures and parameters in neural networks.
+Therefore, studying these structures may reveal the properties of the original systems, and the nature of the interactions between the observer and the observed objects can be objectively reflected by machines.
+Furthermore, the integration of machine learning can assist in establishing a theoretical framework for modeling observers and studying the interactions between the observer and the corresponding observed complex systems.
+This framework allows us not only to explore the hard problems regarding causality and emergence but also to understand the limitations and boundaries of the observer.
+One example of this research is the emergent classicality of quantum systems from the information bottleneck formed by machine learning algorithms Zhang and You (2023).
+5.3 Potential Applications in Complex Systems
+In Section 4, we discussed the techniques related to causal emergence identification, machine learning, and causal inference.
+In this sub-section, we discuss the potential applications of these techniques in complex systems.
+Complex systems can be understood as a large network, with events connected by causal links.
+Automatically discovering the complex causal relationships from data is a challenging problem.
+The research area of causal discovery has tried to address this problem in various ways Glymour et al.
+(2019); Zanga et al.
+(2022); Donges et al.
+(2009); Kalisch and Bühlman (2007); Zheng et al.
+(2018); Zhu et al.
 (2019).
-Therefore, it is commonly assumed there exists an underlying causal mechanism that governs the data generation process, which remains invariant to distributional shifts.
-Formally, by Pearl’s Structural Equation Model Pearl (2009) which causally formulates data generation, the target $Y$ is supposed to be generated from part of the feature $X$, i.e.
-$Y=f(X_{c},\epsilon)$ where $X_{c}$ is selected features or a representation of $X$, and $\epsilon$ is a random noise independent of $(X,Y)$.
-Obviously $X_{c}$ is the true cause of the target, and the causal mechanism $f$ is presumed to be invariant across training and test environments.
-On the other hand, the raw feature $X$ might be a mixture of causal variables and non-causal variables.
-The latter could be produced by confounding, selection bias or anti-causal mechanisms, all of which introduce spurious correlation between the target and non-causal variables.
-The strength and direction of spurious correlation are variant across environments, resulting in distributional shifts that hampers out-of-distribution performance of machine learning techniques.
-As a remedy, a prominent branch of OOD algorithms Arjovsky et al.
-(2019); Huh and Baidya (2022); Zhao et al.
-(2019); Ganin et al.
-(2016) learn an invariant representation space distilling the causal variables from the raw feature space.
-Typically a metric of the representation is devised to verify the stability of correlation between the learned representation space and the target, and is adopted as a regularization for the optimization objective to encourage the representation learner to extract causal variables from features.
-The paradigm for invariant representation learning could be formulated as Wang et al.
-(2022):
-$\min_{f,\phi}\mathcal{R}(f(\phi(X)),Y)+\lambda\ell_{reg}(\phi),$ (54)
-where $\phi(X)$ is the learned representation, $f$ is the learned predictor, $\mathcal{R}$ is the task-specific risk function and $\ell_{reg}$ is the regularization.
-Representation learning based OOD algorithms are diversified in their constraint of the representation space characterized by the regularizers.
-Each metric of the representation fits a specific data generation pattern, and does not guarantee to filter out the causal variables when the setting is violated.
-However, causal emergence potentially offers a united metric of representation for OOD generalization rooted from causality theory.
-Before diving into the application of causal emergence, we would first discuss representative regularizers of representation learning and their limitations.
-Domain invariant representation learning proposes an aligned feature space with the constraint $g(X)\perp\mathcal{E}$, where $\mathcal{E}$ denotes the environment index for training and test environments.
-The regularization is implemented by the metric of Maximum Mean Discrepancy (MMD) distance of representation across environments Li et al.
-(2018) or an adversarial classifier for the environment index Ganin et al.
+However, additional challenges may arise when applying these causal discovery methods in complex systems because circular causal structures and cross-level causation may exist.
+Circular causal structures widely exist in complex dynamical systems due to the feedback effect.
+That is, one variable may affect itself via a direct self-loop feedback or a long chain forming a circle.
+This circular causal structure may pose a challenge to existing methods of causality because most of these methods study directed acyclic causal graphs.
+However, recent progress has seen the development of some methods for discovering these circular structures in a data-driven manner, e.g., Richardson (2013); Lacerda et al.
+(2012); Forré and Mooij (2018); Rantanen et al.
+(2020); Zhang et al.
+(2019); Pamfil et al.
+(2020); Vowels et al.
+(2022); Spirtes (2013).
+Another aspect is that higher-level or cross-level causation may exist if the ingredients of scale and coarse-graining are considered.
+For example, downward causation describes the causal effect between macro-level and micro-level variables.
+Thus, the causal connections may be cross-level.
+If causal emergence occurs in a complex system, some strong causality may be found between macro-variables.
+All of these factors in complex systems present new challenges.
+The machine learning methods discussed in Section 4.1.1 can address the question of whether higher-level causality exists at the macro-level.
+However, it is important to note that this is a global property.
+It must be further developed and extended to find a local causal relationship between macro-variables or macro-variables and micro-variables Varley (2022).
+Moreover, existing methods of causal discovery must be extended to consider the method of grouping a set of variables or coarse-graining a system.
+For example, Cheng et al.
+(2023) proposed a coarse-to-fine causal discovery algorithm based on Granger causality and a graph neural network.
+The grouping process of variables is from coarse to fine, which can improve the efficiency of the algorithm.
+Other multi-level causal discovery methods were proposed and discussed in Wang et al.
+(2023); Zhang et al.
+(2023); Fan et al.
 (2023).
-However, the alignment constraint provides generalization guarantee only for covariate shift Zhao et al.
-(2019) where the distribution shift is restricted to the features $X$ without perturbing $P(Y|X)$.
-On the other hand, invariant risk minimization Arjovsky et al.
-(2019); Koyama and Yamaguchi (2020) pursues an invariant conditional distribution $P(Y|g(X))$ of the representation across environments, which translates to $Y\perp\mathcal{E}\;|\;g(X)$.
-Invariant risk minimization is proven to recover the causal variable in a data generation process where the non-causal variable is anti-causally generated by the target, but is reported to fail under covariate shift Ahuja et al.
-(2021).
-Further, conditional invariant representation Li et al.
-(2018); Huh and Baidya (2022) is designed specifically for image classification where the target is modeled as the cause of the feature instead of the effect.
-The conditional independence $g(X)\perp\mathcal{E}\;|\;Y$ is enforced on the representation space.
-Actually, it is theoretically proven that any statistical independence regularization is valid for a subgroup of Structural Equation Models and it does not generalize across all data generation processes Kaur et al.
-(2022).
-As a result, the metric of representation manages to capture the causation between the representation space and the target variable only if a specific causal graph of data generation is given, which hinders the general applicability of invariant representation learning.
-Worse still, the existence of the environment index $\mathcal{E}$ in the independence regularizers introduces the necessity of multiple environments for training procedure, increasing the data collection and labelling expenses.
-To this end, there is active demand for a unified and general metric of representation for OOD generalization, for which causal emergence might be a good candidate.
-Causal emergence and OOD
-The match between OOD generalization and causal emergence is attributed to the do intervention of Effective Information.
-By definition Pearl (2009), intervening the feature distribution with $do(X)$ modifies the structural equations of all the variables in feature space to a constant value while the equation for the target remains intact as $Y=f(X_{c})$.
-On the one hand, the assumption of causal mechanism’s invariance to distributional shift is respected, since the causal correlation between causal variables and the target is reserved.
-On the other hand, the spurious correlation between non-causal variables and the target are removed by the intervention.
-For example, a non-causal variable anti-causally correlated with the target leads to a structural equation as $X_{n}=h(Y,\epsilon)$, and $do(X_{n})$ replaces the equation by $X_{n}=U$, where $U\sim U_{max}$, and $U_{max}$ represents the maximum entropy distribution on the domain of $X_{n}$.
-As a result, $X_{n}$ is no longer caused by $Y$ and the spurious correlation vanishes.
-For another example, if a non-causal variable is correlated with the target by a confounder $E$ (probably in the feature space), the structural equation takes the form of $X_{n}=\phi(E,\epsilon)$ and $Y=f(X_{c},E,\epsilon)$.
-Then the do intervention would modify the equation to $X_{n}=U$ and confounding ceases to exist.
-In summary, the intervention introduced by EI captures the causal correlation between features and the target whereas suppresses the spurious correlation.
-It renders EI an ideal metric to assess the causality contained by the representation space and thus the generalization of the induced algorithm.
-Recall that an increment in the Effective Information implies the occurrence of causal emergence.
-Therefore, we bring forward a conjecture that out-of-distribution generalization could be achieved simultaneously with maximizing EI (or normalized EI) for representation learning.
-Notably, Shannon’s mutual information holds the property that any representation has no gain of mutual information with the target over the raw feature space, i.e.
-$I(g(X);Y)\leq I(X;Y)$, by data processing theorem.
-Thus, direct fitting inclines to absorb as much information with the target as possible into the representation space.
-In contrast, EI(or normalized EI) would possibly peak at a medium stage of abstraction of raw feature, coinciding with the philosophy of OOD generalization that less could be more.
-Ideally, at the peak of EI(or normalized EI) where causal emergence occurs, all non-causal features are excluded and the causal features are revealed, resulting in the most informative representation while performing invariant to distributional shifts.
-Actually, removing spurious correlation and recovering causal mechanisms by manipulating feature distributions is shared by the reweighting technique widely adopted in OOD literature.
-For instance, Stable Learning Shen et al.
-(2020a, b); Zhang et al.
-(2021) is oriented at the setting where collinearity exists in input variables, causing spurious correlation between non-causal variables and target.
-It reweights training samples to decorrelate the features and thus reducing collinearity and eliminating spurious correlation.
-Similar to EI, Stable Learning is free of environment index labeling.
-Further, sample reweighting and feature decorrelation shares the philosophy of distribution intervention.
-In this sense, EI could be viewed as an information-theoretic abstraction of reweighting-based debiasing techniques for OOD generalization.
-4.3 Other Related Topics in AI
-There are other related topics in artificial intelligence, e.g., abstraction of causal models and world model based reinforcement learning. We will introduce them in detail.
-4.3.1 Abstraction of Causal Models
-In CRL, we have the ability to extract causal relationships and construct causal models from raw data.
-However, in cases where the causal model becomes excessively large and complex, it becomes necessary to employ an abstraction process Rubenstein et al.
-(2017).
-Abstraction in causal modeling refers to the reduction or simplification of the causal model while preserving its essential causal relationships and important features.
-Coarse-graining is the multi-scale description of the state information on different information reduction scales.
-While causal models also can be multi-scale Chalupka et al.
-(2016).
-The causal model in the higher level is an abstraction of the model in the lower level.
-If the original causal model is given, then the key problem is how can we obtain an abstraction of the original model.
-Obviously, the abstraction of the causal model cannot be arbitrary, it must recover the main features of the original causal model.
-Several works proposed different standards for abstract models Rubenstein et al.
-(2017); Beckers and Halpern (2019); Chalupka et al.
-(2016), among which approximate abstraction is a good and reasonable proposal.
-The basic idea behind the method of approximation abstraction is that the differences of causal mechanism both in the abstraction and the original causal model should be minimized Rubenstein et al.
-(2017); Beckers and Halpern (2019); Beckers et al.
-(2020); Beckers (2021).
-Figure 18: An approximate abstraction of the El Niño phenomenon.
-For example, we can build a causal model to describe how the Equatorial Pacific Zonal wind ${W}$ can have a causal effect on the temperature of the Equatorial Pacific ${T}$ Chalupka et al.
-(2016).
-We write the causal variable $W$ as $X_{C}$, the effect variable $T$ as $X_{E}$, and this is the original causal model, which can be denoted as $\mathcal{M}_{m}=((\{{U}\},\{{X_{C}},{X_{E}}\},\mathcal{R}_{m}),\{g_{m},f_{m}\},\mathcal{I}_{m})$, where ${U}$ is the collection of all the unknown variables, ${X_{C}}$ and ${X_{E}}$ are causal variables, $g_{m}$ and $f_{m}$ are causal mechanisms to describe how ${U}$ and ${X_{C}}$ can affect other variables, $\mathcal{I}_{m}$ is the set of interventions.
-However, we care about higher level causality when we consider the El Nino or La Nina phenomena.
-Thus, we can build an abstract causal model $\mathcal{M}_{M}=((\{{U}\},\{Y_{C},Y_{E}\},\mathcal{R}_{M}),\{g_{M},f_{M}\},\mathcal{I}_{M})$, where $Y_{C},Y_{E}$ are new causal variables which are built by the coarse-graining mapping $\phi$, that is $Y_{C}=\phi({X_{C}})$, $Y_{E}=\phi({X_{E}})$.
-And $g_{M}$ and $f_{M}$ are causal mechanisms of $Y_{C},Y_{E}$ on the abstraction model.
-However, if $\phi$ is given, $g_{M}$ and $f_{M}$ are limited otherwise inconsistencies will exist.
-In  Beckers et al.
-(2020), the authors propose a limitation on $g_{M}$ and $f_{M}$.
-We denote $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ as the distance between different level models $\mathcal{M}_{m}$ and $\mathcal{M}_{M}$.
-They claim that $\mathcal{M}_{M}$ is an $\epsilon$-approximate abstraction of $\mathcal{M}_{m}$, if and only if $d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})$ must be less than $\epsilon$ Beckers et al.
-(2020).
-That is,
-$d_{\phi}(\mathcal{M}_{m},\mathcal{M}_{M})=\min_{\phi_{\mathcal{U}}}\max_{do(X=x)\in\mathcal{I}_{m},{u}_{m}\in\mathcal{R}_{m}(\mathcal{U}_{m})}$
-$\big{\{}d_{v_{M}}(\phi(\mathcal{M}_{m}({u}_{m},do(X=x))),\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x))))\big{\}}.$ (55)
-In this equation, $\phi_{\mathcal{U}}$ is a coarse-grained mapping of unknown exogenous variables, $do(X=x)$ is an intervention on certain endogenous variables, $d_{v_{M}}$ is the distance between variables defined at a higher level, and $\omega_{\phi}$ is the mapping of interventions between different causal models at different levels.
-The distance between models is transformed into the distance between variables.
-Given a defined method of measuring the distance between variables (such as Euclidean distance), we can measure the distance between the output of a lower level model after coarse-grained projection$\big{(}\phi(\mathcal{M}_{m}({u}_{m},do(X=x)))\big{)}$ and the variables output by a higher level model$\big{(}\mathcal{M}_{M}(\phi_{\mathcal{U}}({u}_{m}),\omega_{\phi}(do(X=x)))\big{)}$.
-We focus on identifying external variables and intervention methods that maximize the distance, and we do not concern ourselves with the coarse-grained mapping between external variables.
-Therefore, we select the optimal $\phi_{\mathcal{U}}$ and require that it be surjective Beckers et al.
-(2020).
-This yields a distance value that characterizes the inconsistency between different levels of causal models.
-There are similarities between causal emergence identification problem and abstraction of causal models.
-If the original causal mechanism can be understood as a micro-dynamics, then the abstracted mechanism can be regarded as a macro-dynamics.
-Therefore, finding the optimal $\phi$ and causal mechanisms $g_{M}$ and $f_{M}$ is equivalent to finding an optimal coarse-graining strategy and macro-dynamics.
-In the framework of NIS, we set limitations on coarse-graining strategy and macro-dynamics by requiring that the macro-dynamics can recover the micro-dynamics with $\epsilon$ error to exclude trivial cases.
-This requirement is also similar to the approximation requirement for the distance.
-In artificial intelligence, model abstraction is a widely used technique to obtain a compressed model.
-For example, the abstraction of a neural network model can be stored in cell-phone to conduct fast computation Wang et al.
-(2021); Pulina and Tacchella (2010).
-Another article discusses the importance of structural analysis methods for neural network analysis and proposes a new method based on a formal theory of causal abstraction.
-This method aligns neural representations with variables in interpretable causal models and uses interchange interventions to experimentally verify their causal properties.
-The method is utilized in a case study to examine neural models that have been trained on a complex Natural Language Inference (NLI) dataset Geiger et al.
-(2021).
-This dataset comprises numerous artificially processed, linguistically intricate English phrases and sentences.
-The analysis reveals that neural representations effectively capture the compositional structure of the provided examples Geiger et al.
-(2021).
-Model reduction is also important topics in control theory and dynamical systems Antoulas (2005).
-The primary distinction between causal emergence identification and causal model abstraction lies in their approach to abstraction or dimension reduction.
-Causal emergence identification focuses on abstracting or reducing dimensions from the original data, whereas causal model abstraction focuses on abstracting from the micro-level causal models.
-In another word, coarse-graining operates on states or data, while abstractions operate on models.
-In physics literatures Christensen and Moloney (2005), abstractions of models are also called renormalization.
-Another difference is unknown confounder $\mathcal{U}$ being considered in causal model abstraction is ignored in all the discussion of causal emergence.
-The emergence of causality requires the construction of models from a multi-scale perspective, and how to discuss mixed variables at different scales has always been a challenge.
-The theory of approximate abstraction takes into account the coarse-grained mapping of latent confounding variables $\phi_{\mathcal{U}}$, making it compatible with the coarse-grained function of endogenous variables.
-For the study of causal emergence, this theory lays a theoretical foundation for considering the confounding effects brought by environmental factors.
-The approximate abstract framework currently has many limitations.
-Firstly, there is a lack of effective algorithms that can calculate the degree of approximation in the approximate abstract.
-Given the involvement of traversing intervention sets, it is difficult for machines to perform efficient numerical calculations that are not maintainable, and cannot be used as constraint conditions for automatic modeling.
-Additionally, the intervention space is very large, and some interventions are meaningless or impossible, so it is necessary to empirically provide an intervention distribution in advance Beckers et al.
-(2020), making it difficult for this framework to be applied to real-world problems.
-4.3.2 World Models in Reinforcement Learning
-In both the generalized causal emergence identification problem and causal representation learning, a common computational framework is shared, consisting of an encoder-decoder architecture and causal mechanisms or dynamics.
-This framework can also be integrated into a reinforcement learning agent, enabling the agent to learn a world model and make informed decisions and plans.
-Such an agent, known as a world model-based reinforcement learning agent, has demonstrated impressive performance across various environments and tasks Ha and Schmidhuber (2018).
-Reinforcement learning is a significant sub-field of machine learning that possesses several distinctive features when compared to conventional supervised learning techniques Sutton and Barto (2018).
-Firstly, reinforcement learning agents have the capability to interact with the environment and actively acquire new information through interventions.
-Consequently, causal information can be obtained through these interactions Zeng et al.
-(2023).
-Secondly, in contrast to supervised learning where the processes of learning and action are clearly separated, reinforcement learning agents can seamlessly blend learning and action processes.
-Lastly, reinforcement learning agents demonstrate strong performance in environments that have delayed feedback, such as game playing scenarios Mnih et al.
-(2013); Silver et al.
-(2016).
-The reinforcement learning algorithms can be classified into model-based and model-free types, and the world model based reinforcement learning belongs to the former Ha and Schmidhuber (2018); Hafner et al.
-(2023); LeCun (2022); Hafner et al.
-(2019a, b, 2020, 2023).
-As its name suggests, there is a model of the world which can simulate the dynamics of the environment the agent confronts.
-The dynamics of the world model can be learned via the interactions between the agent and the environment, and it can also help the agent to make plans and decisions about the uncertain environment Sekar et al.
-(2020).
-However, in order to represent a complex environment, the world model needs to provide a coarse-grained description of the environment.
-Consequently, a standard architecture of a world model typically includes an encoder and a decoder, as illustrated in Figure 19.
-Figure 19: The architecture of a world model embedded in a model-based reinforcement learning agent, where $X_{t}$ is the environment at time $t$, and $Y_{t}$ is its representation, $H_{t}$ is the internal variable of the agent.
-Different from NIS and causal representation learning agent, another variable, the action in the previous step $A_{t-1}$, should be input to the world model to implement the one step dynamics.
-And the world model can be trained in a supervised learning manner by comparing the predicted representation of the environment and the prediction.
-The world model can also be treated as a macro-dynamics or causal mechanism Zhang et al.
-(2020); Allshire et al.
-(2021); Jiang et al.
-(2022); Zhang et al.
-(2020); Wang et al.
-(2022).
-All the states of the environment can be regarded as macro-states or causal representations.
-These states are constructed to ignore irrelevant information and capture the most important causal features of the environment for the agent to make decisions.
-In the process of planning, the agent can use the world model to simulate the running of the real world dynamics.
-Numeric results show that the reinforcement learning algorithms with world models can perform better than the classic reinforcement learning algorithms.
-World model-based reinforcement learning techniques can also be employed for identifying causal emergence.
-When attempting to identify the causal emergence phenomenon within a highly complex system, extracting macro states and dynamics solely from observational data becomes challenging.
-Explorations or experiments are required to facilitate this process Sekar et al.
-(2020).
-In such cases, a reinforcement learning agent equipped with a trainable world model, specifically a macro-dynamics model, becomes essential.
-By treating the complex system as a unified entity, the agent with the world model can interact with it and derive emergent causal laws from these interactions.
-On the other hand, the exploration of causal emergence can offer valuable insights into reinforcement learning algorithms.
-Firstly, the measure of causal effect, EI, can be applied to world models.
-Techniques of maximizing EI can be employed to enhance the causal effect of world models within reinforcement learning agents Yang et al.
-(2023).
-This, in turn, can greatly benefit planning and problem-solving tasks, improving their efficacy and performance.
-It is like the bi-simulation Kemertas and Aumentado-Armstrong (2021); Zhang et al.
-(2020) which can train the agent with an internal metric.
-Second, if the reinforcement learning agent with a world model can learn the emergent causal mechanisms automatically from a complex system that the agent is interacting with, then it will have more complex behaviors and be more intelligent.
-The agent may use the learnt emergent causal mechanisms to make plans.
-From learning to planning, the whole process may take place continuously.
-Therefore, causal emergence can be observed from outside because new behaviors governed by different causal mechanisms emerge as time goes by.
-This agent seems to have its own purpose or objectives.
-4.4 Summary
-In summary, there are deep connections between causal emergence and some hot topics in artificial intelligence such as CRL, abstractions, and world models.
-On one hand, the machine learning framework for causal emergence identification has similar structures to the frameworks of CRL.
-And even the concepts can have one to one correspondence as shown in Table 1.
-Therefore, the ideas and techniques of both sides can learn from each other.
-On the other hand, the studies of causal emergence can provide new perspectives and insights for causal representations and world model based algorithms.
-For example, we can explain that what the representations and causal mechanisms learned are emergent causal laws.
-Maximization of EI may help to improve the efficiency of learning the causalities for both CRL and reinforcement learning agents.
-Finally, we found a reinforcement learning agent with the ability to learn causal emergence may exhibit the ability of actions for purpose and objectives.
-5 Potential Applications in Complex Systems
-In this section, we will explore how the problems discussed in the previous sections relate to our understanding of complex systems.
-On one hand, the methods of causal emergence identification and the associated AI algorithms can be employed in complex systems to discover emergent patterns and causal relationships.
-On the other hand, the concepts of causal emergence can aid in our comprehension of complex systems, particularly in addressing significant issues such as consciousness and free will.
-These ideas have the potential to provide valuable insights and approaches to tackle these complex and profound problems.
-5.1 Potential Applications
-Complex systems typically consist of a large number of interacting units, and these units and interactions form various hierarchical and self-similar structures in a complex manner Mandelbrot and Mandelbrot (1982).
-In particular, when we examine complex systems from different scales, various structures and patterns may emerge at different scales, and these patterns may further interact with each other.
-Thus, a complex network of causality may form, and the nodes in this network may not only include the individual units of the complex system but also contains the emergent patterns in distinct scales.
-As a result, it is crucial to determine how we can identify and extract emergent patterns and causal relationships from vast amounts of data generated by complex systems.
-5.1.1 Emergence detection
+Another interesting problem is emergence detection.
 In complex dynamical systems, various higher-level patterns, such as waves, periodic oscillations, and solitons, are ubiquitous.
@@
 Identifying these emergent patterns at an early stage is crucial and significant Haugen et al.
-(2023); Mnif and Müller-Schloer (2011).
+(2023); Mnif and Müller-Schloer (2011); O’toole et al.
+(2017); Mnif and Müller-Schloer (2011); O’Toole et al.
+(2014); Pazho et al.
+(2022); Liu et al.
+(2021).
 Therefore, there is an urgent need for a method that can automatically detect these emergent patterns and even provide early warning signals.
-In O’toole et al.
-(2017), the authors regard emergent phenomenon as a specific event happened in a complex system, and detecting emergence is equivalent to detecting the formation and evaporation of this event.
-Therefore, emergence detection problem can be converted as a change-point detection problem of time series data.
-Two methods are discriminated to detect emergence: one is centralized method, the other is decentralized.
-For example, in Mnif and Müller-Schloer (2011), the authors propose that the negative Shannon entropy can be used to measure the orderness of the entire system, and it can also indicate if emergence occurs.
-NIS and NIS+ are also centralized methods.
-Another one is decentralized emergence detection method, in which the detectors are dispersed on the complex system, and implement event detection spontaneously and separately O’Toole et al. (2014).
-As the development of artificial intelligence, some deep learning methods can be used to detect emergence.
-For example, dynamical image segmentation can be used to find interesting patterns in the video automatically Liu et al.
+In this review, we mainly focus on the emergence of causality; however, we are not limited to this specific type of emergence.
+More macroscopic or global-level properties could also be emergent, although causal emergence may be the most important one.
+For example, in Chang et al.
+(2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world.
+The authors then argued that this agent with information closure can be regarded as a kind of consciousness.
+Symmetry may be another interesting property within some complex systems, and it may be emergent.
+For example, any single trajectory of a large number of random walkers in a two-dimensional Euclidean space is random.
+Nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic.
+This kind of symmetry can only be found at the macro-level and, thus, it is emergent.
+It would be interesting and useful to deploy a method for automatically discovering this kind of emergent symmetry.
+The framework for the NIS can be extended to address the problem of emergent symmetry identification if the optimized objective, EI, is replaced with a measure of isotope or more general symmetry Liu and Tegmark (2022).
+Similarly, we can also find other emergent symmetries using this framework once their measurement and optimization are possible.
+However, it is crucial to acknowledge the inherent limitations of applying machine learning techniques to causal discovery and emergence identification problems.
+For instance, in Textor et al.
+(2016), the authors highlighted the existence of ’statistically equivalent’ but causally distinct DAGs, implying that different causal structures can be constructed to account for the same dataset.
+Consequently, when utilizing machine learning techniques to uncover causality and emergent properties, similar challenges need to be addressed.
+Further research in this direction warrants significant attention to overcome these obstacles.
+5.4 Understanding Complex Systems from Causal Emergence
+A profound understanding of causal emergence and emergent causality can provide insights into understanding various mysterious phenomena in complex systems, including living systems, social systems, climate systems, ecosystems, and more.
+Significantly, fundamental questions, such as free will Yurchenko (2023), consciousness Tononi and Koch (2015), and life, are all intricately connected to emergent causation Walker (2014).
+For instance, free will can be viewed as an emergent form of downward causation Yurchenko (2023).
+Social phenomena can be comprehended through the lens of causal emergence Elder-Vass (2010).
+Interestingly, both the EI and $\phi$ID frameworks for causal emergence have connections with one of the theories of consciousness, namely the integrated information theory Tononi et al.
 (2016).
-A complex system can always be represented as a dynamical graph, e.g., a flock of birds, and when emergence occurs, the connectivity of this interaction network always increase.
-Therefore, we can detect the change of the network to find emergence Pazho et al.
-(2022).
-Other deep learning methods could be helpful.
-5.1.2 Discovering emergent causality
-Complex systems can be understood as a large network with events connected by causal links.
-A small disturb of one event may give rise a large roam of changes as butterfly effect indicates.
-While, sometimes, one single event happens with destiny which means it may be the effect of many different causes.
-How to discover the complex causal relationships from data in an automatic way is a challenging problem.
-The research area of causal discovery tries to tackle this problem in many different ways Glymour et al.
-(2019); Zanga et al.
-(2022).
-For example, some studies try to use well-defined statistical indicators to characterize the causality between two variables, e.g., the measures of mutual information Donges et al.
-(2009), transfer entropy Stetter et al.
-(2012), etc.
-Nevertheless, these methods can not distinguish causal relationship from correlations.
-Another widely used method is independent causal discovery method Glymour et al.
-(2019), which tries to find the conditional independence among three variables in data, and derive Markov equivalent sub-graphs accordingly.
-Finally, binary causal interactions can be identified from these sub-graphs Kalisch and Bühlman (2007).
-Score based method tries to maximize or minimize a score function to find the optimal causal structures Zheng et al.
-(2018); Zhu et al.
-(2019).
-These methods always require a large amount of data and computation.
-However, more challenges will be confronted when these causal discovery methods are applied in complex systems because circular causal structures and cross-level causality exist.
-Circular causal structures widely exist in complex dynamical systems due to the feedback effect.
-That is, one variable may take effect on itself via a direct self-loop feedback or a circle with long chain.
-This circular causal structure may challenge the existed study methods of causality because most of these methods study directed acyclic causal graphs.
-Recent progress has developed some methods to discover these circular structures e.g.
-Richardson (2013); Lacerda et al.
-(2012); Forré and Mooij (2018); Rantanen et al.
-(2020); Zhang et al.
-(2019); Pamfil et al.
-(2020); Vowels et al.
-(2022).
-Another aspect is higher level or cross-level causation if the ingredients of scale and coarse-graining are considered.
-For example, downward causation describes the causal effect between macro-level and micro-level variables.
-Thus, the causal connections may be cross-leveled.
-If causal emergence occurs in a complex system, some strong causality may be found between macro-variables.
-All of these factors in complex systems present new challenges.
-The machine learning methods discussed in section 4.1.2 can address the question of whether there exists higher-level causality at the macro-level.
-However, it’s important to note that this is a global property.
-It must be further developed and extended to find local causal relationship between macro-variables or macro-variable and micro-variable Varley (2022).
-While, existed methods of causal discovery must be extended to take the method of grouping a set of variables or coarse-graining a system into consideration.
-For example, the article Cheng et al.
-(2023) proposes a coarse-to-fine causal discovery algorithm based on Granger causality and graph neural network.
-The grouping process of variables is from coarse-to-find which can improve the efficiency of the algorithm.
-Other multi-level causal discovery methods are also proposed and discussed in Wang et al.
-(2023); Zhang et al.
-(2023); Fan et al.
-(2023).
-One of the potential applications of these complicated causal discovery method is to enhance the interpretabiity of large and complex neural networks.
-As we know, interpretability is one of the notorious problem for large neural networks, although they are very useful to solve many real problems.
-Thus, we can apply the methods of causal discovery, especially, the methods to discover emergent causality on these neural network models to find hidden and emergent causal relationships between neurons or group of neurons Chattopadhyay et al.
-(2019); Ren et al.
-(2023).
-If the degrees of causal effect can be computed and compared among these discovered causal relationships, the most important causation can be pointed out to solve a concrete problem.
-This may enhance the interpretability of the neural network.
-In this way, the misery of emergent ability of large language models may be dismissed by uncovering the hidden causation within large models Michaud et al.
-(2023); Wei et al.
-(2022).
-5.1.3 Beyond causality
-In this review, we mainly focus on the emergence of causality, however, we do not need to be bounded on this kind of special emergence, more macroscopic or global level properties could also be emergent although causal emergence may be the most important one.
-For example, individuality is one of the emergent property with no any connection with causality which has been discussed in previous section.
-In  Chang et al.
-(2020), the authors discussed the concept of emergent information closure, which refers to the idea that the encoded information within an agent can form a closed system that is not influenced by the external world.
-The authors then argued that this agent with information closure be regarded as a model of consciousness.
-Symmetry may be another interesting property within some complex systems, and it may be emergent.
-For example, any single trajectory of a large number of random walkers in a two dimensional Euclidean space is random, nevertheless, the Gaussian distribution surface describing how the number of walkers falling in each small region changes with different locations is isotropic.
-This kind of symmetry can only be found on the macro-level, thus, is emergent.
-It should be noted that EI can also be viewed as a unique indicator of symmetry, specifically, it is time reversal symmetry.
-The reason behind this is that when the macro-dynamics approaches a time reversible map (a bijection between states across a time step), EI is increased.
-And the maximum EI can be arrived when the Markov dynamics is reversible.
-This interesting fact should be further studied in the future.
-It is interesting and useful to deploy a machine learning method to automatically discover this kind of emergent symmetry.
-The framework of NIS may be extended to account for the problem of emergent symmetry identification if the optimized objective, EI, is replaced with some measure of isotope or more general symmetry Liu and Tegmark (2022).
-Similarly, we can also find other emergent symmetry with this framework once the symmetry can be measured and optimized.
-5.1.4 Designing Emergence
-We even can design a set of interaction rules for a complex system such that the desired property can emerge from the interactions once the desired property can be defined in a clear and quantitative way.
-The basic idea to realize emergence design is by optimization Adamatzky et al.
-; Kreyssig and Dittrich (2011).
-Actually, the early studies of evolutionary computation or genetic algorithm have solved this problem for some simple and small cases Adamatzky et al.
-.
-For example, M.
-Mitchell et al.
-used genetic algorithm to search the interaction rules of one dimensional cellular automata such that the automata can solve some simple mathematical problems Mitchell et al.
-(1993).
-Recently, as the development of machine learning and automatic differential, searching such interaction rules may become easier and more efficient if the rules can be converted into differentiable operations Mordvintsev et al.
-(2022); Gala et al.
-(2023); Zhang and Walker (2022).
-However, if the space of possible interaction rules is very large, all the methods based on search or optimization may be invalid.
-Thus, some ideas developed in the field of deep learning can be borrowed here to design emergence.
-For example, the framework of generative adversarial network (GAN) can be used to design complex systems to realize some emergent properties Goodfellow et al.
-(2014).
-On one side, we should design a generative model for a complex system or the interaction rules for the system such that a large number of candidate rules or systems can be generated.
-On the other side, a discriminator or a critic can be designed to evaluate the candidates if the desired emergent properties have been realized.
-Similar ideas can be used on other possible generative models, e.g., diffusion model, normalizing flow Cao et al.
-(2022); Yang et al.
-(2022); Papamakarios et al.
-(2021); Kobyzev et al.
-(2020), etc.
-However, the difficulty of these methods lies in the demand for a large amount of training data and computations.
-The lack of data may hamper the applications.
-In general, designing emergence is a useful problem but it is also difficult.
-5.2 Understanding Complex Systems from Causal Emergence
-Deep understanding of causal emergence and emergent causality will help us to understand the hidden mechanism of complex systems including living systems, social systems, climate systems, ecosystems, etc.
-Particularly, some big problems like free will, consciousness Tononi and Koch (2015), and life are all related to emergent causation Walker (2014).
-For example, free will can be understood as an emergent downward causation Yurchenko (2023).
-In Chinese medicine, Yin and Yang can be understood as macro-states and the properties of their interactions are emergent causality.
-Social phenomena can be understood from the viewpoint of causal emergence Elder-Vass (2010).
-For instance, the increase of the rate of suicide may be the emergent causal effect of economic crises.
-Interestingly, both EI and $\phi$ID frameworks on causal emergence have relation with one of the consciousness theory: integrated information theory Tononi et al.
-(2016).
-As a result, further studies of causal emergence and the connections with other big problems are required.
-For a particular system, there are three problems that should be addressed in future studies: 1.
-When does causal emergence occurs?
-2.
-How emergent causality take functional effects on the system?
-3.
-How emergent causality may change when the system is changed to adapt the environment?
-We will discuss these problems one by one in brief.
-We still do not know when will causal emergence occurs and how does the measure of causal emergence change with some key parameters of a system.
+However, understanding these abstract concepts and phenomena in the context of causal emergence is merely an initial step.
+For a specific system, there are three problems that should be addressed in future studies: (1) When does causal emergence occur?
+(2) How does emergent causality have functional effects on the system?
+(3) How does emergent causality change when the system is changed to adapt to the environment?
+We discuss these problems one by one.
+We still do not know when causal emergence will occur and how the measure of causal emergence changes with some key parameters of a system.
 In Yang et al.
-(2023), the authors have shown how the measure of causal emergence change with different noises in an extended Boid model.
-It increases as the observational noise increases because coarse-graining operation can eliminate the observational noise.
-However, the same coarse-graining operation can hardly reduce the intrinsic noise (the noise within the dynamical equations), which leads to the degree of causal emergence decreases.
-It is reasonable to expect that there is a phase transition of the emergent causality within a complex system when some key parameter changes because the causality or the strength of causal effect is also a global property and it may dependent on some order parameters.
-If causal emergence occurs in the system, how does the emergent causation take effects on the part and the whole of the system?
-For example, the emergent "I" can be understood as an emergent macro-variable Yurchenko (2023).
-How does "I" can influence the part, say my foot, to move?
-This problem is non-trivial because it relates to the problem of mind-body interaction.
-One of the difficulty is that how can we build a model for such systems?
-Should we directly model the emergent properties and causation or model the underlying interaction rules and ignore the causation in higher level?
-If we directly model the emergent "I", the model may become very complex because the properties of "I" are difficult.
-However, if we model the system from the micro-level interactions, we can still hardly understand the emergent causality.
-Shall we switch the viewpoints between micro- and macro-?
-If yes, in what manner to switch?
+(2023), the authors showed how the measure of causal emergence changes with different noises in the Boid model (see the relevant discussion in Section 4.1.2).
+It is reasonable to expect that there is a phase transition of emergent causality within a complex system when some key parameter changes because the causality or strength of the causal effect is also a global property, and it may be dependent on some order parameters.
+If causal emergence occurs in the system, how does the emergent causation affect the parts and the whole of the system?
+For example, the emergent “I” can be understood as an emergent macro-variable Yurchenko (2023).
+How does “I” influence the part, say the foot, to move?
+This problem is non-trivial because it relates to the problem of mind–body interaction.
+It is important to study the information flows at both the macro-level and micro-level together to understand this phenomenon.
 Finally, what is the relationship between adaptation and emergence Holland (1995, 2000, 1992)?
-Sometimes, when we refer to a property as emergent, we actually mean that this property can be developed through adaptation.
+Sometimes, when we refer to a property as emergent, we essentially mean that this property can be developed through adaptation.
 Therefore, adaptation or evolution serves as a causal force for certain emergent properties.
@@
 For instance, downward causation, which is commonly found in complex systems, emerges as a result of adaptation and evolution.
-The next crucial question is: how can we evolve an emergent property or causation?
-This problem resembles the issue of designing emergence discussed in Section  5.1.4.
+The next crucial question is, how can we evolve an emergent property or causation?
+This problem resembles the issue of designing emergence Adamatzky et al.
+; Kreyssig and Dittrich (2011); Mitchell et al.
+(1993).
 However, our aim here is to seek an explanation rather than a design.
 We want to understand the specific environment and manner in which the observed emergent causation can be evolved through adaptation.
-All these problems desire further studies in future. However, these problems are just name a few, more interesting problems need to be addressed.
-6 Conclusion
-Quantifying the notion of emergence has been a challenge for a long time.
-Causal emergence serves as an important step in the development of a quantitative theory of emergence, making it a crucial concept in the field of complex system studies.
-In this paper, we provide a systematic review of papers on causal emergence and other related quantitative studies pertaining to emergence and causality.
-Our analysis highlights two crucial problem areas: emergence quantification and causal emergence identification in data.
-We discovered intriguing connections between the latter problem and artificial intelligence, particularly in sub-fields such as causal representation learning and world-model based reinforcement learning.
-As artificial intelligence continues to advance, we anticipate that more puzzles related to causality and cross-level interactions within complex systems will be solved.
-7 Appendix A - Calculation of EI with TPM
+All these problems require further studies in the future. However, these problems are just a few of the more interesting problems that need to be addressed.
+[…]
+Conceptualization, J.Z.; methodology, J.Z., B.Y., P,C,; writing, B.Y., J.Z., A.L., J.W., Z.W., M.Y., K.L., M.M. All authors have read and agreed to the published version of the manuscript.
+[…]
+This work was supported in part by National Natural Science Foundation of China (No. 62141607, U1936219).
+[…]
+Not applicable.
+[…]
+No data available.
+Acknowledgements.
+[…]
+Appendix A Calculation of EI with Transition Probability Matrix
 Calculating effective information ($EI$) requires knowledge of the joint probability distribution of the system.
@@
 $TPM(i,j)=P(X_{t+1}=s_{j}|X_{t}=s_{i})\\ =\frac{P(X_{t+1}=s_{j},X_{t}=s_{i})}{P(X_{t}=s_{i})}$ (53)
-Assuming we can "force" the system to conform to a maximum entropy distribution (i.e., uniform distribution) with $n$ states, the probability of being in any given state $i$ at time $t$, denoted by $P_{u}(X_{t}=s_{i})$, is $1\over{n}$.
+Assuming we can “force” the system to conform to a maximum entropy distribution (i.e., uniform distribution) with $N$ states, the probability of being in any given state $i$ at time $t$, denoted by $P_{u}(X_{t}=s_{i})$, is $1\over{N}$.
 As a result, the TPM can be expressed as the product of the joint probability of the system’s maximum entropy distribution, denoted by $P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})$, and the total number of states, which is $N$, as demonstrated below.
@@
 Hence, the TPM of the system can be used to express both the joint probability and marginal probability at time $t+1$ of the system’s maximum entropy distribution.
+$\therefore P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})=\frac{1}{N}TPM(i,j)$ (55)
+$\therefore P_{u}(X_{t+1}=s_{j})=\sum_{i}P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})\\ =\frac{1}{N}\sum_{i}TPM(i,j)$ (56)
+Using this relationship, $EI$ can be calculated directly with the TPM.
+$EI=I(I_{D};E_{D})$ $=I(X_{t};X_{t+1})|_{do(X_{t})\sim U}$ (57)
+$=\sum_{i,j}P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})\log_{2}\Big{(}\frac{P_{u}(X_{t+1}=s_{j},X_{t}=s_{i})}{P_{u}(X_{t+1}=s_{j})P_{u}(X_{t}=s_{i})}\Big{)}$
+$=\sum_{i,j}\frac{1}{N}TPM(i,j)\log_{2}\Big{(}\frac{\frac{1}{N}TPM(i,j)}{\frac{1}{N}\sum_{k}TPM(k,j)\times\frac{1}{N}}\Big{)}$
+$=\frac{1}{N}\sum_{i,j}TPM(i,j)\log_{2}\Big{(}\frac{N\times TPM(i,j)}{\sum_{k}TPM(k,j)}\Big{)}$
 […]
-[…]
-Using this relationship, $EI$ can be calculated directly with the TPM.
-[…]
-[…]
-[…]
-[…]
-Reference